### v3

# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

Abstract
Negation is a fundamental linguistic phenomenon that poses ongoing challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Current benchmarks often treat negation as a minor detail within broader tasks, such as natural language inference. Consequently, there is a lack of benchmarks specifically designed to evaluate comprehension of negation. In this work, we introduce Thunder-NUBench — a novel benchmark explicitly created to assess sentence-level understanding of negation in LLMs. Thunder-NUBench goes beyond merely identifying surface-level cues by contrasting standard negation with structurally diverse alternatives, such as local negation, contradiction, and paraphrase. This benchmark includes manually curated sentence-negation pairs and a multiple-choice dataset, allowing for a comprehensive evaluation of models’ understanding of negation.
[…]
Thunder-NUBench: A Benchmark for
LLMs’ Sentence-Level Negation Understanding
Yeonkyoung So, Gyuseong Lee, Sungmok Jung, Joonhak Lee, JiA Kang, Sangho Kim, Jaejin Lee Graduate School of Data Science, Seoul National University {kathy1028,ksnannaya,tjdahrwjd,hmjelee,jia6776,ksh4931,jaejin}@snu.ac.kr
1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely associated with various human communicative abilities, such as denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation still plays a significant role in language; approximately 25% of sentences in English texts contain some form of negation (sarabi2016understanding; hossain2020analysis; sep-negation). This prevalence and its impact on meaning make accurate interpretation of negation crucial for several natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI) (khandelwal2020negbert; hosseini2021understanding; singh2023nlms). Recent studies have shown that effectively managing negation is important even for multimodal language models (quantmeyer2024and; alhamoud2025vision; park2025know).
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find it more difficult to process and comprehend negated statements compared to affirmative ones (wales1969so; sarabi2016understanding). Similarly, many studies indicate that pretrained language models (PLMs) struggle to interpret negation accurately. For example, models like BERT (devlin2019bert) and even large language models (LLMs) such as GPT-3 (brown2020language) often have difficulty distinguishing between negated and affirmative statements. These models tend to rely on superficial cues, which can result in incorrect outputs when negation is involved (kassner2020negated; hossain2022analysis; truong2023language).
Figure 1: An example of Thunder-NUBench multiple-choice evaluation task, where the underlined text indicates the main verb phrase of each sentence, and the red text marks the negated part.
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources either treat negation as a minor aspect of broader tasks or focus solely on narrow syntactic detection, often emphasizing encoder-based models (hossain2020analysis; geiger2020neural; truong2022not; anschutz2023not). To address this gap, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset explicitly designed to evaluate LLMs’ sentence-level comprehension of negation. Our benchmark is structured as a multiple-choice question (MCQ) task: given an original sentence, the model must select the correct standard negation from four options. The other three choices—local negation, contradiction, and paraphrase—are carefully designed distractors that test whether models truly grasp semantic scope and logical oppositions.
The contributions of this paper are summarized as follows:
  • We define standard negation within the framework of sentential logic, moving beyond the cue-based and often ambiguous accounts of prior work. Grounding standard negation in logical structure not only clarifies its role in natural language but also supports the evaluation and enhancement of reasoning in LLMs.
  • We create a manually curated benchmark that includes a dataset of sentence-negation pairs for fine-tuning, along with a multiple-choice evaluation task.
  • We conduct systematic evaluations of decoder-based LLMs, assessing their performance under both prompting and supervised fine-tuning. This includes error and confusion analyses that highlight the models’ ongoing challenges with negation.
Thunder-NUBench provides valuable insights into the semantic reasoning abilities of language models and serves as a robust standard for future research focused on understanding negation.
2 Related Work
Negation detection and scope resolution.
Early work in negation detection and scope resolution primarily relied on rule-based systems and handcrafted heuristics, especially in domain-specific contexts like clinical texts. While these systems are effective, they lack flexibility across different domains (chapman2001simple; de2012ucm; ballesteros2012ucm; basile2012ugroningen). Traditional machine learning methods, such as Support Vector Machines (SVMs) (hearst1998support) and Conditional Random Fields (CRFs) (sutton2012introduction), were introduced later; however, they too are limited to narrow domains (morante2008learning; morante2009metalearning; read2012uio1; li2018learning).
More recently, deep learning approaches employing Convolutional Neural Networks (CNNs) (o2015introduction) and Bidirectional Long Short-Term Memory (BiLSTM) networks (siami2019performance) have enhanced performance by providing improved contextual embeddings and sequence modeling (fancellu2016neural; bhatia2019end). Pretrained transformer models like BERT have been employed through transfer learning techniques (e.g., NegBERT (khandelwal2020negbert)), significantly increasing the accuracy of negation detection tasks. Nonetheless, these methods still largely focus on syntactic span detection, leaving deeper semantic understanding of negation a challenging area to tackle.
Dimension Negation Type Definition Example
Scope
Clausal Negation
( = Sentential
Negation)
Negation that applies to the entire clause or
sentence. This typically involves the use of "not",
or its contracted form "n’t" with auxiliary verbs.
He speaks English fluently.
$\rightarrow$ He doesn’t speak English fluently.
Subclausal Negation
( = Constituent /
Local Negation)
Negation that focuses on negating a specific
part of a clause, such as a word or phrase,
rather than the entire clause.
He speaks English fluently.
$\rightarrow$ He speaks English, but not fluently.
Form
Morphological
Negation
Negation expressed through affixes attached to words
such as prefixes like "un-", "in-", "dis-", or suffixes
like "-less".
She is happy.
$\rightarrow$ She is unhappy.
Syntactic Negation
Negation expressed through separate words (particles)
in the syntax, such as "not", "never", "no", etc.
She is happy.
$\rightarrow$ She is not happy.
Target Verbal Negation
Negation that applies directly to the verb
or verb phrase.
They have finished the work.
$\rightarrow$ They have not finished the work.
Non-verbal
Negation
Negation that negates elements other than the verb.
There is milk in the fridge.
$\rightarrow$ There is no milk in the fridge.
Table 1: Typology of negation.
Negation-sensitive subtasks of NLU.
Negation understanding has become increasingly important in natural language understanding (NLU) tasks (hosseini2021understanding). However, existing NLU benchmarks, such as SNLI (bowman2015large) for natural language inference (NLI), CommonsenseQA (talmor2019commonsenseqa) for Question Answering (QA), SST-2 (socher2013recursive) for sentiment analysis, STS-B (cer2017semeval) for textual similarity and paraphrasing, have been criticized for not adequately addressing the semantic impact of negation (hossain2022analysis; rezaei2024paraphrasing). These datasets contain relatively few instances of negation or include negations that are not crucial to task performance, allowing language models to achieve high accuracy even when they completely ignore negation.
Recent studies, including NegNLI (hossain2020analysis), MoNLI (geiger2020neural), and NaN-NLI (truong2022not), have introduced benchmarks for NLU that are sensitive to negation. These studies show that model performance significantly declines when negation plays a crucial role in affecting the outcome (naik2018stress; yanaka2019can; hartmann2021multilingual; hossain2022question; hossain2022leveraging; she2023scone; anschutz2023not). These findings suggest that current language models tend to depend on superficial linguistic patterns rather than a genuine understanding of semantics.
Limitations of distributional semantics.
Distributional semantics (harris1954distributional; sahlgren2008distributional) aims to create models that learn semantic representations based on patterns of word co-occurrences (boleda2020distributional; lenci2022comparative) and capture broad semantic relationships; however, it encounters significant challenges with negation. Negated expressions, such as "not good," often appear in similar contexts as their affirmative counterparts, like "good." As a result, models tend to generate similar vector representations for these expressions, despite their opposing meanings. Previous research has pointed out this limitation, showing that PLMs struggle to capture the subtle semantic differences introduced by antonyms and the reversal of polarity (rimell2017learning; jumelet2018language; niwa2021predicting; jang2022beyond; vahtola2022not). Studies have further suggested that models like BERT find it difficult to distinguish between affirmative and negated contexts (kassner2020negated; ettinger2020bert).
Negations in generative language models.
Recent research on understanding negation has primarily focused on bidirectional models, such as BERT (devlin2019bert) and RoBERTa (liu2019roberta), which have demonstrated strong performance in NLU and negation detection tasks. However, with the emergence of generative foundation models like GPT (radford2018improving) and LLaMA (touvron2023llama), attention has shifted towards evaluating how these models handle negation. Studies have shown that these generative models often exhibit a positive bias and struggle with producing or interpreting negated statements (truong2023language; chen2023say; garcia2023not). Although some benchmarks, such as CONDAQA (ravichander2022condaqa) and ScoNe (she2023scone), reveal these limitations, there is still a lack of robust evaluation resources specifically designed for generative models.
Building on previous studies, this paper assesses whether generative models can comprehend negation in complex sentences and identify semantic differences that extend beyond surface-level patterns.
Type Definition
Base case
If $P$ is an atomic proposition, $\text{Neg}(P)$ is the proposition where the main predicate of $P$ is negated.
Inductive step Conjunction $\text{Neg}(P\;\text{and}\;Q)\equiv\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$,  $\text{Neg}(P\;\text{but}\;Q)\equiv\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$
Disjunction $\text{Neg}(P\;\text{or}\;Q)\equiv\text{Neg}(P)\;\text{and}\;\text{Neg}(Q)$
Implication $\text{Neg}(\text{if }P,Q)\equiv\text{Neg}(\text{Neg}(P)\;\text{or}\;Q)\equiv P\;\text{and}\;\text{Neg}(Q)$ $\text{Neg}(P\;\text{if and only if}\;Q)\equiv\text{Neg}(\text{if }P,Q\;\text{and}\;\text{if }Q,P)$
Table 2: Standard negation. $P$ and $Q$ stand for propositions. In addition to and, or, and if, other natural language connectives such as when are also considered, and their negations follow the same principles depending on their function.
3 Scope and Categorization of Negation
In this work, we aim to clarify the concept of negation by introducing a typology that clearly outlines its semantic boundaries and differentiates it from related, yet distinct, phenomena. This typology organizes various forms of meaning reversal into logically consistent categories, allowing for a more precise and systematic evaluation of how language models handle negation.
3.1 Typology of Negation
Negation is a fundamental semantic and syntactic operation found in natural languages, used to convey denial, rejection, or the absence of a proposition. Hereafter, we denote our negation operation for a sentence $S$ as $\text{Neg}(S)$. In formal logic, negation flips the truth value of a proposition $P$: if $P$ is true, then $\text{Neg}(P)$ is false, and vice versa. Semantically, negation creates a binary opposition between a proposition and its affirmative counterpart, meaning that each one is the opposite of the other (sep-negation).
Negation can be categorized along several dimensions: scope, form, and target (see Table 1). In terms of scope, negation may affect the entire clause (referred to as clausal negation) or only part of it (known as subclausal negation). Regarding form, negation can manifest as bound morphemes, such as prefixes and suffixes (morphological negation), or as separate syntactic elements like "not" or "never" (syntactic negation). Finally, depending on its target, negation can apply to the verb (verbal negation) or to other elements in the sentence (non-verbal negation(zanuttini2001sentential; miestamo2007negation; truong2022not; kletz2023self).
3.2 Negation and Contradiction
Negation and contradiction are closely related concepts that are often conflated in NLP research (jiang2021m). Contradiction refers to the incompatibility of two propositions, meaning that they cannot both be true at the same time. While negation frequently serves as a primary mechanism for creating contradictions—by reversing the truth value of a proposition—contradictions can also arise from antonymy, numeric mismatches, or differences in structure and lexicon (further details can be found in Appendix A). For instance, the statements "An individual was born in France" and "An individual was born in Italy" are contradictory, but they are not negations, as the second statement does not reverse the truth of the first.
Many previous studies have overlooked the possibility that contradictions can exist independently of explicit negation. Recognizing this gap, we specifically examine the ability of LLMs to differentiate between negations and non-negated contradictions, highlighting the nuanced semantic distinctions that are involved.
Type Structure Explanation Local Negation Example
Relative clause
negation
A relative clause is a type of dependent clause that gives
extra details about a noun or noun phrase in the main sentence.
It usually begins with a relative pronoun such as
who, which, that, whom, or whose.
The man who owns the car
is my neighbor. $\rightarrow$ The man
who does not own the car is
my neighbor.
Participle clause
negation
A participle clause is a type of dependent clause that begins with
a participle (a verb form ending in -ing or a past participle).
It acts like an adverb, giving extra details about the main
clause, often showing time, reason, result, or sequence of actions.
Walking through the park, she
found a lost wallet. $\rightarrow$ Not
walking through the park, she
found a lost wallet.
Adverbial clause
negation
An adverbial clause is a dependent clause that acts like an adverb,
modifying a verb, adjective, or adverb. It gives information such
as time, reason, condition, or contrast. These clauses are introduced
by subordinating conjunctions like because, although, or while.
She stayed inside because it was
raining. $\rightarrow$ She stayed inside
because it was not raining.
Compound
sentence
with
local negation
A compound sentence consists of two or more main clauses
joined by coordinating conjunctions such as and, but, or or.
If only one of these clauses is negated, the negation applies
only locally to that clause.
He submitted the report and
attended the meeting. $\rightarrow$ He
submitted the report and did
not attend the meeting.
Table 3: Typology of local negation.
3.3 Standard Negation
Standard negation refers to the typical form of negation applied to the declarative verbal main clause. It specifically negates the verb in a main clause (miestamo2000towards). A main clause can function as a complete sentence on its own, consisting at minimum of a subject and a predicate. This definition is grounded in the notion that the verb acts as the head of the clause (miller2011critical).
Building on this traditional understanding, we treat standard negation as the process of reversing the truth value of the verb phrase in the main clause, which we will refer to as the main predicate in this paper. A verb phrase is headed by a verb and can consist of a single verb or a combination of auxiliaries, complements, and modifiers (e.g., "will call" and "is being promoted") (lakoff1966criterion). Since the main predicate conveys the core action or state of the clause, negating it effectively reverses the proposition of the entire sentence. In this context, we treat standard negation as a truth-functional operation that maps the main predicate to its complement set within the semantic space.
We further clarify the scope of standard negation within the typology presented in Table 1. Standard negation includes both clausal negation and verbal negation, as it reverses the meaning of the entire sentence by negating the main predicate. In terms of form, standard negation can employ both syntactic and morphological negation.
Syntactically, standard negation often uses explicit negation particles, such as "not." Morphologically, it can involve complementary antonyms (for example, "alive" vs. "dead" or "true" vs. "false"), which occupy mutually exclusive semantic spaces, thus reversing the truth value of the proposition.
In contrast, other types of antonyms, such as gradable antonyms (e.g., "happy" vs. "unhappy") and relational antonyms (e.g., "buy" vs. "sell") (lehrer1982antonymy), do not strictly reverse truth values. Therefore, they are classified as contradictions rather than standard negation in this paper.
Atomic propositions.
While this characterization effectively defines standard negation for atomic propositions (elementary sentences that cannot be further decomposed) (davis2004semantics), its application to complex sentences with multiple clauses requires a more thorough approach. In this paper, we treat an atomic proposition as a sentence that contains a single main predicate.
Complex propositions.
Specifically, for propositions composed of multiple logically connected atomic statements, the method for reversing the truth value of the entire complex proposition can be ambiguous. In natural language, such logical structures typically appear as coordinated clauses (e.g., "$P$ and $Q$ or RR") or comma-separated lists connected by "and" or "or" (e.g., "$P,Q,$ and RR"). We treat these as equivalent to a sequence of binary conjunctions or disjunctions.
Figure 2: Dataset generation process.
Definition of standard negation.
In this paper, standard negation refers to natural-language sentential negation, which is formally treated as logical negation within the framework of sentential logic (enderton2001). To address the complexities involved, we define standard negation recursively by applying it pairwise over the logical structure of a sentence until only atomic propositions remain, ensuring that the truth value of the entire sentence is reversed even when it contains multiple coordinated clauses. Our definition of $\text{Neg}(\cdot)$ is presented in Table 2. Conditionals of the form "if $P$, QQ" are equivalent to $\text{Neg}(P)\;\text{or}\;Q$" in logic, and we adhere to this equivalence when defining their negation (more details can be found in Appendix C).
Our definition of standard negation is inspired by the framework of sentential logic (enderton2001). However, it should be viewed as an operational definition rather than a strict mathematical formulation. Natural language sentences often lack explicit structural markers, such as parentheses, which are vital in well-formed logical formulas. Furthermore, coordination can appear with or without commas. Unlike formal languages, natural languages do not adhere to strict formation rules, making it challenging to map their structures unambiguously to logical formulas. Consequently, our definition cannot perfectly align with the forms used in sentential logic. Nevertheless, it provides a clear operational account of standard negation in natural languages.
3.4 Local Negation
We define local negation as a form of negation that specifically targets a verb phrase outside the main clause. While the term is often used interchangeably with subclausal negation, our focus is solely on local negation relating to subclausal and verbal negation. This concept applies to four types of sentence structures: relative clauses, participle clauses, adverbial clauses, and compound sentences (refer to Table 3 for more details).
In particular, conditional clauses, such as the "if PP" part in "if $P$, QQ" are categorized as adverbial clauses. In compound sentences, standard negation requires all main clauses to be negated in order to achieve sentence-level negation. If only a subset of the clauses is negated, this is considered local negation.
Local negation, in terms of structure, resembles standard negation, typically using explicit negation markers like "not." However, its scope is confined to a specific part of the sentence rather than encompassing the entire main clause. Because explicit cues such as "not" are still present, models that depend on shallow cue detection may be misled, failing to distinguish between standard negation and local negation.
4 Thunder-NUBench Dataset
Category Description
Standard Negation This category involves reversing the truth value of the main clause, which is the primary focus of the benchmark.
Local Negation In this case, negation is applied to a subordinate clause or a partial structure, which does not reverse the entire sentence.
Contradiction This category introduces conflicts with the original meaning through semantic changes, such as the use of antonyms, different numbers, or other entities, without employing explicit negation.
Paraphrase Here, the original meaning is preserved while the surface form is altered. Examples of paraphrases are intentionally constructed to vary the sentence structure and word choice significantly, ensuring that no additional information is added. As a result, the original sentence still entails its paraphrase. This category tests whether models mistakenly consider different surface forms as meaning reversals, even when the semantic meanings remain equivalent.
Table 4: Multiple choice categories included in Thunder-NUBench.
We construct the Thunder-NUBench dataset through three main stages: (1) pre-processing, (2) generation, and (3) review. The overall workflow is illustrated in Figure 2.
Pre-processing.
We begin by extracting sentences from two primary corpora: (1) the Hover dataset (jiang2020hover), designed for multi-hop fact extraction and claim verification, and (2) the Wikipedia Summary dataset (scheepers2017compositionality), which contains concise summaries from English Wikipedia. We chose these datasets because their factual content and complex sentence structures are well-suited for developing a dataset aimed at understanding standard negation in complex, sufficiently lengthy sentences. Additionally, we automatically correct any grammatical errors and merge or split sentences as needed to create well-formed single-sentence units.
Generation.
We create two types of datasets from the pre-processed sentences: the sentence-negation pair dataset and the multiple choice dataset. In the sentence-negation pair dataset, each original sentence is paired with a manually crafted standard negation, as detailed in Section 3.3. In the multiple-choice dataset, each original sentence is presented with four options: a standard negation, a local negation, a contradiction, and a paraphrase. Each of them are described in Table 4. Together, these categories assess whether models truly understand semantic negation rather than relying on superficial cues.
Standard and local negation options are manually created rather than generated by LLMs. We have observed that LLMs often struggle to produce correct standard negations, frequently resulting in subclausal or local negations instead. They can also generate incorrect local negations, even when explicitly prompted to do otherwise. Since precise negation is essential to our benchmark, these options must be developed by humans to ensure the quality of the dataset. In contrast, contradiction and paraphrase options are initially created automatically using carefully designed prompts with the OpenAI API (OpenAI) and are then refined during the review process.
Review.
All constructed data undergo a multi-stage human review process (see Appendix I). A different author, separated from the creator, cross-checks each instance, and any disagreements are addressed in regular meetings to ensure consistency. Options for contradictions are reviewed only after the corresponding standard and local negations are finalized, as they must not overlap semantically. Consequently, the earlier negations are re-examined during the contradiction review and are cross-checked by multiple authors.
The guidelines for data generation and review are continuously updated, and any previously created data are revised accordingly (see Appendix J). This protocol ensures rigorous quality control and consistency throughout the benchmark.
Dataset statistics.
The final dataset includes a training set of sentence-negation pairs and a multiple-choice evaluation set (see Table 5). For few-shot prompting, we construct a demonstration set of 50 examples. These are carefully selected to have unique Wikipedia page indices to avoid any overlap with the test set. Furthermore, to provide the model with a balanced overview of the task, we match the distribution of local negation types (choice2_type) in the demonstration set to that of the overall dataset. This ensures that the demonstrations are representative and prevents the model from developing a biased strategy for specific negation types.
Dataset Split Count
Sentence-Negation Train 3,772
Multiple Choice Demonstration 50
Test 1,261
Total 5,083
Table 5: Thunder-NUBench statistics.
5 Experiments
5.1 Evaluation Setup
We evaluate models under two common Multiple-Choice Question Answering (MCQA) evaluation settings: (1) a completion-based evaluation, where the model assigns probabilities to each candidate by appending it as a continuation of the prompt, and (2) an option-selection evaluation, where the model selects from labeled options (A/B/C/D). In the completion-based evaluation, we report performance using accuracy, while in the option-selection evaluation, we use exact match. To mitigate known issues, such as selection or position bias in the option-selection evaluation, we randomly shuffle the order of the options (using random seed 42). Details of the specific prompt templates and formatting can be found in Appendix O.
Models.
We evaluate two main groups of model sizes: those with 2-3 billion parameters and those with 7-8 billion parameters (team2024gemma; grattafiori2024llama; qwen2025qwen25technicalreport; jiang2023mistral7b). Each group includes both pretrained models and instruction-tuned models. To also examine larger models, we incorporate API models, which are assessed only in the option-selection setting  (achiam2023gpt; hurst2024gpt; TheC3). This is because recent APIs do not provide the log-probability outputs needed for completion-based evaluation.
  • 2-3B models:
    • Pretrained: gemma-2b, Llama-3.2-3B, Qwen2.5-3B
    • Instruction-tuned: gemma-1.1-2b-it, Llama-3.2-3B-Instruct, Qwen2.5-3B-Instruct
  • 7-8B models:
    • Pretrained: gemma-7b, Llama-3.1-8B, Mistral-7B-v0.3, Qwen2.5-7B
    • Instruction-tuned: gemma-1.1-7b-it, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen2.5-7B-Instruct
  • API models: GPT-4o mini, GPT-4.1 mini, Claude Haiku 3.5
Zero-shot and few-shot evaluation.
For each model, we evaluate performance in both zero-shot and few-shot settings using the Language Model Evaluation Harness (eval-harness). In the few-shot scenario, we use examples from the demonstration set as in-context demonstrations. Results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and are reported for one, five, and ten examples from the demonstration set (1-shot, 5-shot, and 10-shot). We present the performance results on the test set for each model and prompt configuration.
Supervised fine-tuning.
We conduct Supervised Fine-Tuning (SFT) using the LLaMA-Factory framework (zheng2024llamafactory) on the Sentence-Negation Pair dataset from Thunder-NUBench. The dataset is formatted in the Alpaca instruction style (alpaca). To achieve parameter-efficient training, we apply Low-Rank Adaptation (LoRA) (hu2022lora) with a rank of 8, targeting all linear layers. The fine-tuning process is carried out for three epochs, using a batch size of 1, a gradient accumulation step of 8, cosine learning rate scheduling, and bfloat16 precision. After the SFT, we evaluate the model’s zero-shot performance to directly assess its ability to generalize from instruction tuning without being influenced by in-context examples. It is important to note that API models are not fine-tuned, as they are not compatible with the LLaMA-Factory framework.
Figure 3: Model performance on Thunder-NUBench. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
5.2 Model Performance on Thunder-NUBench
Figure 3 displays evaluation results from all models across three settings: the zero-shot baseline, the few-shot baseline, and the zero-shot after SFT using Thunder-NUBench. It summarizes the overall trends concerning model sizes, training configurations (including whether it is instruction-tuned), and evaluation settings. Complete results are reported in Appendix P.
In the completion-based evaluation, performance steadily improves from zero-shot to few-shot and further after SFT, with gaps between smaller and larger models narrowing after SFT. In the option-selection evaluation, performance also rises with more shots, but SFT yields smaller gains than few-shot prompting. Meanwhile, instruction-tuned models remain relatively stable across conditions. API models achieve the best overall performance, though their performance doesn’t consistently improve with more shots, except at 10-shot.
5.3 Analysis of Negation Understanding Performance
We analyze model errors to evaluate the ability of our models to differentiate standard negation from similar semantic variants. Each type of local negation in our dataset is explicitly labeled based on its sentence structure: relative clause, participle clause, compound sentence, and adverbial clause, as defined in Table 3.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
completion- based Llama- 3.1-8B Baseline zeroshot 0.562 70.62 21.33 8.05 25.64 30.84 64.29 43.87
5shot 0.393 83.87 14.31 1.81 20.83 22.40 47.62 45.81
SFT zeroshot 0.203 85.55 12.50 1.95 8.97 11.69 18.71 32.26
Llama- 3.1-8B- Instruct Baseline zeroshot 0.488 70.57 26.18 3.25 24.36 25.97 54.76 37.74
5shot 0.347 80.55 18.76 0.69 18.91 19.16 39.80 37.74
SFT zeroshot 0.244 86.04 12.34 1.62 12.18 12.66 28.23 33.87
option- selection Llama- 3.1-8B Baseline zeroshot 0.541 47.95 6.74 45.31 23.08 24.03 41.50 19.03
5shot 0.362 78.95 6.58 14.47 30.45 21.75 40.82 25.16
SFT zeroshot 0.444 58.79 27.34 13.87 20.83 20.13 37.41 20.65
Llama- 3.1-8B- Instruct Baseline zeroshot 0.256 76.53 15.11 8.36 16.67 18.83 19.05 23.23
5shot 0.269 74.04 21.83 4.13 20.83 16.88 25.17 19.35
SFT zeroshot 0.254 72.32 19.03 8.65 15.06 17.86 21.09 14.52
Table 6: Error distribution and confusion analysis of pretrained and instruction-tuned Llama-3.1-8B models across various evaluation settings: zero-shot baseline, five-shot baseline, and SFT zero-shot.
To identify which subtypes of local negation are most frequently confused with standard negation, we calculate the confusion rate. This is defined as the proportion of examples within each subtype where the model incorrectly selects the local negation option instead of the correct standard negation. For example, if 320 items are labeled as participle clause negation and the model incorrectly chooses the local negation option instead of the correct standard negation option in 32 of these cases, the confusion rate for participle clause negation would be 10%. Complete analysis results are provided in Appendix R.
We focus on the results of Llama-3.1-8B and its instruction-tuned version, as shown in Table 6. In the completion-based setting, error rates generally decrease from zero-shot to 5-shot and continue to improve after SFT, with most errors concentrated in local negation options. Within local negation, compound sentences exhibit the highest confusion but also show the largest relative improvement after SFT.
In the option-selection setting, the base model demonstrates unusually high errors for paraphrase options in the zero-shot scenario (nearly 40%). Although these errors decrease after SFT, they are partly offset by increases in errors for local negation and contradiction options. The instruction-tuned model consistently achieves lower overall error states and confusion rates than the base version. Compared to the completion-based setting, the option-selection setting results in lower confusion rates for adverbial clause negation.
Overall, these patterns highlight how different evaluation settings and model configurations lead to distinct types of errors, and how the addition of more examples or SFT affects error distribution.
6 Conclusion
In this work, we introduce Thunder-NUBench, a benchmark designed to evaluate LLMs’ sentence-level understanding of negation, going beyond surface cue detection. By distinguishing between standard negation, local negation, contradiction, and paraphrase, Thunder-NUBench offers a comprehensive assessment of semantic comprehension. Our experiments demonstrate that while supervised fine-tuning and in-context learning can help reduce specific errors, these approaches still struggle to differentiate standard negation from closely related semantic variants. Thunder-NUBench serves as a valuable diagnostic tool for analyzing the limitations of models’ understanding of negation and stands as a robust benchmark for future research. Its design enables evaluation across diverse model families and settings, making it broadly applicable for studying semantic reasoning in LLMs.
Limitations
Thunder-NUBench is exclusively constructed in English, despite negation being a universal linguistic phenomenon demonstrated in diverse forms across languages. The syntactic and semantic expressions of negation may vary in other languages, meaning that our current findings may not generalize to multilingual or cross-lingual settings. In future work, we aim to extend the research to a broader range of languages to enable cross-linguistic evaluation of negation understanding in language models.
Although we built the dataset using two distinct sources (HoVer and Wikipedia summaries), both are derived from encyclopedic, formal domains, which may not fully represent the variety of sentence structures and informal language found in real-world use cases. Moreover, while all examples were systematically generated and reviewed, some bias may persist due to subjective decisions in the human curation process. We attempt to mitigate this through cross-checking by an independent group of authors, but some residual bias may remain.
Thunder-NUBench primarily focuses on standard (sentence-level) negation and its distinction from local negation, contradiction, and paraphrase. Other important negation phenomena, such as double negation and negative polarity items (NPIs), are not directly addressed in this benchmark. Our current focus is establishing a strong foundation for evaluating models’ understanding of standard negation. However, we aim to expand the evaluation to a broader range of negation phenomena in future work.
Ethical Considerations
This work does not involve the use of crowd-sourcing methods. Instead, all data included in the Thunder-NUBench benchmark was carefully reviewed by the authors to ensure quality, relevance, and adherence to ethical standards. The datasets and tools used for training and evaluation are publicly available and used in compliance with their respective licenses.
When leveraging OpenAI’s text generation models, we take additional care to avoid generating or including any content that is harmful, biased, or violates privacy. All generated examples are manually reviewed to meet ethical and safety standards. We ensure no personally identifiable information or offensive content is present in the final dataset.
The Thunder-NUBench dataset will be released under the Creative Commons Attribution Non-Commercial Share Alike 4.0, ensuring transparency, reproducibility, and accessibility for future research.
Acknowledgments
This work was partially supported by the National Research Foundation of Korea (NRF) under Grant No. RS-2023-00222663 (Center for Optimizing Hyperscale AI Models and Platforms), and by the Institute for Information and Communications Technology Promotion (IITP) under Grant No. 2018-0-00581 (CUDA Programming Environment for FPGA Clusters) and No. RS-2025-02304554 (Efficient and Scalable Framework for AI Heterogeneous Cluster Systems), all funded by the Ministry of Science and ICT (MSIT) of Korea. It was also partially supported by the Korea Health Industry Development Institute (KHIDI) under Grant No. RS-2025-25454559 (Frailty Risk Assessment and Intervention Leveraging Multimodal Intelligence for Networked Deployment in Community Care), funded by the Ministry of Health and Welfare (MOHW) of Korea. Additional support was provided by the BK21 Plus Program for Innovative Data Science Talent Education (Department of Data Science, Seoul National University, No. 5199990914569) and the BK21 FOUR Program for Intelligent Computing (Department of Computer Science and Engineering, Seoul National University, No. 4199990214639), both funded by the Ministry of Education (MOE) of Korea. This work was also partially supported by the Artificial Intelligence Industrial Convergence Cluster Development Project, funded by the MSIT and Gwangju Metropolitan City. Research facilities were provided by the Institute of Computer Technology (ICT) at Seoul National University.
Appendix A Typology of Contradiction
Contradictions in natural language can arise in diverse ways that go beyond simple negation. Following the typology of de2008finding, contradictions can be grouped into seven categories: antonymy, explicit negation, numeric mismatch, factive/modal inconsistencies, structural reversals, lexical incompatibilities, and conflicts based on world knowledge. These categories reflect the fact that contradiction covers a broader semantic scope than negation alone. Table 7 summarizes these types with definitions and examples.
Contradiction Type Definition Example
Antonym
Contradiction caused by opposing
meanings of aligned words.
The policy was a success.
$\rightarrow$ The policy was a failure.
Negation
One sentence explicitly negates
a statement in the other.
She attended the meeting.
$\rightarrow$ She did not attend the meeting.
Numeric
Inconsistent numbers, dates, or
quantities in related statements.
Totally, ten people were injured.
$\rightarrow$ Totally, five people were injured.
Factive/Modal
Conflict in implied facts or modal
possibilities due to verbs or auxiliaries.
He managed to enter the building.
$\rightarrow$ He did not enter the building.
Structure
Syntactic rearrangement or argument
swapping causes contradiction.
Alice hired Bob.
$\rightarrow$ Bob hired Alice.
Lexical
Contradiction through incompatible
verbs or phrases, not strictly antonyms.
The manager praised her performance.
$\rightarrow$ The manager expressed
     disappointment in her performance.
World
Knowledge
Contradiction relies on common-sense
or background knowledge.
The Eiffel Tower is in Paris.
$\rightarrow$ The Eiffel Tower is in Berlin.
Table 7: Contradiction types from de2008finding. Contradiction covers a broader scope than negation.
Appendix B Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (hengeveld1986copular).
As discussed in Section 3.3, standard negation in this work targets the main predicate of a clause. For sentences with copular verbs, this means that the entire verb phrase, including the copular verb and its complement, is subject to negation. For example, in the sentence "She is a doctor," the main predicate is "is a doctor." Negating this sentence results in "She is not a doctor," where the negation applies to the entire predicate, not just the verb "is."
Negation of a verb phrase including a copular verb can be realized either syntactically (e.g., "is not an expert") or by replacing the complement with its complementary antonym (e.g., "is a non-expert"), both of which result in the reversal of the main predicate’s truth value. Although such constructions may superficially appear to be non-verbal negation, especially when the complement is a noun or adjective, they are, in fact, instances of verbal negation, since the negation applies to the predicate as a whole.
Appendix C Negation of Implications
Negating implications presents challenges, as natural language intuitions often diverge from the rules of formal logic. Let’s say there is a conditional statement, "If I study hard, I will pass the bar exam." Formally, let $P$ denote "I study hard" and $Q$ denote "I will pass the bar exam." In classical logic, the conditional "if $P$, QQ" can be false only when $P$ is true and $Q$ is false. This implies that the negation of the conditional is $P\;\text{and}\;\text{Neg}(Q)$" ("I study hard and I won’t pass the bar exam,) while the conditional itself is equivalent to $\text{Neg}(P)\;\text{or}\;Q$" ("I don’t study hard or I will pass the exam) (nguyen2023negation).
Psychological studies confirm that people often accept both "if $P$, Neg(QQ)" ("If I study hard, I won’t pass the exam.") and "if Neg($P$), QQ" ("If I don’t study hard, I will pass the exam."). However, the former can be interpreted as $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$", and the latter $P\;\text{or}\;Q$", both of which are not equivalent to the original statement’s negation, $P\;\text{and}\;\text{Neg}(Q)$". "if Neg($P$), Neg(QQ)" ("If I don’t study hard, I won’t pass the exam.") is not the correct negation as well, as it is equivalent to $P\;\text{or}\;\text{Neg}(Q)$(espino2012not).
While humans often struggle to distinguish the correct negation of a conditional from invalid alternatives, the logical form is unambiguous. We therefore include conditional statements in our benchmark to test whether language models, like humans, are prone to intuitive but invalid interpretations, or whether they can correctly apply truth-functional reasoning.
Note that implications in natural language are not limited to the explicit "if $P$, QQ" form, but may also appear with connectives such as when, as long as, or unless, which functionally convey conditional meaning and are treated under the same negation principle.
Appendix D Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (gleitman1965coordinating).
Coordinating conjunctions connect elements of equal grammatical rank. The seven common ones in English are: for, and, nor, but, or, yet, so (often remembered as FANBOYS). Among these, and, or, and but are indisputably used to coordinate clauses. The others can be ambiguous or function in non-coordinating roles(e.g., indicating cause or result rather than logical structure). These are the examples using and, or, and but to connect sentences equally.
  • "She studied hard, and she passed the exam."
  • "I wanted to go, but it was raining."
  • "You can call me, or you can send an email."
We consider only the coordinating conjunctions and, or, and but as indicators of compound sentences, in which two or more independent clauses are equally connected. Although but introduces a contrast semantically, in terms of logical structure, it functions as a conjunction equivalent to and; therefore, its negation follows the same principle.
Appendix E Local Negation Constructions Excluded
In constructing the Thunder-NUBench dataset, we consider various types of local (i.e., subclausal) negation, where negation applies to a phrase or constituent rather than the main predicate. However, several constructions are excluded due to their semantic ambiguity, syntactic irregularity, or misalignment with the benchmark’s focus on verbal negation.
Infinitive Phrase Negation.
Infinitive phrases (e.g., "to go") can be negated with "not" (e.g., "not to go" or "to not go"). Unlike the clause-level structures that define our local negation category, infinitive phrases are not full clauses but simply part of a verb phrase, making them less compatible with our definition. Moreover, although grammatically correct, this construction is relatively rare and sounds awkward depending on the context.
  • Original: George wants to go to the park.
  • Negated (infinitive): George wants not to go / George wants to not go to the park.
For these reasons, we exclude infinitive phrase negation from the benchmark.
Appositive Clause Negation.
Appositive clauses are noun phrases that provide descriptive clarification. Attempting to negate an appositive typically involves lexical replacement rather than syntactic negation.
  • Original: My brother, a talented musician, plays the guitar.
  • Negated (appositive): My brother, not a talented musician, plays the guitar.
Such changes alter descriptive content rather than reversing the meaning of the predicate, and often fall into the domain of contradiction. Accordingly, they are excluded from the dataset.
Prepositional Phrase Negation.
Negating a prepositional phrase often involves replacing the preposition with its antonym (e.g., "with" → "without", "in" → "outside"), which results in a sentence that differs in content, rather than reversing the meaning of the predicate.
  • Original: She went to the park with her bird.
  • Negated (preposition): She went to the park without her bird.
Since such modifications do not negate the verb but instead change the nature of an adjunct or argument, they fall outside the scope of standard negation or local negation in this work and are excluded.
In all of the above cases, the negation does not target the whole verb phrase but rather peripheral elements within the sentence. As the Thunder-NUBench is designed to evaluate verbal negation, these local or phrase-level forms of negation were intentionally left out.
Appendix F Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (deprez2015double).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(van1996litotes). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we have generated paraphrase candidates for our dataset using such double negation patterns. For example,
  • Sentence: His characteristic style fuses samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    Double Negation: His characteristic style does not fail to fuse samba, funk, rock, and bossa nova with lyrics that blend humor and satire with often esoteric topics.
  • Sentence: It covers a broad range of fields, including the humanities, social sciences, exact sciences, applied sciences, and life sciences.
    Double Negation: It does not exclude a broad range of fields, including the humanities, social sciences, exact sciences, applied sciences, and life sciences.
  • Sentence: Sanders was honoured to meet with many world dignitaries and representatives of UNESCO member nations, and delighted when delegates from UNESCO, visited Toowoomba in 2018 in return.
    Double Negation: Sanders was not unhappy to meet with many world dignitaries and representatives of UNESCO member nations, and not displeased when delegates from UNESCO visited Toowoomba in 2018 in return.
However, upon closer examination, these paraphrase candidates do not always preserve the exact meaning of the original sentence. The antonyms used (e.g., "exclude" for "cover," "unhappy" for "honoured") are not always true complementary antonyms, which does not effectively negate the meaning. Moreover, the litotes construction ("does not fail to fuse") tends to add an emphatic nuance, rather than being a perfect semantic equivalent. Therefore, the boundary between paraphrasing and double negation is ambiguous, and their relationship requires more careful analysis. Given these issues, and because our primary focus is on standard negation, we ultimately decide to exclude double negation constructions as paraphrase candidates from our dataset.
Appendix G HoVer Dataset
Column Detail Example
id Unique claim identifier 0
uid
User/annotator identifier
330ca632-e83f-4011-b11b-0d0158145036
claim
The statement to be verified, often requiring multi-article evidence
Skagen Painter Peder Severin Krøyer favored
naturalism along with Theodor Esbern Philipsen
and the artist Ossian Elgström studied
with in the early 1900s.
supporting_facts
List of Wikipedia article titles and sentence indices providing evidence
[ { "key": "Kristian Zahrtmann", "value": 0 },
{ "key": "Kristian Zahrtmann", "value": 1 },
{ "key": "Peder Severin Krøyer", "value": 1 },
{ "key": "Ossian Elgström", "value": 2 } ]
label
Whether the claim is supported
1: SUPPORTED or 0: NOT_SUPPORTED
num_hops
Number of articles required for verification
[…]
hpqa_id
Reference to the original HotpotQA pair
5ab7a86d5542995dae37e986
Table 8: Details of HoVer dataset structure with examples.
The HoVer (Hoppy Verification) dataset is developed for the tasks of multi-hop evidence retrieval and factual claim verification. In HoVer, each claim requires supporting evidence that spans multiple English Wikipedia articles to determine whether the claim is substantiated or not. The dataset is distributed under a CC BY-SA 4.0 License, and it can be accessed via its official homepage [^1] https://hover-nlp.github.io/. Table 8 offers an overview of the dataset’s structure. The data is split into training, validation, and test sets, containing 18,171, 4,000, and 4,000 examples respectively.
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (yang2018hotpotqa). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (jiang2020hover).
Appendix H Wikipedia Summary Dataset
Column Detail Example
title Article title from Wikipedia. Alain Connes
description A brief description or category for the article (when available). French mathematician
summary
The extracted summary or introduction section of the article,
typically more concise than the full text.
Alain Connes (; born 1 April 1947) is a French mathematician…
full_text
The complete article text (when included), encompassing the
full body of the Wikipedia page.
Alain Connes (; born 1 April 1947) is a French mathematician…
__index_level_0__ Index number for each entry in the dataset. 3
Table 9: Details of Wikipedia Summary dataset structure with examples.
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^2] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^3] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^4] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 9 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (scheepers2017compositionality).
Appendix I Human Review Protocol
To ensure high-quality data construction, we implement a rigorous quality control protocol that combines generation, independent review, and iterative consensus building. The process involves the following key steps:
  • Task allocation and independence. Authors are assigned distinct portions of the dataset, but no author is permitted to review the data they have generated. This ensures that each instance is subject to at least one independent review.
  • Sequential authoring across choices. For the multiple-choice dataset, construction proceeds in four stages: standard negation, local negation, contradiction, and paraphrase. At each stage, different authors are responsible for creating the new option, while reviewers who have not authored that option perform the verification.
  • Cross-checking and layered review. Each newly created option is reviewed by at least one other author, and reviewers also revisit earlier options in the same instance. For example, when reviewing the paraphrased sentence, the reviewer also checks that standard negation, local negation, and contradiction sentences are correct. As a result, every instance undergoes multiple rounds of verification across stages, such that all authors ultimately examine data they have not created themselves.
  • Guideline refinement and retroactive correction. Generation and reviewing guidelines are continuously updated based on discussion of ambiguous or problematic cases. Whenever the guidelines changes, all previously created data are revisited to ensure compliance, promoting consistency across the dataset.
  • Consensus and adjudication. Disagreements are discussed in weekly meetings and, if necessary, adjudicated by a lead reviewer, ensuring that no instance remains unresolved.
Overall, this iterative and layered procedure ensures that every instance in the multiple-choice dataset is independently reviewed multiple times, leading to stable guidelines and a consistent dataset.
Appendix J Detailed Principles and Examples of the Thunder-NUBench
While the main text already defines the core notions of standard and local negation (Section 3) and explains how they are applied throughout dataset construction (Section 4), here we provide more detailed illustrations.
Paraphrasing before Negation.
Before negating, the main verb or other components may be paraphrased with synonyms, provided that the sentence’s tense, structure, and meaning remain strictly equivalent before applying standard negation. Authors refer to the Merriam-Webster Thesaurus  [^5] https://www.merriam-webster.com/. For example,
  • Original Sentence: Toumour is a village and rural commune in Niger located near the Niger–Nigeria border.
    • Paraphrased Sentence: Toumour is a village and rural commune in Niger that is found close to the Niger–Nigeria boundary.
      $\rightarrow$ Standard Negation after Paraphrase: Toumour isn’t a village and rural commune in Niger that is found close to the Niger–Nigeria boundary.
    • Explanation: In this example, the participle clause "located near the Niger–Nigeria border" is rephrased as a relative clause "that is found close to the Niger–Nigeria boundary." Since both constructions serve as modifiers and preserve the same semantic role, we treat them as equivalent in meaning for the purpose of standard negation.
  • Original Sentence: The armed forces said Boko Haram attacked their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
    • Paraphrased Sentence: The armed forces stated that Boko Haram assaulted their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
      $\rightarrow$ Standard Negation after Paraphrase: The armed forces didn’t state that Boko Haram assaulted their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
    • Explanation: In this example, the reporting verb “said” is paraphrased as “stated,” and the verb “attacked” is replaced with the synonym “assaulted.” These substitutions preserve the original tense and meaning, allowing standard negation to be applied without altering the semantic content of the sentence.
Negation of Simple Sentences.
For simple, declarative sentences, standard negation is achieved by inserting "not" after the auxiliary or main verb, or by replacing the predicate with its complementary antonym. For example, "She is happy." $\rightarrow$ "She is not happy."; "The room is occupied." $\rightarrow$ "The room is unoccupied."
Negation in Compound Sentences.
When multiple clauses or propositions are coordinated (e.g., with "and", "or", "but"), standard negation is logically applied, governed by De Morgan’s laws. Here, "but" is treated as a coordinating conjunction equivalent to "and" in terms of logical structure, so its negation follows the same principle.
  • Conjunction "$P$ and/but QQ": the negation is $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$".
  • Disjunction "$P$ or QQ": the negation is $\text{Neg}(P)\;\text{and}\;\text{Neg}(Q)$".
For example, "He passed the test and received an award." is negated as "He did not pass the test or did not receive an award."
When application of logical negation produces unnatural language, sentences may be split or slightly rephrased for fluency, provided logical meaning is preserved. For example,
  • Original: "He finished the report and submitted the assignment."
  • Standard Negation: "He did not finish the report or did not submit the assignment."
  • Standard Negation, but Splitted: "He did not finish the report. Or, he did not submit the assignment."
Coordinated Elements in the Sentence.
When a sentence contains coordinated elements (such as subjects, objects, or predicates connected by "and" or "or"), standard negation typically follows logical principles derived from De Morgan’s Laws. However, whether logical negation applies to each individual component or to the entire predicate as a whole depends on whether the coordination expresses multiple independent propositions or a single collective event.
  • If the coordination introduces semantically distinct propositions, that is, each conjunct could form a complete sentence on its own, negation must be applied to each proposition individually. For example, "My sister and I studied hard."
    This sentence can be interpreted as: "My sister studied hard and I studied hard."
    Therefore, the correct standard negation is: "My sister did not study hard, or I did not study hard."
  • Conversely, if the coordination connects elements that jointly participate in a single action or state (e.g., a shared subject or a collective predicate), then the sentence is treated as a simple clause, and the predicate as a whole is negated. Logical decomposition is not appropriate. For example, "My sister and I share clothes."
    This expresses a single collective action involving both participants.
    Therefore, the correct standard negation is: "My sister and I do not share clothes."
    (NOT: "My sister does not share clothes, or I do not share clothes.")
  • This distinction is crucial: even if two noun phrases are coordinated, if the sentence semantically decomposes into separate atomic propositions, standard negation must apply to each atomic proposition. Otherwise, it applies to the whole predicate as one unit.
  • Other examples of semantically collective predicates where logical splitting is not appropriate include: "be the same", "have in common", "do something together", "combine", "unite", etc. These describe inherently joint or relational properties, not independent propositions. For example, "Clarence Brown and Peter Glenville are from the same country." should be negated as "Clarence Brown and Peter Glenville are not from the same country."
Use of Antonyms.
When replacing predicates with antonyms in standard negation, only complementary antonyms are appropriate, as they provide a clear binary opposition, ensuring logical consistency of negation. Gradable and relational antonyms are unsuitable for standard negation because their antonyms do not represent the logical complement of the original predicate. In other words, replacing a predicate $p$ with its antonym does not produce $\neg p$ in a truth-conditional sense.
Specifically, unlike complementary antonyms, which form mutually exclusive pairs (i.e., $p\cup\neg p=U$ and $p\cap\neg p=\emptyset$), gradable and relational antonyms do not partition the meaning space cleanly, and thus fail to reverse the truth value reliably.
  • Complementary Antonyms: Also called binary/contradictory antonyms. These antonyms represent mutually exclusive pairs with no intermediate states. The presence of one implies the absence of the other. Examples include:
    • alive / dead
    • true / false
    • present / absent
    • occupied / vacant
    Using complementary antonyms in negation ensures a direct and unambiguous reversal of the original proposition’s truth value.
  • Gradable Antonyms: These antonyms exist on a continuum and allow for varying degrees between the two extremes. Negating one does not necessarily affirm the other. Examples include:
    • hot / cold
    • happy / sad
    • tall / short
    • young / old
    Due to their scalar nature, gradable antonyms are inappropriate for standard negation, as they do not provide a definitive binary opposition.
  • Relational Antonyms: Also known as converse antonyms, these pairs describe a reciprocal relationship where one implies the existence of the other. Examples include:
    • parent / child
    • teacher / student
    • buy / sell
    • employer / employee
    Relational antonyms are context-dependent and do not represent direct opposites in a binary sense, making them unsuitable for standard negation purposes.
General Principles of Standard Negation.
  • The negated sentence must preserve all elements (subject, tense, objects, adjuncts, etc.) of the original, except for the truth value of the main predicate.
  • When naturalness and logical negation conflict, logical correctness takes priority, but minimal rephrasing is allowed for fluency.
  • If the negated clause creates a contradiction with other parts of the sentence, the contradictory clause must be removed. For example, the standard negation of the sentence "While the spatial size of the entire universe is unknown, it is possible to measure the size of the observable universe, which is approximately 93 billion light-years in diameter."
    will be "While the spatial size of the entire universe is unknown, it isn’t possible to measure the size of the observable universe."
    The relative clause must be removed because its content directly contradicts the negated main clause.
Common Negation Errors and Corrections.
  • Original sentence: His characteristic style fuses samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Incorrect negation: His distinctive style doesn’t fuse samba, funk, rock or bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Correct negation: His distinctive style doesn’t fuse samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Explanation: The verb "fuse" implies a combination of all listed elements. "and" must be preserved.
  • Original sentence: The mascot of Avon Center School is the "Koalaty Kid," while the mascot at Prairieview is an eagle and the mascot at Woodview is an owl.
    • Incorrect negation: Avon Center School’s mascot is not the "Koalaty Kid," Prairieview’s mascot is not an eagle, or Woodview’s mascot is not an owl.
    • Correct negation: Avon Center School’s mascot is not the "Koalaty Kid," while the mascot at Prairieview is an eagle and the mascot at Woodview is an owl.
    • Explanation: Two clauses connected by while are not coordinated propositions (as with and or or), but instead express contrastive information. Therefore, applying logical negation across both clauses is incorrect. Negation should apply only to the main clause (here, the first statement), while the contrasting clause remains affirmative.
Appendix K Code for Data Construction
K.1 Sentence-Negation Pair Dataset
To construct the sentence-negation pair dataset, we begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective.
We select different model versions depending on the complexity of each task. For sentence merging, which demands nuanced contextual understanding and complex syntax, we use GPT-4. For grammar correction, where edits are more straightforward, GPT-3.5 is sufficient.
1def grammar_fix(claim):
2 messages = [{"role": "system", "content": "Fix grammatical errors."},
[…]
4 response = client.chat.completions.create(model="gpt-3.5-turbo", messages=messages)
5 fixed_text = response.choices[0].message.content.strip()
6 return fixed_text
Listing 1: Fixing Grammar with OpenAI API.
1def merge_sentences_with_gpt(claim):
2 messages = [{"role": "system", "content": "Merge sentences into a single one."},
[…]
4 response = client.chat.completions.create(model="gpt-4-turbo-preview", messages=messages)
5 merged_text = response.choices[0].message.content.strip()
6 return merged_text
Listing 2: Merging Sentences with OpenAI API.
K.2 Multiple Choice Dataset
To construct the multiple-choice dataset, we first segment the "summary" column of the Wikipedia Summary dataset, which often contains multiple sentences in a single entry, into individual sentences. To focus on the challenges of negation in complex sentences, we filter out sentences that are too short. This process is done with Python code.
Since conditional sentences (e.g., "If $P$, QQ") are rarely present in the Wikipedia summary dataset, we adopt a two-step approach: (1) prompting the model to generate conditional variants from given sentences (using OpenAI API, GPT-4o-mini), and (2) manually filtering or lightly editing the results to obtain valid conditionals.
Subsequently, we automatically generate contradictions and paraphrases for each sentence via the OpenAI API (GPT-4o) as well, followed by human review. The following scripts illustrate the procedures.
1import pandas as pd
2import re
3from datasets import load_dataset
4import random
5
6df = pd.DataFrame(load_dataset("jordiclive/wikipedia-summary-dataset")[’train’].shuffle(seed=42).select(range(10000))
7df = df.drop(columns=[’full_text’])
8
9def split_into_sentences(text):
10 sentences = re.split(r’(?<=[.!?]) +’, text)
11 return sentences
12
13df[’sentence’] = df[’summary’].apply(split_into_sentences)
14df = df.explode(’sentence’)
15df = $df[$’sentence’].apply(lambda x: len(x.split()) >= 30)]
16df = df.reset_index(drop=True)
17df.to_csv("file/wikipedia_summary_sentences.csv", index=False)
Listing 3: Sentence extraction and preprocessing from Wikipedia summaries.
1def generate_conditionals(sentence):
2 prompt = f"""
3 Based on the sentence below, write a conditional sentence that uses the main topic of the sentence.
4 The conditional sentence should express a hypothetical situation or cause-effect relationship related to the topic. It can be slightly complex in structure.
5 For example:
6 - If it rains tomorrow, I will stay home.
7 Sentence:
8 ’{sentence}’
9 """
10
11 completion = client.chat.completions.create(
12 model="gpt-4o-mini",
13 messages=[
14 {"role": "system", "content": "You are a helpful assistant that specializes in generating conditional sentences."},
15 {"role": "user", "content": prompt}
16 ]
17 )
18
19 return completion.choices[0].message.content
Listing 4: Conditionals sentence generation.
1def generate_contradiction(sentence):
2 prompt = f"""
3 You will be given a sentence. Generate a contradictory sentence that directly conflicts with the original sentence without using standard negation.
4
5 Definitions:
6 - Standard negation: Directly negating the main verb or using words like ’not’, ’no’, ’never’, or negative contractions such as \"isn’t\", \"doesn’t\", or \"can’t\".
7 - Contradiction: A sentence that logically conflicts with the original statement. The contradiction must be such that both sentences cannot logically be true at the same time under any circumstances.
8
9 Important:
10 - Do not change the main verb from the original sentence.
11 - Do not use ’never’ or other negative words to form the contradiction.
12 - Ensure the contradicted sentence logically excludes the possibility of the original sentence being true simultaneously.
13
14 Examples:
15 Original sentence: \"The tallest student won the award.\"
16 Contradicted sentence: \"The shortest student won the award.\"
17
18 Original sentence: \"The room was completely dark.\"
19 Contradicted sentence: \"The room was brightly lit.\"
20
21 Original sentence: \"The event took place in the morning.\"
22 Contradicted sentence: \"The event took place in the evening.\"
23
24 Original sentence: \"All people are dying.\"
25 Contradicted sentence: \"Some people are dying.\"
26
27 Now, generate a contradictory sentence without standard negation, without changing the main verb, and ensuring the two sentences are logically incompatible, for the following:
28
29 Original sentence: \"{sentence}\"
30
31 Contradicted sentence:
32 """
33
34 completion = client.chat.completions.create(
35 model="gpt-4o",
36 messages=[
37 {"role": "system", "content": "You are a helpful assistant tasked with generating logical contradictions. Do not use negation to make contradiction."},
38 {"role": "user", "content": prompt}
39 ]
40 )
41
42 return completion.choices[0].message.content
Listing 5: Contradiction generation.
1def generate_paraphrase(sentence):
2 prompt = f"""
3 Paraphrase the following sentence using synonyms or slight structural variations without changing its meaning.
4 Do not add or remove any main verbs. Keep the original intent of the sentence intact.
5
6 Original sentence: "{sentence}"
7
8 Paraphrased sentence:
9 """
10
11 completion = client.chat.completions.create(
12 model="gpt-4o",
13 messages=[
14 {"role": "system", "content": "You are a helpful assistant skilled at generating paraphrases while keeping the meaning of sentences unchanged."},
15 {
16 "role": "user",
17 "content": prompt
18 }
19 ]
20 )
21
22 return completion.choices[0].message.content
Listing 6: Paraphrase generation.
Appendix L Thunder-NUBench Dataset Structure
Thunder-NUBench consists of two subsets: a sentence-negation pair dataset for supervised fine-tuning and a multiple-choice dataset for evaluation. Both datasets are built on English text and reviewed by authors following strict guidelines.
L.1 Sentence-Negation Pair Dataset
This subset contains pairs of affirmative and corresponding standard negation sentences. It includes the following fields:
  • index: the index of the data.
  • premise: the original sentence.
  • hypothesis: its logically negated form.
L.2 Multiple-Choice Dataset
This evaluation set presents each original sentence with four candidate transformations.
  • wikipedia_index: the original index of the Wikipedia Summary dataset.
  • index: the index of the data.
  • sentence: the original sentence.
  • choice1: standard negation (correct answer).
  • choice2: local negation (subclausal negation).
  • choice2_type: specifies the type of local negation.
  • choice2_element: a short description of the phrase or clause that was negated (e.g., "being built", "which crashed").
  • choice3: contradiction (non-negated, semantically incompatible).
  • choice4: paraphrase (semantically equivalent).
choice2_type Definition
Demonstration
Set
Test Set
relative_part
negation inside relative clauses
(e.g., "who did not attend…").
12 312
pp_part
negation in participle clauses
(e.g., "not walking through the park…").
12 308
adverb_part
negation in adverbial clauses
(e.g., "because it was not raining").
12 310
compound_part negation applied to one clause within a compound sentence. 12 294
non-applicable
used when the sentence structure does not support
a valid local negation variant under our definition.
2 37
Total 50 1,261
Table 10: Choice 2 Types and Distributions.
The details of choice2_type and distribution on demonstration and test sets are described in Table 10. It follows the definition in Table 3.
In addition to the Wikipedia Summary dataset, we supplement the evaluation set with conditionals (e.g., If $P$, $Q$) by manually searching Wikipedia articles where such constructions are more likely to occur (e.g., Newton’s laws of motion). Among the 100 conditional sentences included across demonstration and test sets, 20 are collected through manual search (marked with indices beginning with "S" in wikipedia_index). Meanwhile, the remaining 80 are sampled from the Wikipedia Summary dataset and converted into conditional form using the script in 4 (Appendix K).
Appendix M Incorporating Thunder-NUBench into LM Evaluation Harness
This section describes how Thunder-NUBench is integrated into the LM Evaluation Harness (eval-harness) for zero-shot and few-shot evaluation, both in completion-based and option-selection settings.
1task: nubench_completion
2dataset_path: {dataset_path}/NUBench
3dataset_name: multiple-choice
4output_type: multiple_choice
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_completion
8doc_to_text: "{{query}}"
9doc_to_target: "{{gold}}"
10doc_to_choice: "choices"
11metric_list:
12 - metric: acc
13 aggregation: mean
14 higher_is_better: true
Listing 7: NUBench/NUBench_completion.yaml
1task: nubench_option
2dataset_path: {dataset_path}/NUBench
3dataset_name: multiple-choice
4output_type: generate_until
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_option
8doc_to_text: "{{query}}"
9doc_to_target: "{{answerLetter}}"
10generation_kwargs:
11 until:
12 - "</s>"
[…]
14metric_list:
15 - metric: exact_match
16 aggregation: mean
17 higher_is_better: true
18 ignore_punctuation: true
19 ignore_case: true
20filter_list:
21 - name: get_response
22 filter:
23 - function: "regex"
[…]
25 - function: remove_whitespace
26 - function: "regex"
[…]
28 - function: take_first
29dataset_kwargs:
30 trust_remote_code: true
Listing 8: NUBench/NUBench_option.yaml
1import re
2import datasets
3import random
4
5def process_docs_completion(dataset: datasets.Dataset) -> datasets.Dataset:
6 def _process_doc(doc):
[…]
8
9 if doc.get("choice2_type", "") == "non-applicable":
10 choices = [doc["choice1"], doc["choice3"], doc["choice4"]]
11 else:
12 choices = [doc["choice1"], doc["choice2"], doc["choice3"], doc["choice4"]]
13
14 return {
15 "query": prompt,
16 "choices": choices,
17 "gold": 0
18 }
19 return dataset.map(_process_doc)
20
21
22def process_docs_option(dataset: datasets.Dataset, seed: int = 42) -> datasets.Dataset:
23 rng = random.Random(seed)
24
25 def _process_doc(doc):
[…]
27
28 c1 = doc.get("choice1", "")
29 c2 = doc.get("choice2", "")
30 c3 = doc.get("choice3", "")
31 c4 = doc.get("choice4", "")
32 c2_na = str(doc.get("choice2_type", "")).lower() == "non-applicable"
33
34 gathered = []
35 keys = []
36
37 if c1:
38 gathered.append(c1); keys.append("choice1")
39 if c2 and not c2_na:
40 gathered.append(c2); keys.append("choice2")
41 if c3:
42 gathered.append(c3); keys.append("choice3")
43 if c4:
44 gathered.append(c4); keys.append("choice4")
45
46 assert len(gathered) >= 3, "Need at least 3 choices for multiple-choice."
47
48 paired = list(zip(gathered, keys))
49 rng.shuffle(paired)
50 choices, choice_keys = zip(*paired)
51
52 try:
53 answer_idx = choice_keys.index("choice1")
54 except ValueError:
55 answer_idx = 0
56
57 label_map = ["A", "B", "C", "D"]
58 labels = label_map[:len(choices)]
59
60 lines = [
61 "Given the following instruction and candidate answers, choose the single best answer.",
62 f"Instruction: {initial_prompt}",
63 ]
64 for lab, ch in zip(labels, choices):
65 lines.append(f"{lab}. {ch}")
66 joined = ", ".join(labels)
67 lines += ["", f"Your response should be one of {joined}.", "Only output the letter.", "Answer:"]
[…]
69
70 return {
71 "query": prompt,
72 "choices": list(choices),
73 "choice_keys": list(choice_keys),
74 "answerKey": answer_idx,
75 "answerLetter": labels[answer_idx],
76 }
77
78 return dataset.map(_process_doc)
Listing 9: NUBench/utils.py
Appendix N Finetuning via LLaMA-Factory
We detail our supervised fine-tuning setup using LLaMA-Factory (zheng2024llamafactory) with LoRA (hu2022lora) on Thunder-NUBench training data, including configuration of the fine-tuning and instruction-based examples in Alpaca format (alpaca).
The YAML configuration provided in Listing 10 is specific to the LLaMA-3.1-8B model. Other models (e.g., Qwen or Mistral) can be fine-tuned similarly by modifying the model_name_or_path and template fields in the configuration file accordingly.
1### model
2model_name_or_path: Llama-3.1-8B
3trust_remote_code: true
4
5### method
6stage: sft
7do_train: true
8finetuning_type: lora
9lora_rank: 8
10lora_target: all
11
12### dataset
13dataset: nubench_train
14template: llama3
15cutoff_len: 512
16max_samples: 5000
17overwrite_cache: true
18preprocessing_num_workers: 16
19dataloader_num_workers: 4
20
21### output
22output_dir: lora/sft/Llama-3.1-8B
23logging_steps: 10
24save_steps: 500
25plot_loss: true
26overwrite_output_dir: true
27save_only_model: false
28report_to: none # choices: [none, wandb, tensorboard, swanlab, mlflow]
29
30### train
31per_device_train_batch_size: 1
32gradient_accumulation_steps: 8
33learning_rate: 1.0e-4
34num_train_epochs: 3.0
35lr_scheduler_type: cosine
36warmup_ratio: 0.1
37bf16: true
38ddp_timeout: 180000000
39resume_from_checkpoint: null
Listing 10: Llama-3.1-8B_lora_sft.yaml
1[
2 {
3 "instruction": "Negate the sentence.",
4 "input": "Sentence: Eddie Vedder was born before Nam Woo-hyun.",
5 "output": "Eddie Vedder wasnt born before Nam Woo-hyun."
6 },
7 {
8 "instruction": "Negate the sentence.",
9 "input": "Sentence: Halestorm is from Pennsylvania, while Say Anything is from California.",
10 "output": "Halestorm is not from Pennsylvania, while Say Anything is from California."
11 },
12 (...)
13]
Listing 11: Sentence-Negation Pair dataset for training in alpaca format.
Appendix O Prompt Selection for In-Context Learning
We explore a range of prompt types for in-context learning, from minimal instructions to more detailed variants (zhao2021calibrate; li2023practical; wan2023better). Specifically, we evaluate three prompt styles:
  1. 1.
    Simple prompt: a minimal instruction, “Negate the sentence.”
  2. 2.
    Definition prompt: a concise description of the task, “Negate the main predicate of the main clause so that the proposition is logically reversed.”
  3. 3.
    Detail prompt: an extended instruction with explicit step-by-step guidelines on identifying the main predicate, preserving other sentence elements, handling antonyms, and applying negation consistently across logical operators (see Listing 12 for the full format).
"Reverse the truth value of the main predicate (verbal phrase) in the main clause, while preserving all other elements of the main clause unchanged.
1) Identify the main clause and its main verb (main predicate). Ignore subordinate clauses.
2) Preserve all other main-clause content.
3) Insert a negative particle such as "not" into the main verb, or replace it with a complementary antonym only if it forms an absolute binary (e.g., alive/dead, true/false, possible/impossible).
4) If the sentence contains multiple propositions connected by logical operators (e.g., and, or, conditional constructions), negate it in a way that reverses the entire proposition (e.g., A and B => not A or not B; If A then B => A and not B).
Sentence: {doc[’sentence’]}"
Listing 12: Detail prompt format.
Across both completion-based and option-selection evaluation settings, we found that the simple prompt consistently achieved the highest average performance in zero-shot and few-shot settings (See results in Appendix P.) While the definition-based and detail prompts occasionally provided more explicit guidance, they did not improve performance overall. Therefore, we adopt the simple prompt as our default setting in the main experiments. The complete prompt formats, including prompt–response structures for each evaluation setting, are provided in Listings 13 and 14, based on the simple prompt.
Prompt] "Negate the sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
Negation:"
Response] "Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells."
Listing 13: Completion-based Format.
Prompt] "Given the following instruction and candidate answers, choose the single best answer.
Instruction: Negate the sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
A. Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells.
B. Chromosome 2 is the second-largest human chromosome, which doesnt span more than 242 million base pairs or represent nearly eight percent of the whole DNA in human cells.
C. Chromosome 2 is the smallest human chromosome, spanning fewer than 50 million base pairs and representing less than two percent of the total DNA in human cells.\
D. Chromosome 2 is the second-biggest human chromosome, with over 242 million base pairs, making up nearly 8% of all DNA in human cells.
Your response should be one of A, B, C, D.
Only output the letter.
Answer:"
Response] "A"
Listing 14: Option-selection Format.
Appendix P Total Model Performance on Thunder-NUBench
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.437 0.197
0.437
(±0.005)
0.248
(±0.005)
0.524
(±0.005)
0.276
(±0.007)
0.572
(±0.004)
0.267
(±0.011)
0.732 0.259
gemma-1.1-
2b-it
0.388 0.264
0.428
(±0.005)
0.347
(±0.004)
0.526
(±0.009)
0.357
(±0.008)
0.540
(±0.008)
0.336
(±0.004)
0.750 0.007
gemma-7b 0.454 0.388
0.465
(±0.001)
0.631
(±0.012)
0.561
(±0.002)
0.691
(±0.010)
0.620
(±0.010)
0.720
(±0.006)
0.797 0.722
gemma-1.1-
7b-it
0.663 0.645
0.689
(±0.009)
0.684
(±0.006)
0.730
(±0.007)
0.714
(±0.003)
0.759
(±0.007)
0.707
(±0.004)
0.786 0.733
Llama-3.2-3B 0.444 0.313
0.481
(±0.006)
0.417
(±0.011)
0.574
(±0.006)
0.463
(±0.009)
0.601
(±0.007)
0.501
(±0.009)
0.733 0.342
Llama-3.2-3B-
Instruct
0.472 0.530
0.487
(±0.009)
0.540
(±0.011)
0.557
(±0.006)
0.526
(±0.006)
0.587
(±0.004)
0.532
(±0.005)
0.745 0.604
Llama-3.1-8B 0.439 0.459
0.493
(±0.007)
0.549
(±0.012)
0.596
(±0.006)
0.635
(±0.008)
0.635
(±0.006)
0.667
(±0.011)
0.797 0.556
Llama-3.1-8B-
Instruct
0.512 0.744
0.571
(±0.007)
0.708
(±0.007)
0.646
(±0.007)
0.726
(±0.008)
0.675
(±0.004)
0.729
(±0.011)
0.756 0.747
Mistral-7B-v0.3 0.425 0.376
0.471
(±0.003)
0.603
(±0.008)
0.548
(±0.012)
0.702
(±0.010)
0.597
(±0.002)
0.720
(±0.013)
0.773 0.386
Mistral-7B-
Instruct-v0.3
0.619 0.642
0.635
(±0.007)
0.630
(±0.007)
0.664
(±0.005)
0.664
(±0.004)
0.690
(±0.008)
0.667
(±0.007)
0.765 0.650
Qwen2.5-3B 0.458 0.500
0.470
(±0.008)
0.538
(±0.008)
0.520
(±0.005)
0.560
(±0.008)
0.554
(±0.006)
0.567
(±0.009)
0.724 0.629
Qwen2.5-3B-
Instruct
0.534 0.620
0.599
(±0.007)
0.584
(±0.003)
0.633
(±0.009)
0.600
(±0.007)
0.656
(±0.007)
0.609
(±0.006)
0.700 0.726
Qwen2.5-7B 0.480 0.623
0.474
(±0.008)
0.608
(±0.007)
0.504
(±0.004)
0.622
(±0.007)
0.556
(±0.006)
0.629
(±0.003)
0.761 0.656
Qwen2.5-7B-
Instruct
0.542 0.676
0.558
(±0.007)
0.637
(±0.005)
0.620
(±0.004)
0.623
(±0.005)
0.659
(±0.005)
0.614
(±0.005)
0.718 0.793
gpt-4o-mini - 0.643 -
0.682
(±0.002)
0.687
(±0.005)
0.704
(±0.008)
gpt-4.1-mini - 0.797 -
0.822
(±0.005)
0.841
(±0.008)
0.846
(±0.005)
claude-3-5-
haiku-latest
- 0.722 -
0.505
(±0.016)
0.476
(±0.009)
0.633
(±0.004)
2-3B models
average
0.456 0.404 0.484 0.446 0.556 0.464 0.585 0.469 0.731 0.428
7-8B models
average
0.517 0.569 0.545 0.631 0.609 0.672 0.649 0.682 0.769 0.655
api models
average
- 0.721 - 0.670 - 0.668 - 0.728 - -
base models
average
0.448 0.408 0.470 0.513 0.547 0.564 0.591 0.582 0.760 0.507
instruct-tuned
models average
0.533 0.589 0.567 0.59 0.625 0.601 0.652 0.599 0.746 0.608
total average 0.491 0.538 0.518 0.573 0.586 0.598 0.622 0.615 0.753 0.558
Table 11: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the simple prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.443 0.176
0.427
(±0.002)
0.260
(±0.002)
0.515
(±0.005)
0.272
(±0.009)
0.560
(±0.007)
0.266
(±0.014)
0.720 0.259
gemma-1.1-
2b-it
0.402 0.263
0.436
(±0.004)
0.322
(±0.006)
0.542
(±0.006)
0.319
(±0.004)
0.570
(±0.005)
0.309
(±0.007)
0.753 0.334
gemma-7b 0.501 0.404
0.484
(±0.006)
0.568
(±0.009)
0.581
(±0.008)
0.638
(±0.008)
0.630
(±0.002)
0.675
(±0.006)
0.803 0.709
gemma-1.1-
7b-it
0.656 0.614
0.680
(±0.006)
0.627
(±0.009)
0.728
(±0.006)
0.657
(±0.006)
0.748
(±0.005)
0.668
(±0.006)
0.775 0.692
Llama-3.2-3B 0.416 0.324
0.471
(±0.006)
0.381
(±0.008)
0.561
(±0.007)
0.451
(±0.008)
0.590
(±0.008)
0.491
(±0.009)
0.754 0.282
Llama-3.2-3B-
Instruct
0.454 0.516
0.477
(±0.006)
0.464
(±0.010)
0.546
(±0.005)
0.452
(±0.008)
0.589
(±0.004)
0.450
(±0.006)
0.769 0.386
Llama-3.1-8B 0.462 0.477
0.499
(±0.005)
0.512
(±0.013)
0.598
(±0.006)
0.591
(±0.008)
0.645
(±0.005)
0.630
(±0.009)
0.817 0.608
Llama-3.1-8B-
Instruct
0.519 0.684
0.557
(±0.005)
0.640
(±0.004)
0.628
(±0.002)
0.663
(±0.004)
0.665
(±0.007)
0.681
(±0.014)
0.784 0.762
Mistral-7B-v0.3 0.446 0.297
0.475
(±0.006)
0.571
(±0.013)
0.560
(±0.011)
0.683
(±0.010)
0.608
(±0.002)
0.694
(±0.007)
0.781 0.427
Mistral-7B-
Instruct-v0.3
0.596 0.630
0.635
(±0.006)
0.622
(±0.007)
0.685
(±0.005)
0.630
(±0.011)
0.707
(±0.009)
0.632
(±0.007)
0.765 0.629
Qwen2.5-3B 0.427 0.531
0.474
(±0.007)
0.490
(±0.007)
0.540
(±0.003)
0.499
(±0.006)
0.574
(±0.007)
0.500
(±0.007)
0.722 0.648
Qwen2.5-3B-
Instruct
0.474 0.585
0.530
(±0.008)
0.511
(±0.006)
0.628
(±0.007)
0.512
(±0.006)
0.665
(±0.010)
0.513
(±0.006)
0.696 0.713
Qwen2.5-7B 0.454 0.628
0.483
(±0.009)
0.606
(±0.005)
0.541
(±0.007)
0.622
(±0.007)
0.587
(±0.002)
0.639
(±0.007)
0.745 0.706
Qwen2.5-7B-
Instruct
0.536 0.636
0.549
(±0.007)
0.634
(±0.009)
0.634
(±0.010)
0.634
(±0.008)
0.674
(±0.006)
0.636
(±0.004)
0.734 0.764
gpt-4o-mini - 0.621 -
0.659
(±0.008)
0.688
(±0.005)
0.700
(±0.009)
gpt-4.1-mini - 0.842 -
0.845
(±0.007)
0.857
(±0.005)
0.867
(±0.005)
claude-3-5-
haiku-latest
- 0.688 -
0.495
(±0.006)
0.470
(±0.008)
0.613
(±0.013)
2-3B models
average
0.436 0.399 0.469 0.405 0.555 0.418 0.591 0.422 0.736 0.437
7-8B models
average
0.521 0.546 0.545 0.598 0.619 0.640 0.658 0.657 0.775 0.662
api models
average
- 0.717 - 0.666 - 0.672 - 0.727 - -
base models
average
0.450 0.405 0.473 0.484 0.557 0.537 0.599 0.556 0.763 0.520
instruct-tuned
models average
0.520 0.561 0.552 0.546 0.627 0.552 0.660 0.556 0.754 0.611
total average 0.485 0.524 0.513 0.542 0.592 0.567 0.629 0.586 0.758 0.566
Table 12: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the definition prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
zero-shot 1-shot 5-shot 10-shot After SFT
evaluation
setting
comple-
tion
option
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
(±SD)
option
(±SD)
comple-
tion
option
gemma-2b 0.404 0.257
0.423
(±0.003)
0.253
(±0.005)
0.511
(±0.003)
0.257
(±0.012)
0.547
(±0.004)
0.145
(±0.009)
0.496 0.514
gemma-1.1-
2b-it
0.394 0.260
0.440
(±0.007)
0.282
(±0.005)
0.524
(±0.009)
0.241
(±0.007)
0.561
(±0.009)
0.242
(±0.009)
0.519 0.476
gemma-7b 0.462 0.410
0.472
(±0.005)
0.483
(±0.008)
0.564
(±0.004)
0.531
(±0.007)
0.607
(±0.010)
0.549
(±0.008)
0.731 0.635
gemma-1.1-
7b-it
0.588 0.472
0.653
(±0.008)
0.510
(±0.009)
0.708
(±0.009)
0.501
(±0.004)
0.728
(±0.006)
0.533
(±0.011)
0.725 0.616
Llama-3.2-3B 0.445 0.306
0.463
(±0.005)
0.317
(±0.011)
0.550
(±0.008)
0.365
(±0.010)
0.582
(±0.006)
0.407
(±0.008)
0.567 0.553
Llama-3.2-3B-
Instruct
0.454 0.429
0.440
(±0.006)
0.387
(±0.009)
0.501
(±0.007)
0.381
(±0.008)
0.560
(±0.005)
0.389
(±0.006)
0.586 0.595
Llama-3.1-8B 0.442 0.402
0.466
(±0.004)
0.394
(±0.008)
0.566
(±0.006)
0.462
(±0.012)
0.624
(±0.005)
0.518
(±0.010)
0.726 0.642
Llama-3.1-8B-
Instruct
0.508 0.539
0.529
(±0.008)
0.515
(±0.010)
0.627
(±0.005)
0.554
(±0.004)
0.669
(±0.003)
0.580
(±0.006)
0.673 0.681
Mistral-7B-v0.3 0.454 0.236
0.485
(±0.004)
0.416
(±0.014)
0.551
(±0.010)
0.597
(±0.009)
0.591
(±0.007)
0.631
(±0.007)
0.586 0.545
Mistral-7B-
Instruct-v0.3
0.588 0.586
0.608
(±0.013)
0.551
(±0.010)
0.662
(±0.005)
0.556
(±0.008)
0.685
(±0.005)
0.558
(±0.005)
0.672 0.646
Qwen2.5-3B 0.415 0.463
0.461
(±0.005)
0.399
(±0.002)
0.532
(±0.005)
0.324
(±0.006)
0.568
(±0.006)
0.313
(±0.003)
0.662 0.615
Qwen2.5-3B-
Instruct
0.467 0.548
0.492
(±0.009)
0.473
(±0.009)
0.653
(±0.008)
0.227
(±0.008)
0.671
(±0.008)
0.495
(±0.011)
0.674 0.572
Qwen2.5-7B 0.443 0.555
0.464
(±0.005)
0.515
(±0.010)
0.537
(±0.008)
0.542
(±0.005)
0.577
(±0.003)
0.566
(±0.004)
0.691 0.689
Qwen2.5-7B-
Instruct
0.507 0.545
0.511
(±0.006)
0.527
(±0.008)
0.621
(±0.012)
0.526
(±0.007)
0.662
(±0.004)
0.544
(±0.003)
0.674 0.705
gpt-4o-mini - 0.595 -
0.620
(±0.014)
0.647
(±0.003)
0.669
(±0.006)
gpt-4.1-mini - 0.858 -
0.864
(±0.004)
0.876
(±0.003)
0.870
(±0.005)
claude-3-5-
haiku-latest
- 0.604 -
0.472
(±0.009)
0.537
(±0.009)
0.689
(±0.013)
2-3B models
average
0.430 0.377 0.453 0.352 0.545 0.299 0.582 0.332 0.584 0.554
7-8B models
average
0.499 0.468 0.524 0.489 0.605 0.534 0.643 0.560 0.685 0.645
api models
average
- 0.686 - 0.652 - 0.687 - 0.743 - -
base models
average
0.438 0.376 0.462 0.397 0.544 0.440 0.585 0.447 0.637 0.599
instruct-tuned
models average
0.501 0.483 0.525 0.464 0.614 0.427 0.648 0.477 0.646 0.613
total average 0.469 0.474 0.493 0.469 0.579 0.478 0.617 0.512 0.642 0.606
Table 13: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the detail prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
Following the classification of prompts introduced in Appendix O, we report results on Thunder-NUBench separately for the simple, definition, and detail prompt styles. The experiments are conducted under three evaluation regimes: (i) zero-shot, (ii) few-shot (1-, 5-, and 10-shot, averaged across seeds), and (iii) supervised fine-tuning (SFT). For SFT, we assess models only in the zero-shot setting in order to directly measure the effect of task-specific training without the influence of in-context examples.
Figure 4: Model performance on Thunder-NUBench with definition prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Figure 5: Model performance on Thunder-NUBench with detail prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Table 11, Table 12, and Table 13 present the complete numerical results for the simple, definition, and detail prompts, respectively. The same results are also summarized in graphical form in Figure 3, Figure 4, and Figure 5.
Across prompt types, we observe the following trends:
  • Simple prompts. Results for simple prompts are described in Section 5.2; we summarize them here only for completeness.
  • Definition prompts. Performance patterns closely resemble those of simple prompts. However, in the option-selection evaluation, 2-3B models show smaller gains from additional shots, while 7-8B models tend to achieve slightly higher accuracy after SFT compared to 10-shot prompting.
  • Detail prompts. For smaller (2-3B) models in the option-selection evaluation, adding more shots can actually reduce performance, reflecting difficulties with long and complex prompt instructions to these small models. Nonetheless, these models, along with larger ones, show larger SFT gains under detail prompts than under simple or definition prompts.
Overall, these findings show that prompt formulation plays a non-trivial role in shaping performance trends. In all cases, SFT consistently boosts zero-shot performance.
Appendix Q General Benchmark Performance after SFT
tasks
arc_
challenge
arc_
easy
gsm8k hellaswag
openbook-
qa
wino-
grande
avg
metric acc acc
exact_
match
acc_norm acc_norm acc
Before SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.432 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.756 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.702 0.607
After SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.431 0.746 0.365 0.713 0.411 0.671 0.556
7-8B models
average
0.524 0.819 0.536 0.798 0.453 0.721 0.642
base models
average
0.474 0.788 0.428 0.772 0.437 0.706 0.601
instruct-tuned
models average
0.495 0.787 0.497 0.751 0.432 0.693 0.609
total average 0.484 0.788 0.463 0.762 0.435 0.699 0.605
Before SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.433 0.745 0.341 0.713 0.407 0.669 0.551
7-8B models
average
0.528 0.819 0.545 0.799 0.455 0.725 0.645
base models
average
0.477 0.785 0.419 0.772 0.436 0.709 0.600
instruct-tuned
models average
0.498 0.790 0.496 0.753 0.433 0.694 0.611
total average 0.487 0.788 0.457 0.762 0.434 0.701 0.605
Before SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.429 0.746 0.359 0.713 0.409 0.668 0.554
7-8B models
average
0.523 0.816 0.551 0.8 0.458 0.723 0.645
base models
average
0.469 0.785 0.424 0.773 0.439 0.704 0.599
instruct-tuned
models average
0.496 0.787 0.513 0.753 0.434 0.695 0.613
total average 0.483 0.786 0.469 0.763 0.437 0.699 0.606
Table 14: Average performance of 2–3B / 7–8B, and base / instruction-tuned models on six general benchmarks before and after SFT on Thunder-NUBench. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer.
To assess whether supervised fine-tuning (SFT) on Thunder-NUBench affects performance on broader natural language understanding tasks, we evaluate models on six widely used benchmarks: ARC-Challenge, ARC-Easy, GSM8K, HellaSwag, OpenBookQA, and WinoGrande (clark2018think; cobbe2021training; zellers2019hellaswag; mihaylov2018can; sakaguchi2021winogrande). These datasets cover a diverse range of domains, including commonsense reasoning, scientific knowledge, mathematics, and reading comprehension.
Following the classification of prompts introduced in Appendix O, SFT is conducted separately for the simple, definition, and detail prompt styles. We then report post-SFT zero-shot performance across the six benchmarks to verify whether task-specific training on Thunder-NUBench preserves general capabilities.
Table 14 summarizes the results. We find that performance on general benchmarks remains broadly stable after SFT, indicating that fine-tuning on Thunder-NUBench does not substantially harm general capabilities.
Appendix R Total Analysis of Model Predictions on the Thunder-NUBench
This appendix extends the error analysis presented in Section 5.3, providing the complete results. In particular, we examine (i) incorrect choice distributions and (ii) confusion rates for local negation categories, comparing models of different sizes (2-3B vs. 7-8B) and training paradigms (pretrained vs. instruction-tuned) under both zero-shot and few-shot conditions.
We report few-shot results using a fixed random seed (1234), which corresponds to the default seed used in the LM Evaluation Harness framework. Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make fine-grained confusion analysis less interpretable.
For consistency with the main text, we report results only under the simple prompt, which serves as the default evaluation setting throughout the paper.
We organize the complete prediction analysis by model family.
  • Results for the Gemma family are reported in Table 15, Table 16.
  • Results for the LLaMA family are reported in Table 17, Table 18.
  • Results for the Mistral family are reported in Table 19, Table 20.
  • Results for the Qwen family are reported in Table 21, Table 22.
  • Results for API models are reported in Table 23.
Each table follows the same format, reporting error rates, incorrect choice distributions, and confusion rates across local negation subcategories under zero-shot, few-shot, and SFT conditions.
In the option-selection evaluation setting, we also track cases labeled as “Answer Format Wrong.” This category captures instances where the model’s output does not follow the required answer format (selecting strictly one of A, B, C, or D). Because such responses cannot be mapped to a specific incorrect option, they are not included in the incorrect choice distribution or confusion rate. Instead, we report their raw counts alongside the other error statistics. This also serves as an indicator of the model’s ability to follow output-format instructions.
Overall, we observe consistent trends that larger models (7-8B) achieve lower error rates than smaller ones (2-3B), and introduction-tuned variants generally outperform base models. Supervised fine-tuning (SFT) tends to reduce error rates. Meanwhile, incorrect choice distributions and local negation confusion rates vary across models and settings, showing that this analysis can serve as a useful tool to identify which aspects of negation remain particularly challenging.
Notably, we also find some unusual behaviors. For example, the Gemma-1.1-2B-it model after SFT produces 1,239 outputs with incorrect answer formats in the option-selection setting, indicating formatting issues rather than genuine reasoning errors. Similarly, for API models, Claude 3.5 Haiku shows increasing formatting errors as more shots are added, which directly degrades performance.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2b baseline zero-shot 0.563 74.08 21.41 4.51 29.17 32.14 67.69 44.19
1-shot 0.558 74.54 21.34 4.13 32.37 31.49 63.27 45.16
5-shot 0.474 81.61 15.89 2.51 26.28 30.19 55.78 48.06
10-shot 0.421 82.49 16.01 1.51 23.72 23.38 48.64 48.06
sft zero-shot 0.268 84.91 14.20 0.89 13.14 18.51 26.87 35.48
gemma-1.1- 2b-it baseline zero-shot 0.612 59.20 32.25 8.55 31.09 28.25 55.44 35.48
1-shot 0.565 66.76 28.47 4.77 32.37 29.55 55.10 39.35
5-shot 0.471 66.50 27.44 6.06 25.64 24.35 47.96 31.94
10-shot 0.468 68.98 26.44 4.58 26.28 22.40 49.66 35.48
sft zero-shot 0.250 81.27 17.78 0.95 11.86 15.26 26.87 30.00
gemma-7b baseline zero-shot 0.546 72.28 24.24 3.48 29.49 32.14 59.18 42.90
1-shot 0.535 77.74 18.10 4.15 30.77 31.82 57.82 51.61
5-shot 0.437 84.39 13.79 1.81 25.64 26.95 47.96 51.94
10-shot 0.369 84.95 13.33 1.72 22.44 23.05 35.71 48.06
sft zero-shot 0.203 81.64 16.80 1.56 10.26 11.36 18.03 28.71
gemma-1.1- 7b-it baseline zero-shot 0.337 55.06 38.35 6.59 9.29 14.29 26.53 26.77
1-shot 0.310 69.57 26.09 4.35 14.10 13.31 26.53 35.16
5-shot 0.279 73.01 25.00 1.99 13.14 13.64 20.07 37.10
10-shot 0.232 73.72 24.91 1.37 10.58 7.79 15.31 36.77
sft zero-shot 0.214 80.74 17.78 1.48 10.90 11.04 17.35 31.94
Table 15: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2b baseline zero-shot 0.803 274 33.29 33.42 33.29 20.51 23.05 21.43 15.48
1-shot 0.750 0 31.18 29.6 39.22 22.44 26.62 23.47 23.87
5-shot 0.722 0 30.63 23.82 45.55 20.19 23.70 28.23 19.35
10-shot 0.736 0 29.42 22.74 47.84 21.47 23.70 23.81 20.32
sft zero-shot 0.741 6 32.54 34.48 32.97 22.44 27.92 25.51 22.90
gemma-1.1- 2b-it baseline zero-shot 0.738 4 34.23 27.86 37.90 24.04 28.25 25.85 25.48
1-shot 0.654 0 46.67 18.91 34.42 29.81 35.71 30.61 29.68
5-shot 0.631 0 48.68 14.59 36.73 32.05 37.34 27.55 29.35
10-shot 0.660 0 43.87 11.06 45.07 29.49 34.42 25.85 29.35
sft zero-shot 0.994 1,239 85.71 0 14.29 0.96 0.97 1.02 0.97
gemma-7b baseline zero-shot 0.612 195 57.89 9.88 32.24 17.31 19.81 44.90 28.06
1-shot 0.351 0 69.07 12.19 18.74 20.51 14.61 40.82 24.84
5-shot 0.307 0 77.52 15.50 6.98 22.76 17.86 35.03 22.90
10-shot 0.288 0 81.54 13.50 4.96 24.36 16.23 31.29 25.16
sft zero-shot 0.278 1 70.57 12.57 16.86 11.54 14.29 28.91 26.45
gemma-1.1- 7b-it baseline zero-shot 0.355 0 64.51 20.09 15.40 16.99 15.26 40.14 22.90
1-shot 0.307 0 63.05 15.76 21.19 16.03 11.36 34.01 19.03
5-shot 0.283 0 63.87 18.77 17.37 16.99 9.42 29.59 19.03
10-shot 0.289 0 65.93 17.31 16.76 19.23 8.77 30.61 20.32
sft zero-shot 0.269 2 76.26 16.91 6.820 22.12 14.61 27.21 20.32
Table 16: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.2- 3B baseline zero-shot 0.556 70.04 23.11 6.85 24.68 30.19 62.59 44.19
1-shot 0.519 75.99 20.49 3.52 25.32 30.19 60.54 47.42
5-shot 0.419 78.41 18.94 2.65 20.19 25.32 46.60 43.87
10-shot 0.401 80.20 16.83 2.97 22.44 24.03 43.20 43.23
sft zero-shot 0.267 82.79 15.43 1.78 14.42 17.53 25.85 33.55
Llama-3.2- 3B-Instruct baseline zero-shot 0.528 72.67 23.87 3.45 23.72 29.55 50.68 54.84
1-shot 0.502 75.36 23.22 1.42 28.85 27.27 48.98 51.29
5-shot 0.441 76.08 22.84 1.08 25.64 26.62 39.80 46.45
10-shot 0.410 76.79 21.86 1.35 25.64 24.68 35.03 44.52
sft zero-shot 0.255 83.80 14.95 1.25 12.82 13.31 27.55 34.52
Llama-3.1- 8B baseline zero-shot 0.562 70.62 21.33 8.05 25.64 30.84 64.29 43.87
1-shot 0.507 77.78 18.15 4.07 25.96 29.87 59.18 48.39
5-shot 0.393 83.87 14.31 1.81 20.83 22.40 47.62 45.81
10-shot 0.358 84.51 13.72 1.77 20.19 23.05 40.82 41.29
sft zero-shot 0.203 85.55 12.50 1.95 8.97 11.69 18.71 32.26
Llama-3.1- 8B-Instruct baseline zero-shot 0.488 70.57 26.18 3.25 24.36 25.97 54.76 37.74
1-shot 0.419 73.86 24.43 1.70 22.76 25.97 44.56 34.84
5-shot 0.347 80.55 18.76 0.69 18.91 19.16 39.80 37.74
10-shot 0.329 81.20 18.31 0.48 19.55 20.45 34.01 36.45
sft zero-shot 0.244 86.04 12.34 1.62 12.18 12.66 28.23 33.87
Table 17: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.2- 3B baseline zero-shot 0.688 0 42.21 25.95 31.83 22.12 33.12 32.65 31.94
1-shot 0.570 0 62.59 14.74 22.67 31.09 32.14 45.58 38.71
5-shot 0.548 0 74.10 14.33 11.58 34.62 35.06 50.34 47.74
10-shot 0.500 0 78.13 13.00 8.87 35.90 31.82 43.88 49.68
sft zero-shot 0.658 0 39.40 33.01 27.59 23.40 31.17 27.89 24.52
Llama-3.2- 3B-Instruct baseline zero-shot 0.474 8 63.90 10.34 25.76 19.55 28.25 60.20 16.77
1-shot 0.461 0 71.60 12.56 15.83 23.40 25.97 61.22 26.77
5-shot 0.472 0 74.62 10.92 14.45 29.81 28.25 63.27 25.16
10-shot 0.467 0 73.85 10.36 15.79 33.01 25.00 62.24 23.23
sft zero-shot 0.437 87 61.85 7.11 31.03 11.22 18.51 43.54 21.61
Llama-3.1- 8B baseline zero-shot 0.541 0 47.95 6.74 45.31 23.08 24.03 41.50 19.03
1-shot 0.454 0 67.66 4.55 27.80 27.56 28.25 47.62 23.87
5-shot 0.362 0 78.95 6.58 14.47 30.45 21.75 40.82 25.16
10-shot 0.316 0 82.46 6.02 11.53 30.77 21.75 31.97 23.23
sft zero-shot 0.444 48 58.79 27.34 13.87 20.83 20.13 37.41 20.65
Llama-3.1-8B- Instruct baseline zero-shot 0.256 12 76.53 15.11 8.36 16.67 18.83 19.05 23.23
1-shot 0.290 0 71.51 21.37 7.12 19.23 18.18 29.25 19.03
5-shot 0.269 0 74.04 21.83 4.13 20.83 16.88 25.17 19.35
10-shot 0.290 0 80.00 14.52 5.48 25.00 20.78 26.87 22.90
sft zero-shot 0.254 31 72.32 19.03 8.65 15.06 17.86 21.09 14.52
Table 18: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B- v0.3 baseline zero-shot 0.575 76.97 18.48 4.55 26.60 32.14 66.33 58.39
1-shot 0.532 78.39 17.59 4.02 26.92 33.12 59.52 53.23
5-shot 0.439 83.57 14.98 1.44 22.12 25.65 51.70 52.58
10-shot 0.401 84.39 13.83 1.78 21.15 24.68 47.28 47.10
sft zero-shot 0.227 88.81 9.09 2.10 8.65 13.31 29.93 31.61
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.381 67.29 30.83 1.88 13.46 18.51 30.95 42.90
1-shot 0.365 72.61 26.09 1.30 15.38 20.45 34.69 39.03
5-shot 0.331 72.66 26.86 0.48 14.10 15.58 31.97 37.74
10-shot 0.304 73.63 25.85 0.52 12.18 14.29 30.61 35.48
sft zero-shot 0.236 85.52 13.80 0.67 10.26 12.01 27.21 33.87
Table 19: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B- v0.3 baseline zero-shot 0.624 351 53.21 21.56 25.23 16.99 18.18 26.53 14.52
1-shot 0.410 0 67.12 16.63 16.25 30.77 24.68 28.91 29.03
5-shot 0.306 0 83.16 11.92 4.92 26.28 24.35 24.49 29.68
10-shot 0.272 0 85.42 8.45 6.12 23.40 22.73 21.09 28.39
sft zero-shot 0.614 0 40.05 29.46 30.49 22.44 27.92 28.57 22.58
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.358 1 65.56 25.78 8.67 27.24 23.05 27.55 18.71
1-shot 0.370 0 61.03 34.90 4.07 27.24 20.78 26.19 19.03
5-shot 0.335 0 68.96 29.86 1.18 28.85 19.81 26.19 20.32
10-shot 0.331 0 71.22 27.10 1.68 28.53 22.08 25.51 20.97
sft zero-shot 0.351 1 67.57 21.32 11.11 23.08 23.05 33.67 18.06
Table 20: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen2.5- 3B baseline zero-shot 0.542 67.69 26.02 6.29 24.68 30.19 55.10 42.26
1-shot 0.522 67.63 25.53 6.84 24.36 28.25 55.78 38.06
5-shot 0.481 70.84 23.89 5.27 25.00 28.57 50.00 37.74
10-shot 0.437 73.14 22.14 4.72 23.40 25.32 45.58 38.06
sft zero-shot 0.276 85.34 13.22 1.44 14.74 20.78 30.95 30.97
Qwen2.5- 3B-Instruct baseline zero-shot 0.466 68.03 28.57 3.40 19.23 28.25 42.86 40.97
1-shot 0.393 67.68 28.28 4.04 16.67 20.45 42.52 30.65
5-shot 0.355 77.01 21.88 1.12 19.55 22.40 39.46 31.94
10-shot 0.351 77.83 21.04 1.13 21.15 23.05 35.71 32.90
sft zero-shot 0.300 85.19 13.49 1.32 15.71 22.40 33.33 34.19
Qwen2.5- 7B baseline zero-shot 0.520 70.73 24.54 4.73 24.68 31.17 53.06 43.55
1-shot 0.534 71.47 25.56 2.97 29.17 31.17 56.12 41.61
5-shot 0.492 71.13 25.81 3.06 24.36 28.25 51.70 40.65
10-shot 0.447 74.07 23.98 1.95 24.04 24.68 45.58 42.58
sft zero-shot 0.240 87.75 11.59 0.66 15.06 15.91 22.45 33.23
Qwen2.5- 7B-Instruct baseline zero-shot 0.458 65.16 31.54 3.29 18.27 29.22 39.80 36.13
1-shot 0.438 69.02 30.07 0.91 22.12 23.70 43.88 35.48
5-shot 0.378 71.85 27.94 0.21 19.55 19.16 37.41 36.13
10-shot 0.335 73.76 25.77 0.47 19.23 18.18 30.61 34.19
sft zero-shot 0.282 88.76 10.96 0.28 16.03 19.48 31.29 36.77
Table 21: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen2.5- 3B baseline zero-shot 0.500 0 38.99 19.49 41.52 23.08 19.16 26.87 11.61
1-shot 0.463 0 45.89 27.23 26.88 25.00 21.43 26.53 14.84
5-shot 0.431 0 44.75 25.60 29.65 24.04 19.81 23.13 12.58
10-shot 0.439 0 45.85 23.10 31.05 25.00 21.10 25.17 11.94
sft zero-shot 0.373 2 51.07 13.25 35.68 17.31 15.58 26.19 19.35
Qwen2.5- 3B-Instruct baseline zero-shot 0.380 0 49.27 40.29 10.44 20.19 15.91 24.15 17.10
1-shot 0.419 0 53.22 32.39 14.39 26.28 19.16 28.23 18.39
5-shot 0.398 2 54.40 31.60 14.00 27.24 19.16 29.25 13.55
10-shot 0.388 0 56.44 29.24 14.31 26.60 20.13 30.61 13.23
sft zero-shot 0.274 1 59.13 34.49 6.38 12.18 13.31 21.43 20.00
Qwen2.5- 7B baseline zero-shot 0.377 0 68.21 15.79 16.00 28.53 22.73 36.39 18.71
1-shot 0.392 0 70.65 15.59 13.77 33.97 26.30 38.10 16.13
5-shot 0.377 0 72.00 15.16 12.84 33.65 23.05 38.78 16.77
10-shot 0.372 0 73.13 14.93 11.94 32.69 23.05 40.48 16.45
sft zero-shot 0.344 0 82.03 13.13 4.84 30.77 24.35 43.54 18.39
Qwen2.5-7B- Instruct baseline zero-shot 0.324 0 62.84 29.34 7.82 23.72 22.73 23.47 14.19
1-shot 0.363 0 61.79 31.66 6.55 30.45 21.10 27.21 13.87
5-shot 0.368 0 66.38 26.51 7.11 31.41 19.81 34.69 15.16
10-shot 0.390 0 67.07 25.81 7.11 34.62 20.78 35.03 17.74
sft zero-shot 0.207 0 78.54 17.62 3.83 18.59 20.13 17.69 10.65
Table 22: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gpt-4o- mini baseline zero-shot 0.357 0 34.67 65.33 0 17.31 8.44 13.27 11.94
1-shot 0.317 0 40.00 59.75 0.25 19.87 7.47 13.95 10.97
5-shot 0.316 0 38.35 61.15 0.50 19.23 7.14 11.90 11.61
10-shot 0.298 0 42.02 57.98 0 19.23 8.12 12.93 11.29
gpt-4.1- mini baseline zero-shot 0.203 0 51.95 48.05 0 14.10 6.49 8.16 14.52
1-shot 0.184 0 64.66 35.34 0 16.35 6.17 10.20 16.13
5-shot 0.156 0 65.48 34.52 0 13.46 6.17 7.14 15.16
10-shot 0.153 0 64.77 35.23 0 14.74 5.19 4.76 15.81
claude-3-5- haiku-latest baseline zero-shot 0.278 0 67.81 30.48 1.71 22.76 14.29 15.99 24.52
1-shot 0.477 434 83.93 15.48 0.60 20.19 6.49 10.54 8.71
5-shot 0.527 574 95.56 4.44 0 13.78 3.57 5.44 5.16
10-shot 0.368 324 94.29 5.00 0.71 20.83 6.17 5.44 10.32
Table 23: Error rates, incorrect choice distributions, and local negation confusion rates for the API models under zero-shot and few-shot conditions, evaluated in the option-selection setting.