### v4

# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

Abstract
Negation is a fundamental linguistic phenomenon that poses ongoing challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Current benchmarks often treat negation as a minor detail within broader tasks, such as natural language inference. Consequently, there is a lack of benchmarks specifically designed to evaluate comprehension of negation. In this work, we introduce Thunder-NUBench — a novel benchmark explicitly created to assess sentence-level understanding of negation in LLMs. Thunder-NUBench goes beyond identifying surface-level cues by contrasting standard negation with structurally diverse alternatives, such as local negation, contradiction, and paraphrase. This benchmark includes manually created sentence-negation pairs and a multiple-choice dataset, allowing for a comprehensive evaluation of models’ understanding of negation.
Thunder-NUBench: A Benchmark for LLMs’
Sentence-Level Negation Understanding
Yeonkyoung So1, Gyuseong Lee1, Sungmok Jung1, Joonhak Lee1, JiA Kang1, Sangho Kim1, Jaejin Lee1,2 1Graduate School of Data Science, Seoul National University 2Dept. of Computer Science and Engineering, Seoul National University {kathy1028,ksnannaya,tjdahrwjd,hmjelee,jia6776,ksh4931,jaejin}@snu.ac.kr
1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely associated with various human communicative abilities, such as denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation still plays a significant role in language; approximately 25% of sentences in English texts contain some form of negation (Sarabi and Blanco, 2016; Hossain et al., 2020; Horn and Wansing, 2025). This prevalence and its impact on meaning make accurate interpretation of negation crucial for several natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI) (Khandelwal and Sawant, 2020; Hosseini et al., 2021; Singh et al., 2023). Recent studies have shown that effectively managing negation is important even for multi-modal language models (Quantmeyer et al., 2024; Alhamoud et al., 2025; Park et al., 2025).
Figure 1: Illustration of the Thunder-NUBench task. Only the standard negation reverses the truth value of the original sentence, while other options differ in scope or semantics.
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find it more difficult to process and comprehend negated statements compared to affirmative ones (Wales and Grieve, 1969; Sarabi and Blanco, 2016). Similarly, many studies indicate that pretrained language models (PLMs) struggle to interpret negation accurately. For example, models like BERT (Devlin et al., 2019) and even large language models (LLMs) such as GPT-3 (Brown et al., 2020) often have difficulty distinguishing between negated and affirmative statements. These models tend to rely on superficial cues, which can result in incorrect outputs when negation is involved (Kassner and Schütze, 2020; Hossain et al., 2022a; Truong et al., 2023).
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources either treat negation as a minor aspect of broader tasks or focus solely on narrow syntactic detection, often emphasizing encoder-based models (Hossain et al., 2020; Geiger et al., 2020; Truong et al., 2022; Anschütz et al., 2023). To address this gap, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset explicitly designed to evaluate LLMs’ sentence-level comprehension of negation. [^1] Thunder-NUBench is publicly available at https://huggingface.co/datasets/thunder-research-group/SNU_Thunder-NUBench. Our benchmark is structured as a multiple-choice question (MCQ) task: given an original sentence, the model must select the correct standard negation from four options. The other three choices (local negation, contradiction, and paraphrase) are carefully designed distractors that test whether models truly grasp semantic scope and logical oppositions.
The contributions of this paper are summarized as follows:
  • We define standard negation within the framework of sentential logic. Grounding standard negation in logical structure not only clarifies its role in natural language but also supports the evaluation and enhancement of reasoning in LLMs.
  • We introduce a manually created benchmark that includes a dataset of sentence-negation pairs for fine-tuning, along with a multiple-choice evaluation task.
  • We conduct systematic evaluations of decoder-based LLMs across model families, scales, and training methods to analyze variation in negation understanding.
Thunder-NUBench provides valuable insights into the semantic reasoning abilities of language models and serves as a robust standard for future research focused on understanding negation.
2 Related Work
Dimension Negation Type Definition Example
Scope
Clausal Negation
( = Sentential
Negation)
Negation that applies to the entire clause or
sentence. This typically involves the use of "not",
or its contracted form "n’t" with auxiliary verbs.
He speaks English fluently.
$\rightarrow$ He doesn’t speak English fluently.
Subclausal Negation
( = Constituent /
Local Negation)
Negation that focuses on negating a specific
part of a clause, such as a word or phrase,
rather than the entire clause.
He speaks English fluently.
$\rightarrow$ He speaks English, but not fluently.
Form
Morphological
Negation
Negation expressed through affixes attached to words
such as prefixes like "un-", "in-", "dis-", or suffixes
like "-less".
She is happy.
$\rightarrow$ She is unhappy.
Syntactic Negation
Negation expressed through separate words (particles)
in the syntax, such as "not", "never", "no", etc.
She is happy.
$\rightarrow$ She is not happy.
Target Verbal Negation
Negation that applies directly to the verb
or verb phrase.
They have finished the work.
$\rightarrow$ They have not finished the work.
Non-verbal
Negation
Negation that negates elements other than the verb.
There is milk in the fridge.
$\rightarrow$ There is no milk in the fridge.
Table 1: Typology of negation.
Negation detection and scope resolution.
Early work in negation detection and scope resolution primarily relied on rule-based systems and handcrafted heuristics, especially in domain-specific contexts like clinical texts. While these systems are effective, they lack flexibility across different domains (Chapman et al., 2001; de Albornoz et al., 2012; Ballesteros et al., 2012; Basile et al., 2012). Traditional machine learning methods, such as Support Vector Machines (SVMs) (Hearst et al., 1998) and Conditional Random Fields (CRFs) (Sutton et al., 2012), were introduced later; however, they too are limited to narrow domains (Morante et al., 2008; Morante and Daelemans, 2009; Read et al., 2012; Li and Lu, 2018).
More recently, deep learning approaches employing Convolutional Neural Networks (CNNs) (O’shea and Nash, 2015) and Bidirectional Long Short-Term Memory (BiLSTM) networks (Siami-Namini et al., 2019) have enhanced performance by providing improved contextual embeddings and sequence modeling (Fancellu et al., 2016; Bhatia et al., 2019). Pretrained transformer models like BERT have been employed through transfer learning techniques (e.g., NegBERT (Khandelwal and Sawant, 2020)), significantly increasing the accuracy of negation detection tasks. Nonetheless, these methods still largely focus on syntactic span detection, leaving deeper semantic understanding of negation a challenging area to tackle.
Negation-sensitive subtasks of NLU.
Negation understanding has become increasingly important in natural language understanding (NLU) tasks (Hosseini et al., 2021). However, existing NLU benchmarks, such as SNLI (Bowman et al., 2015) for natural language inference (NLI), CommonsenseQA (Talmor et al., 2019) for Question Answering (QA), SST-2 (Socher et al., 2013) for sentiment analysis, STS-B (Cer et al., 2017) for textual similarity and paraphrasing, have been criticized for not adequately addressing the semantic impact of negation (Hossain et al., 2022a; Rezaei and Blanco, 2024). These datasets contain relatively few instances of negation or include negations that are not crucial to task performance, allowing language models to achieve high accuracy even when they completely ignore negation.
Recent studies, including NegNLI (Hossain et al., 2020), MoNLI (Geiger et al., 2020), NaN-NLI (Truong et al., 2022), NoFEVER-ML and NoSNLI-ML (Vrabcová et al., 2025), have introduced benchmarks for NLU that includes negation. These studies show that model performance significantly declines when negation plays a crucial role in affecting the outcome (Naik et al., 2018; Yanaka et al., 2019; Hartmann et al., 2021; Hossain et al., 2022b; Hossain and Blanco, 2022; She et al., 2023; Anschütz et al., 2023). These findings suggest that current language models tend to depend on superficial linguistic patterns rather than a genuine understanding of semantics.
Limitations of distributional semantics.
Distributional semantics (Harris, 1954; Sahlgren, 2008) aims to create models that learn semantic representations based on patterns of word co-occurrences (Boleda, 2020; Lenci et al., 2022) and capture broad semantic relationships; however, it encounters significant challenges with negation. Negated expressions, such as "not good," often appear in similar contexts as their affirmative counterparts, like "good." As a result, models tend to generate similar vector representations for these expressions, despite their opposing meanings. Previous research has pointed out this limitation, showing that PLMs struggle to capture the subtle semantic differences introduced by antonyms and the reversal of polarity (Rimell et al., 2017; Jumelet and Hupkes, 2018; Niwa et al., 2021; Jang et al., 2022; Vahtola et al., 2022). Studies have further suggested that models like BERT find it difficult to distinguish between affirmative and negated contexts (Kassner and Schütze, 2020; Ettinger, 2020).
Negations in generative language models.
Recent research on understanding negation has primarily focused on bidirectional models, such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), which have demonstrated strong performance in NLU and negation detection tasks. However, with the emergence of generative foundation models like GPT (Radford et al., 2018) and LLaMA (Touvron et al., 2023), attention has shifted towards evaluating how these models handle negation. Studies have shown that these generative models often exhibit a positive bias and struggle with producing or interpreting negated statements (Truong et al., 2023; Chen et al., 2023; García-Ferrero et al., 2023). Although some benchmarks, such as CONDAQA (Ravichander et al., 2022) and ScoNe (She et al., 2023), reveal these limitations, there is still a lack of robust evaluation resources specifically designed for negation understanding of generative models.
Building on previous studies, this paper focuses on sentence-level negation as a core linguistic operation. While prior negation-related NLI or QA benchmarks evaluate negation within broader inference or comprehension tasks, Thunder-NUBench specifically tests whether models can distinguish standard negation, the logical reversal of the sentence, from closely related distractors. As a result, Thunder-NUBench enables a more direct examination of sentence-level negation than is afforded by NLI or QA benchmarks, where negation is evaluated only as part of broader inference tasks.
Type Definition
Base case
If $P$ is an atomic proposition, $\text{Neg}(P)$ is the proposition where the main predicate of $P$ is negated.
Inductive step Conjunction $\text{Neg}(P\;\text{and}\;Q)\equiv\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$,  $\text{Neg}(P\;\text{but}\;Q)\equiv\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$
Disjunction $\text{Neg}(P\;\text{or}\;Q)\equiv\text{Neg}(P)\;\text{and}\;\text{Neg}(Q)$
Implication $\text{Neg}(\text{if }P,Q)\equiv\text{Neg}(\text{Neg}(P)\;\text{or}\;Q)\equiv P\;\text{and}\;\text{Neg}(Q)$ $\text{Neg}(P\;\text{if and only if}\;Q)\equiv\text{Neg}(\text{if }P,Q\;\text{and}\;\text{if }Q,P)$
Table 2: Standard negation. $P$ and $Q$ stand for propositions. In addition to and, or, and if, other natural language connectives such as when are also considered, and their negations follow the same principles depending on their function.
3 Scope and Categorization of Negation
In this work, we aim to clarify the concept of negation by introducing a typology that clearly outlines its semantic boundaries and differentiates it from related, yet distinct, phenomena. This typology organizes various forms of meaning reversal into logically consistent categories, allowing for a more precise and systematic evaluation of how language models handle negation.
3.1 Typology of Negation
Negation is a fundamental semantic and syntactic operation found in natural languages, used to convey denial, rejection, or the absence of a proposition. Hereafter, we denote our negation operation for a sentence $S$ as $\text{Neg}(S)$. In formal logic, negation flips the truth value of a proposition $P$: if $P$ is true, then $\text{Neg}(P)$ is false, and vice versa. Semantically, negation creates a binary opposition between a proposition and its affirmative counterpart, meaning that each one is the opposite of the other (Horn and Wansing, 2025).
Negation can be categorized along several dimensions: scope, form, and target (see Table 1). In terms of scope, negation may affect the entire clause (referred to as clausal negation) or only part of it (known as subclausal negation). Regarding form, negation can manifest as bound morphemes, such as prefixes and suffixes (morphological negation), or as separate syntactic elements like "not" or "never" (syntactic negation). Finally, depending on its target, negation can apply to the verb (verbal negation) or to other elements in the sentence (non-verbal negation(Zanuttini, 2001; Miestamo, 2007; Truong et al., 2022; Kletz et al., 2023).
3.2 Negation and Contradiction
Negation and contradiction are closely related concepts that are often conflated in NLP research (Jiang et al., 2021). Contradiction refers to the incompatibility of two propositions, meaning that they cannot both be true at the same time. While negation frequently serves as a primary mechanism for creating contradictions (by reversing the truth value of a proposition), contradictions can also arise from antonymy, numeric mismatches, or differences in structure and lexicon (further details can be found in Appendix A). For instance, the statements "An individual was born in France" and "An individual was born in Italy" are contradictory, but they are not negations, as the second statement does not reverse the truth of the first.
Many previous studies have overlooked the possibility that contradictions can exist independently of explicit negation. Recognizing this gap, we specifically examine the ability of LLMs to differentiate between negations and non-negated contradictions, highlighting the nuanced semantic distinctions that are involved.
Type Structure Explanation Local Negation Example
Relative clause
negation
A relative clause is a type of dependent clause that gives
extra details about a noun or noun phrase in the main sentence.
It usually begins with a relative pronoun such as
who, which, that, whom, or whose.
The man who owns the car is my
    neighbor.
$\rightarrow$ The man who does not own the car
     is my neighbor.
Participle clause
negation
A participle clause is a type of dependent clause that begins with
a participle (a verb form ending in -ing or a past participle).
It acts like an adverb, giving extra details about the main
clause, often showing time, reason, result, or sequence of actions.
Walking through the park, she found
    a lost wallet.
$\rightarrow$ Not walking through the park, she
     found a lost wallet.
Adverbial clause
negation
An adverbial clause is a dependent clause that acts like an adverb,
modifying a verb, adjective, or adverb. It gives information such
as time, reason, condition, or contrast. These clauses are introduced
by subordinating conjunctions like because, although, or while.
She stayed inside because it was
    raining.
$\rightarrow$ She stayed inside because it was
    not raining.
Compound
sentence
with
local negation
A compound sentence consists of two or more main clauses
joined by coordinating conjunctions such as and, but, or or.
If only one of these clauses is negated, the negation applies
only locally to that clause.
He submitted the report and
    attended the meeting.
$\rightarrow$ He submitted the report and did
     not attend the meeting.
Table 3: Typology of local negation.
3.3 Standard Negation
Standard negation refers to the typical form of negation applied to the declarative verbal main clause. It specifically negates the verb in a main clause (Miestamo, 2000). A main clause can function as a complete sentence on its own, consisting at a minimum of a subject and a predicate. This definition is grounded in the notion that the verb acts as the head of the clause (Miller and Miller, 2011).
Building on this traditional understanding, we treat standard negation as the process of reversing the truth value of the verb phrase in the main clause, which we will refer to as the main predicate in this paper. A verb phrase is headed by a verb and can consist of a single verb or a combination of auxiliaries, complements, and modifiers (e.g., "will call" and "is being promoted") (Lakoff, 1966). Since the main predicate conveys the core action or state of the clause, negating it effectively reverses the proposition of the entire sentence. In this context, this paper treats standard negation as a truth-functional operation that maps the main predicate to its complement set within the semantic space.
We further clarify the scope of standard negation within the typology presented in Table 1. Standard negation includes both clausal negation and verbal negation, as it reverses the meaning of the entire sentence by negating the main predicate. In terms of form, standard negation can employ both syntactic and morphological negation. Syntactically, standard negation often uses explicit negation particles, such as "not." Morphologically, it can involve complementary antonyms (for example, "alive" vs. "dead" or "true" vs. "false"), which occupy mutually exclusive semantic spaces, thus reversing the truth value of the proposition. In contrast, other types of antonyms, such as gradable antonyms (e.g., "happy" vs. "unhappy") and relational antonyms (e.g., "buy" vs. "sell") (Lehrer and Lehrer, 1982), do not strictly reverse truth values. Therefore, they are classified as contradictions rather than standard negation in this paper.
Atomic and Complex Propositions.
While this characterization of standard negation effectively defines standard negation for atomic propositions (elementary sentences that cannot be further decomposed) (Davis and Gillon, 2004), its application to complex sentences with multiple clauses requires a more thorough approach. In this paper, we treat an atomic proposition as a sentence that contains a single main predicate. Specifically, for propositions composed of multiple logically connected atomic statements, the method for reversing the truth value of the entire complex proposition can be ambiguous. In natural language, such logical structures typically appear as coordinated clauses (e.g., "$P$ and $Q$ or RR") or comma-separated lists connected by "and" or "or" (e.g., "$P,Q,$ and RR"). We treat these as equivalent to a sequence of binary conjunctions or disjunctions.
Definition of standard negation.
In this paper, standard negation refers to natural-language sentential negation, which is formally treated as logical negation within the framework of sentential logic (Enderton, 2001). To address the complexities involved, we define standard negation recursively by applying it pairwise over the logical structure of a sentence until only atomic propositions remain, ensuring that the truth value of the entire sentence is reversed even when it contains multiple coordinated clauses. Our definition of $\text{Neg}(\cdot)$ is presented in Table 2. Conditionals of the form "if $P$, QQ" are equivalent to $\text{Neg}(P)\;\text{or}\;Q$" in logic, and we adhere to this equivalence when defining their negation (more details can be found in Appendix C).
Figure 2: Dataset generation process.
3.4 Local Negation
We define local negation as a form of negation that specifically targets a verb phrase outside the main clause. While the term is often used interchangeably with subclausal negation, our focus is solely on local negation relating to subclausal and verbal negation. This concept applies to four types of sentence structures: relative clauses, participle clauses, adverbial clauses, and compound sentences (refer to Table 3 for more details).
In particular, conditional clauses, such as the "if PP" part in "if $P$, QQ" are categorized as adverbial clauses. In compound sentences, standard negation requires all main clauses to be negated in order to achieve sentence-level negation. If only a subset of the clauses is negated, this is considered local negation.
Local negation, in terms of structure, resembles standard negation, typically using explicit negation markers like "not." However, its scope is confined to a specific part of the sentence rather than encompassing the entire main clause. Because explicit cues such as "not" are still present, models that depend on shallow cue detection may be misled, failing to distinguish between standard negation and local negation.
4 Thunder-NUBench Dataset
We construct the Thunder-NUBench dataset through three main stages: (1) pre-processing, (2) generation, and (3) review. The overall workflow is illustrated in Figure 2.
Pre-processing.
We begin by extracting sentences from two primary corpora: (1) the Hover dataset (Jiang et al., 2020), designed for multi-hop fact extraction and claim verification, and (2) the Wikipedia Summary dataset (Scheepers, 2017), which contains concise summaries from English Wikipedia. We chose these datasets because their factual content and complex sentence structures are well-suited for developing a dataset aimed at understanding standard negation in complex, sufficiently lengthy sentences. Additionally, we automatically correct any grammatical errors and merge or split sentences as needed to create well-formed single-sentence units.
Generation.
We create two types of datasets from the pre-processed sentences: the sentence-negation pair dataset and the multiple choice dataset. In the sentence-negation pair dataset, each original sentence is paired with a manually crafted standard negation, as detailed in Section 3.3. In the multiple-choice dataset, each original sentence is presented with four options: a standard negation, a local negation, a contradiction, and a paraphrase. Each of them is described in Table 4. Together, these categories assess whether models truly understand semantic negation rather than relying on superficial cues.
Standard and local negation options are manually created rather than generated by LLMs. We have observed that LLMs often struggle to produce correct standard negations, frequently resulting in subclausal or local negations instead. They can also generate incorrect local negations, even when explicitly prompted to do otherwise. Since precise negation is essential to our benchmark, these options must be developed by humans to ensure the quality of the dataset. In contrast, contradiction and paraphrase options are initially created automatically using carefully designed prompts with the OpenAI API (OpenAI, 2025) and are then refined during the review process. Details of the GPT models and code used at each stage of the data generation process are provided in Appendix K.
Category Description
Standard Negation This category involves reversing the truth value of the main clause, which is the primary focus of the benchmark.
Local Negation In this case, negation is applied to a subordinate clause or a partial structure, which does not reverse the entire sentence.
Contradiction This category introduces conflicts with the original meaning through semantic changes, such as the use of antonyms, different numbers, or other entities, without employing explicit negation.
Paraphrase Here, the original meaning is preserved while the surface form is altered. Examples of paraphrases are intentionally constructed to vary the sentence structure and word choice significantly, ensuring that no additional information is added. As a result, the original sentence still entails its paraphrase. This category tests whether models mistakenly consider different surface forms as meaning reversals, even when the semantic meanings remain equivalent.
Table 4: Multiple choice categories included in Thunder-NUBench.
Review.
All constructed data undergo a multi-stage human review process (see Appendix I). A different author, separated from the creator, cross-checks each instance, and any disagreements are addressed in regular meetings to ensure consistency. Options for contradictions are reviewed only after the corresponding standard and local negations are finalized, as they must not overlap semantically. Consequently, the earlier negations are re-examined during the contradiction review and are cross-checked by multiple authors.
The guidelines for data generation and review are continuously updated, and any previously created data are revised accordingly (see Appendix J). This protocol ensures rigorous quality control and consistency throughout the benchmark.
Dataset statistics.
The final dataset includes a training set of sentence-negation pairs and a multiple-choice evaluation set (see Table 5). For few-shot prompting, we construct a demonstration set of 50 examples. These are carefully selected to have unique Wikipedia page indices to avoid any overlap with the test set. Furthermore, to provide the model with a balanced overview of the task, we match the distribution of local negation types (choice2_type) in the demonstration set to that of the overall dataset. This ensures that the demonstrations are representative and prevents the model from developing a biased strategy for specific negation types.
Dataset Split Count
Sentence-Negation Train 3,772
Multiple Choice Demonstration 50
Test 1,261
Total 5,083
Table 5: Thunder-NUBench statistics.
5 Experiments
5.1 Evaluation Setup
We evaluate models under two common Multiple-Choice Question Answering (MCQA) settings: (1) a completion-based evaluation, where the model assigns probabilities to each candidate by appending it as a continuation of the prompt, and (2) an option-selection evaluation, where the model selects from labeled options (A/B/C/D). To mitigate known issues such as selection or position bias in the option-selection setting, we randomly shuffle the order of the options (using random seed 42). In both settings, we use two instruction variants: a definition-based instruction (referred to as the definition instruction) and a step-by-step instruction (referred to as the detailed instruction). Details of the prompt templates and formatting are provided in Appendix M.
We report results for a diverse set of pretrained and instruction-tuned models, evaluated under zero-shot and few-shot settings with 1, 5, and 10 examples, as well as supervised fine-tuning (SFT). For SFT, models are trained on the sentence-negation pair dataset introduced in Thunder-NUBench. The full list of models, evaluation protocols, and fine-tuning details is provided in Appendix N.
5.2 Overall Performance and Key Findings
Effect of Model Size.
Performance tends to improve as model size increases, indicating that negation understanding benefits from larger models. The improvement is more pronounced in the option-selection evaluation than in the completion-based setting, with larger models showing higher accuracy gains (as seen in Figure 3). This suggests that option selection benefits more from larger models, since it requires comparing multiple candidate answers rather than generating a single continuation.
Figure 3: Zero-shot accuracy of instruction-tuned Qwen3 models under the detailed prompt setting.
Effect of Instruction Tuning.
Instruction tuning generally improves performance across most model families and evaluation settings. As shown in Figure 4, this improvement is observed in both completion-based and option-selection evaluations. We further observe that instruction tuning yields larger improvements under the detailed prompt setting than under the definition-based instruction. This suggests that instruction-tuned models follow step-by-step instructions more reliably, leading to larger gains in negation understanding.
Figure 4: Boxplots showing the distribution of zero-shot accuracy for pretrained and instruction-tuned models across evaluation settings, using both definition-based and detailed prompt instructions.
However, the effect of instruction tuning is not uniform across all models. In particular, some Qwen3 models exhibit comparable or higher performance in their pretrained versions, indicating that instruction tuning does not always guarantee gains for negation understanding.
Effect of Few-Shot Learning.
Overall, increasing the number of in-context demonstrations tends to improve performance across few-shot settings. This indicates that few-shot learning generally benefits sentence-level negation understanding. Figure 5 illustrates this trend for the Llama-3.1-8B-Instruct model under definition-based instruction. The initial introduction of demonstrations leads to a clear performance improvement, while the incremental benefits decrease as more examples are added.
This trend, however, is not uniform. For some API-based models under the detailed instruction, additional demonstrations do not always yield further gains. One possible reason is that the detailed instructions already provide strong guidance for API-based large models, and adding demonstrations may introduce noise or encourage unnecessary pattern following.
Figure 5: Mean accuracy of Llama-3.1-8B-Instruct under the definition instruction across different numbers of in-context demonstrations.
Evaluation Formats Instruction Formats Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
completion- based definition baseline zero-shot 0.536 73.82 23.96 2.22 25.00 31.82 58.50 48.71
5-shot 0.367 85.31 14.25 0.43 18.91 22.40 40.14 48.06
after SFT zero-shot 0.229 86.51 12.46 1.04 10.90 14.29 23.13 33.55
detailed baseline zero-shot 0.468 71.86 26.44 1.69 23.08 29.87 47.62 38.71
5-shot 0.347 84.70 14.84 0.46 19.55 23.05 37.76 41.29
after SFT zero-shot 0.211 87.97 10.90 1.13 8.97 13.31 21.09 33.23
option- selection definition baseline zero-shot 0.462 70.67 18.52 10.81 27.56 24.68 53.40 30.00
5-shot 0.332 81.62 11.22 7.16 25.64 18.83 40.48 27.42
after SFT zero-shot 0.356 71.94 15.37 12.69 25.00 19.81 44.56 17.10
detailed baseline zero-shot 0.486 71.45 14.85 13.70 33.01 25.97 56.12 29.03
5-shot 0.353 77.53 11.24 11.24 27.56 19.81 39.12 26.77
after SFT zero-shot 0.270 65.69 21.70 12.61 14.74 14.61 29.25 15.16
Table 6: Error distribution and confusion analysis of Llama-3.1-8B-Instruct model across various evaluation settings.
Effect of Supervised Fine-Tuning.
Supervised fine-tuning (SFT) generally leads to improved zero-shot performance across most models, indicating that task-specific training further enhances negation understanding. The performance gains from SFT are more pronounced in the completion-based evaluation than in the option-selection setting, suggesting that the effects of fine-tuning may interact with the evaluation format. We observe that SFT does not degrade models’ general language abilities, as confirmed by evaluations on general benchmarks (Appendix P).
Detailed results for all models and training settings are provided in Appendix O. In addition, we conducted a human evaluation, showing that although humans perform well on average, distinguishing standard negation from closely related alternatives is not uniformly easy across participants. Detailed results are provided in Appendix R.
5.3 Error Analysis
We analyze model errors to evaluate the ability of our models to differentiate standard negation from similar semantic variants. Each type of local negation in our dataset is explicitly labeled based on its sentence structure, as defined in Table 3.
We measure the confusion rate, defined as the proportion of examples within each subtype where the model incorrectly selects the local negation option instead of the correct standard negation. For example, if 320 items are labeled as participle clause negation and the model incorrectly chooses the local negation option instead of the correct standard negation option in 32 of these cases, the confusion rate for participle clause negation would be 10%. Complete analysis results are provided in Appendix Q.
We focus on the results of Llama-3.1-8B-Instruct model as shown in Table 6. In the completion-based setting, error rates generally decrease from zero-shot to 5-shot and continue to improve after SFT, with most errors concentrated in local negation options. Within local negation, compound sentences exhibit the highest confusion but also show the largest relative improvement after SFT.
In the option-selection setting, errors are likewise dominated by local negation, while the model shows a higher tendency to confuse paraphrase options compared to the completion-based setting. Comparatively, the option-selection setting results in lower confusion rates for adverbial clause negation. In addition, performance improvements in the option-selection setting do not consistently follow the expected progression across training configurations. In some cases, SFT yields higher error rates than the few-shot baselines, and this pattern is observed in some other models as well.
Overall, these patterns highlight how different evaluation settings and model configurations lead to distinct types of errors, and how the addition of more examples or SFT affects error distribution.
6 Conclusion
In this work, we introduce Thunder-NUBench, a benchmark designed to evaluate LLMs’ sentence-level understanding of negation, going beyond surface cue detection. By distinguishing between standard negation, local negation, contradiction, and paraphrase, Thunder-NUBench offers a comprehensive assessment of semantic comprehension. Our experiments demonstrate that while supervised fine-tuning and in-context learning can help reduce specific errors, these approaches still struggle to differentiate standard negation from closely related semantic variants. Thunder-NUBench serves as a valuable diagnostic tool for analyzing the limitations of models’ understanding of negation and stands as a robust benchmark for future research. Its design enables evaluation across diverse model families and settings, making it broadly applicable for studying semantic reasoning in LLMs.
Limitations
Thunder-NUBench is exclusively constructed in English, despite negation being a universal linguistic phenomenon demonstrated in diverse forms across languages. The syntactic and semantic expressions of negation may vary in other languages, meaning that our current findings may not generalize to multilingual or cross-lingual settings. In future work, we aim to extend the research to a broader range of languages to enable cross-linguistic evaluation of negation understanding in language models.
Although we built the dataset using two distinct sources (HoVer and Wikipedia summaries), both are derived from encyclopedic, formal domains, which may not fully represent the variety of sentence structures and informal language found in real-world use cases. Moreover, while all examples were systematically generated and reviewed, some bias may persist due to subjective decisions in the human review process. We attempt to mitigate this through cross-checking by an independent group of authors, but some residual bias may remain.
Thunder-NUBench primarily focuses on standard (sentence-level) negation and its distinction from local negation, contradiction, and paraphrase. Other important negation phenomena, such as double negation and negative polarity items (NPIs), are not directly addressed in this benchmark. Our current focus is on establishing a strong foundation for evaluating models’ understanding of standard negation. However, we aim to expand the evaluation to a broader range of negation phenomena in future work.
Ethical Considerations
This work does not involve the use of crowd-sourcing methods. Instead, all data included in the Thunder-NUBench benchmark has been carefully reviewed by the authors to ensure quality, relevance, and adherence to ethical standards. The datasets and tools used for training and evaluation are publicly available and used in compliance with their respective licenses.
When leveraging OpenAI’s text generation models, we take additional care to avoid generating or including any content that is harmful, biased, or violates privacy. All generated examples are manually reviewed to meet ethical and safety standards. We ensure no personally identifiable information or offensive content is present in the final dataset.
The Thunder-NUBench dataset is released under the CC BY-NC-SA 4.0 license, ensuring transparency, reproducibility, and accessibility for future research. We believe our work contributes positively to developing trustworthy and interpretable language models.
Acknowledgments
We thank the anonymous reviewers and the meta-reviewer for their valuable feedback on this paper. We also sincerely thank Sungeun Hahm, Suyoung Park, Jongmin Kim, Yelim Ahn, Hyunji M. Park, Seorin Kim, and Jisoo Kim for their valuable contributions to the initial design of the negation dataset.
This work was partially supported by the National Research Foundation of Korea (NRF) under Grant No. RS-2023-00222663 (Center for Optimizing Hyperscale AI Models and Platforms), and by the Institute for Information and Communications Technology Promotion (IITP) under Grant No. 2018-0-00581 (CUDA Programming Environment for FPGA Clusters) and No. RS-2025-02304554 (Efficient and Scalable Framework for AI Heterogeneous Cluster Systems), all funded by the Ministry of Science and ICT (MSIT) of Korea. It was also partially supported by the Korea Health Industry Development Institute (KHIDI) under Grant No. RS-2025-25454559 (Frailty Risk Assessment and Intervention Leveraging Multimodal Intelligence for Networked Deployment in Community Care), funded by the Ministry of Health and Welfare (MOHW) of Korea. Additional support was provided by the BK21 Plus Program for Innovative Data Science Talent Education (Department of Data Science, Seoul National University, No. 5199990914569) and the BK21 FOUR Program for Intelligent Computing (Department of Computer Science and Engineering, Seoul National University, No. 4199990214639), both funded by the Ministry of Education (MOE) of Korea. This work was also partially supported by the Artificial Intelligence Industrial Convergence Cluster Development Project, funded by the MSIT and Gwangju Metropolitan City. Research facilities were provided by the Institute of Computer Technology (ICT) at Seoul National University.
Appendix A Typology of Contradiction
Contradiction Type Definition Example
Antonym
Contradiction caused by opposing
meanings of aligned words.
The policy was a success.
$\rightarrow$ The policy was a failure.
Negation
One sentence explicitly negates
a statement in the other.
She attended the meeting.
$\rightarrow$ She did not attend the meeting.
Numeric
Inconsistent numbers, dates, or
quantities in related statements.
Totally, ten people were injured.
$\rightarrow$ Totally, five people were injured.
Factive/Modal
Conflict in implied facts or modal
possibilities due to verbs or auxiliaries.
He managed to enter the building.
$\rightarrow$ He did not enter the building.
Structure
Syntactic rearrangement or argument
swapping causes contradiction.
Alice hired Bob.
$\rightarrow$ Bob hired Alice.
Lexical
Contradiction through incompatible
verbs or phrases, not strictly antonyms.
The manager praised her performance.
$\rightarrow$ The manager expressed
     disappointment in her performance.
World
Knowledge
Contradiction relies on common-sense
or background knowledge.
The Eiffel Tower is in Paris.
$\rightarrow$ The Eiffel Tower is in Berlin.
Table 7: Contradiction types from De Marneffe et al. (2008). Contradiction covers a broader scope than negation.
Contradictions in natural language can arise in diverse ways that go beyond simple negation. Following the typology of De Marneffe et al. (2008), contradictions can be grouped into seven categories: antonymy, explicit negation, numeric mismatch, factive/modal inconsistencies, structural reversals, lexical incompatibilities, and conflicts based on world knowledge. These categories reflect the fact that contradiction covers a broader semantic scope than negation alone. Table 7 summarizes these types with definitions and examples.
Appendix B Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (Hengeveld, 1986).
As discussed in Section 3.3, standard negation in this work targets the main predicate of a clause. For sentences with copular verbs, this means that the entire verb phrase, including the copular verb and its complement, is subject to negation. For example, in the sentence "She is a doctor," the main predicate is "is a doctor." Negating this sentence results in "She is not a doctor," where the negation applies to the entire predicate, not just the verb "is."
Negation of a verb phrase including a copular verb can be realized either syntactically (e.g., "is not an expert") or by replacing the complement with its complementary antonym (e.g., "is a non-expert"), both of which result in the reversal of the main predicate’s truth value. Although such constructions may superficially appear to be non-verbal negation, especially when the complement is a noun or adjective, they are, in fact, instances of verbal negation, since the negation applies to the predicate as a whole.
Appendix C Negation of Implications
Negating implications presents challenges, as natural language intuitions often diverge from the rules of formal logic. Let’s say there is a conditional statement, "If I study hard, I will pass the bar exam." Formally, let $P$ denote "I study hard" and $Q$ denote "I will pass the bar exam." In classical logic, the conditional "if $P$, QQ" can be false only when $P$ is true and $Q$ is false. This implies that the negation of the conditional is $P\;\text{and}\;\text{Neg}(Q)$" ("I study hard and I won’t pass the bar exam,) while the conditional itself is equivalent to $\text{Neg}(P)\;\text{or}\;Q$" ("I don’t study hard or I will pass the exam) (Nguyen et al., 2023).
Psychological studies confirm that people often accept both "if $P$, Neg(QQ)" ("If I study hard, I won’t pass the exam.") and "if Neg($P$), QQ" ("If I don’t study hard, I will pass the exam."). However, the former can be interpreted as $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$", and the latter $P\;\text{or}\;Q$", both of which are not equivalent to the original statement’s negation, $P\;\text{and}\;\text{Neg}(Q)$". "if Neg($P$), Neg(QQ)" ("If I don’t study hard, I won’t pass the exam.") is not the correct negation as well, as it is equivalent to $P\;\text{or}\;\text{Neg}(Q)$(Espino and Byrne, 2012).
While humans often struggle to distinguish the correct negation of a conditional from invalid alternatives, the logical form is unambiguous. We therefore include conditional statements in our benchmark to test whether language models, like humans, are prone to intuitive but invalid interpretations, or whether they can correctly apply truth-functional reasoning.
Note that implications in natural language are not limited to the explicit "if $P$, QQ" form, but may also appear with connectives such as when, as long as, or unless, which functionally convey conditional meaning and are treated under the same negation principle.
Appendix D Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (Gleitman, 1965).
Coordinating conjunctions connect elements of equal grammatical rank. The seven common ones in English are: for, and, nor, but, or, yet, so (often remembered as FANBOYS). Among these, and, or, and but are indisputably used to coordinate clauses. The others can be ambiguous or function in non-coordinating roles(e.g., indicating cause or result rather than logical structure). These are the examples using and, or, and but to connect sentences equally.
  • "She studied hard, and she passed the exam."
  • "I wanted to go, but it was raining."
  • "You can call me, or you can send an email."
We consider only the coordinating conjunctions and, or, and but as indicators of compound sentences, in which two or more independent clauses are equally connected. Although but introduces a contrast semantically, in terms of logical structure, it functions as a conjunction equivalent to and; therefore, its negation follows the same principle.
Appendix E Local Negation Constructions Excluded
In constructing the Thunder-NUBench dataset, we consider various types of local (i.e., subclausal) negation, where negation applies to a phrase or constituent rather than the main predicate. However, several constructions are excluded due to their semantic ambiguity, syntactic irregularity, or misalignment with the benchmark’s focus on verbal negation.
Infinitive Phrase Negation.
Infinitive phrases (e.g., "to go") can be negated with "not" (e.g., "not to go" or "to not go"). Unlike the clause-level structures that define our local negation category, infinitive phrases are not full clauses but simply part of a verb phrase, making them less compatible with our definition. Moreover, although grammatically correct, this construction is relatively rare and sounds awkward depending on the context.
  • Original: George wants to go to the park.
  • Negated (infinitive): George wants not to go / George wants to not go to the park.
For these reasons, we exclude infinitive phrase negation from the benchmark.
Appositive Clause Negation.
Appositive clauses are noun phrases that provide descriptive clarification. Attempting to negate an appositive typically involves lexical replacement rather than syntactic negation.
  • Original: My brother, a talented musician, plays the guitar.
  • Negated (appositive): My brother, not a talented musician, plays the guitar.
Such changes alter descriptive content rather than reversing the meaning of the predicate, and often fall into the domain of contradiction. Accordingly, they are excluded from the dataset.
Prepositional Phrase Negation.
Negating a prepositional phrase often involves replacing the preposition with its antonym (e.g., "with" → "without", "in" → "outside"), which results in a sentence that differs in content, rather than reversing the meaning of the predicate.
  • Original: She went to the park with her bird.
  • Negated (preposition): She went to the park without her bird.
Since such modifications do not negate the verb but instead change the nature of an adjunct or argument, they fall outside the scope of standard negation or local negation in this work and are excluded.
In all of the above cases, the negation does not target the whole verb phrase but rather peripheral elements within the sentence. As the Thunder-NUBench is designed to evaluate verbal negation, these local or phrase-level forms of negation were intentionally left out.
Appendix F Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (Déprez et al., 2015).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(Van der Wouden, 1996). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we have generated paraphrase candidates for our dataset using such double negation patterns. For example,
  • Sentence: His characteristic style fuses samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    Double Negation: His characteristic style does not fail to fuse samba, funk, rock, and bossa nova with lyrics that blend humor and satire with often esoteric topics.
  • Sentence: It covers a broad range of fields, including the humanities, social sciences, exact sciences, applied sciences, and life sciences.
    Double Negation: It does not exclude a broad range of fields, including the humanities, social sciences, exact sciences, applied sciences, and life sciences.
  • Sentence: Sanders was honoured to meet with many world dignitaries and representatives of UNESCO member nations, and delighted when delegates from UNESCO, visited Toowoomba in 2018 in return.
    Double Negation: Sanders was not unhappy to meet with many world dignitaries and representatives of UNESCO member nations, and not displeased when delegates from UNESCO visited Toowoomba in 2018 in return.
However, upon closer examination, these paraphrase candidates do not always preserve the exact meaning of the original sentence. The antonyms used (e.g., "exclude" for "cover," "unhappy" for "honoured") are not always true complementary antonyms, which does not effectively negate the meaning. Moreover, the litotes construction ("does not fail to fuse") tends to add an emphatic nuance, rather than being a perfect semantic equivalent. Therefore, the boundary between paraphrasing and double negation is ambiguous, and their relationship requires more careful analysis. Given these issues, and because our primary focus is on standard negation, we ultimately decide to exclude double negation constructions as paraphrase candidates from our dataset.
Appendix G HoVer Dataset
Column Detail Example
id Unique claim identifier 0
uid
User/annotator identifier
330ca632-e83f-4011-b11b-0d0158145036
claim
The statement to be verified,
often requiring multi-article evidence
Skagen Painter Peder Severin Krøyer favored
naturalism along with Theodor Esbern Philipsen
and the artist Ossian Elgström studied
with in the early 1900s.
supporting_facts
List of Wikipedia article titles and sentence indices
providing evidence
[ { "key": "Kristian Zahrtmann", "value": 0 },
{ "key": "Kristian Zahrtmann", "value": 1 },
{ "key": "Peder Severin Krøyer", "value": 1 },
{ "key": "Ossian Elgström", "value": 2 } ]
label
Whether the claim is supported
1: SUPPORTED or 0: NOT_SUPPORTED
num_hops
Number of articles required for verification
[…]
hpqa_id
Reference to the original HotpotQA pair
5ab7a86d5542995dae37e986
Table 8: Details of HoVer dataset structure with examples.
The HoVer (Hoppy Verification) dataset is developed for the tasks of multi-hop evidence retrieval and factual claim verification. In HoVer, each claim requires supporting evidence that spans multiple English Wikipedia articles to determine whether the claim is substantiated or not. The dataset is distributed under a CC BY-SA 4.0 License, and it can be accessed via its official homepage [^2] https://hover-nlp.github.io/. Table 8 offers an overview of the dataset’s structure. The data is split into training, validation, and test sets, containing 18,171, 4,000, and 4,000 examples respectively.
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (Yang et al., 2018). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (Jiang et al., 2020).
Appendix H Wikipedia Summary Dataset
Column Detail Example
title Article title from Wikipedia. Alain Connes
description
A brief description or category for the article
(when available).
French mathematician
summary
The extracted summary or introduction section
of the article, typically more concise than the full text.
Alain Connes (; born 1 April 1947)
is a French mathematician…
full_text
The complete article text (when included),
encompassing the full body of the Wikipedia page.
Alain Connes (; born 1 April 1947)
is a French mathematician…
__index_level_0__ Index number for each entry in the dataset. 3
Table 9: Details of Wikipedia Summary dataset structure with examples.
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^3] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^5] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 9 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (Scheepers, 2017).
Appendix I Human Review Protocol
To ensure high-quality data construction, we implement a rigorous quality control protocol that combines generation, independent review, and iterative consensus building. The process involves the following key steps:
  • Task allocation and independence. Authors are assigned distinct portions of the dataset, but no author is permitted to review the data they have generated. This ensures that each instance is subject to at least one independent review.
  • Sequential authoring across choices. For the multiple-choice dataset, construction proceeds in four stages: standard negation, local negation, contradiction, and paraphrase. At each stage, different authors are responsible for creating the new option, while reviewers who have not authored that option perform the verification.
  • Cross-checking and layered review. Each newly created option is reviewed by at least one other author, and reviewers also revisit earlier options in the same instance. For example, when reviewing the paraphrased sentence, the reviewer also checks that standard negation, local negation, and contradiction sentences are correct. As a result, every instance undergoes multiple rounds of verification across stages, such that all authors ultimately examine data they have not created themselves.
  • Guideline refinement and retroactive correction. Generation and reviewing guidelines are continuously updated based on discussion of ambiguous or problematic cases. Whenever the guidelines changes, all previously created data are revisited to ensure compliance, promoting consistency across the dataset.
  • Consensus and adjudication. Disagreements are discussed in weekly meetings and, if necessary, adjudicated by a lead reviewer, ensuring that no instance remains unresolved.
Overall, this iterative and layered procedure ensures that every instance in the multiple-choice dataset is independently reviewed across multiple stages, leading to stable guidelines and a consistent dataset.
Edit-based reliability and quality control statistics.
Because our review process is revision-based rather than label-based, standard inter-annotator agreement metrics (e.g., Cohen’s $\kappa$) are not directly applicable. Instead of assigning categorical labels, reviewers have directly edited or deleted problematic instances to enforce the dataset guidelines.
We therefore report edit-based quality control statistics as a proxy for annotation reliability, including (i) the number of instances escalated to adjudication meetings and (ii) the number of instances that required edits after cross-checking. All cases requiring edits have been discussed in meetings and resolved by consensus, ensuring consistency across the dataset.
Subset # items
# Brought
to meeting
# Edited
Edit
rate
Sentence-
Negation
Standard
Negation
3,992 469 72 1.80%
Multiple Choice
Standard/
Local Negation
1,102 126 34 3.09%
Contradiction/
Paraphrase
1,102 409 117 10.62%
Implication
(add-on)
201 49 16 7.96%
Table 10: Edits include rewriting or deleting instances that violate guidelines.
Overall, the relatively low edit rates after cross-checking indicate that most instances satisfied the guidelines upon independent review, while meeting-based adjudication played a critical role in resolving ambiguous or complex cases.
Appendix J Detailed Principles and Examples of the Thunder-NUBench
While the main text already defines the core notions of standard and local negation (Section 3) and explains how they are applied throughout dataset construction (Section 4), here we provide more detailed illustrations.
Paraphrasing before Negation.
Before negating, the main verb or other components may be paraphrased with synonyms, provided that the sentence’s tense, structure, and meaning remain strictly equivalent before applying standard negation. Authors refer to the Merriam-Webster Thesaurus  [^6] https://www.merriam-webster.com/. For example,
  • Original Sentence: Toumour is a village and rural commune in Niger located near the Niger–Nigeria border.
    • Paraphrased Sentence: Toumour is a village and rural commune in Niger that is found close to the Niger–Nigeria boundary.
      $\rightarrow$ Standard Negation after Paraphrase: Toumour isn’t a village and rural commune in Niger that is found close to the Niger–Nigeria boundary.
    • Explanation: In this example, the participle clause "located near the Niger–Nigeria border" is rephrased as a relative clause "that is found close to the Niger–Nigeria boundary." Since both constructions serve as modifiers and preserve the same semantic role, we treat them as equivalent in meaning for the purpose of standard negation.
  • Original Sentence: The armed forces said Boko Haram attacked their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
    • Paraphrased Sentence: The armed forces stated that Boko Haram assaulted their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
      $\rightarrow$ Standard Negation after Paraphrase: The armed forces didn’t state that Boko Haram assaulted their military post on March 15, 2020, which they responded to by repelling the attack, killing 50 insurgents.
    • Explanation: In this example, the reporting verb “said” is paraphrased as “stated,” and the verb “attacked” is replaced with the synonym “assaulted.” These substitutions preserve the original tense and meaning, allowing standard negation to be applied without altering the semantic content of the sentence.
Negation of Simple Sentences.
For simple, declarative sentences, standard negation is achieved by inserting "not" after the auxiliary or main verb, or by replacing the predicate with its complementary antonym. For example, "She is happy." $\rightarrow$ "She is not happy."; "The room is occupied." $\rightarrow$ "The room is unoccupied."
Negation in Compound Sentences.
When multiple clauses or propositions are coordinated (e.g., with "and", "or", "but"), standard negation is logically applied, governed by De Morgan’s laws. Here, "but" is treated as a coordinating conjunction equivalent to "and" in terms of logical structure, so its negation follows the same principle.
  • Conjunction "$P$ and/but QQ": the negation is $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$".
  • Disjunction "$P$ or QQ": the negation is $\text{Neg}(P)\;\text{and}\;\text{Neg}(Q)$".
For example, "He passed the test and received an award." is negated as "He did not pass the test or did not receive an award."
When application of logical negation produces unnatural language, sentences may be split or slightly rephrased for fluency, provided logical meaning is preserved. For example,
  • Original: "He finished the report and submitted the assignment."
  • Standard Negation: "He did not finish the report or did not submit the assignment."
  • Standard Negation, but Splitted: "He did not finish the report. Or, he did not submit the assignment."
Coordinated Elements in the Sentence.
When a sentence contains coordinated elements (such as subjects, objects, or predicates connected by "and" or "or"), standard negation typically follows logical principles derived from De Morgan’s Laws. However, whether logical negation applies to each individual component or to the entire predicate as a whole depends on whether the coordination expresses multiple independent propositions or a single collective event.
  • If the coordination introduces semantically distinct propositions, that is, each conjunct could form a complete sentence on its own, negation must be applied to each proposition individually. For example, "My sister and I studied hard."
    This sentence can be interpreted as: "My sister studied hard and I studied hard."
    Therefore, the correct standard negation is: "My sister did not study hard, or I did not study hard."
  • Conversely, if the coordination connects elements that jointly participate in a single action or state (e.g., a shared subject or a collective predicate), then the sentence is treated as a simple clause, and the predicate as a whole is negated. Logical decomposition is not appropriate. For example, "My sister and I share clothes."
    This expresses a single collective action involving both participants.
    Therefore, the correct standard negation is: "My sister and I do not share clothes."
    (NOT: "My sister does not share clothes, or I do not share clothes.")
  • This distinction is crucial: even if two noun phrases are coordinated, if the sentence semantically decomposes into separate atomic propositions, standard negation must apply to each atomic proposition. Otherwise, it applies to the whole predicate as one unit.
  • Other examples of semantically collective predicates where logical splitting is not appropriate include: "be the same", "have in common", "do something together", "combine", "unite", etc. These describe inherently joint or relational properties, not independent propositions. For example, "Clarence Brown and Peter Glenville are from the same country." should be negated as "Clarence Brown and Peter Glenville are not from the same country."
Use of Antonyms.
When replacing predicates with antonyms in standard negation, only complementary antonyms are appropriate, as they provide a clear binary opposition, ensuring logical consistency of negation. Gradable and relational antonyms are unsuitable for standard negation because their antonyms do not represent the logical complement of the original predicate. In other words, replacing a predicate $p$ with its antonym does not produce $\neg p$ in a truth-conditional sense.
Specifically, unlike complementary antonyms, which form mutually exclusive pairs (i.e., $p\cup\neg p=U$ and $p\cap\neg p=\emptyset$), gradable and relational antonyms do not partition the meaning space cleanly, and thus fail to reverse the truth value reliably.
  • Complementary Antonyms: Also called binary/contradictory antonyms. These antonyms represent mutually exclusive pairs with no intermediate states. The presence of one implies the absence of the other. Examples include:
    • alive / dead
    • true / false
    • present / absent
    • occupied / vacant
    Using complementary antonyms in negation ensures a direct and unambiguous reversal of the original proposition’s truth value.
  • Gradable Antonyms: These antonyms exist on a continuum and allow for varying degrees between the two extremes. Negating one does not necessarily affirm the other. Examples include:
    • hot / cold
    • happy / sad
    • tall / short
    • young / old
    Due to their scalar nature, gradable antonyms are inappropriate for standard negation, as they do not provide a definitive binary opposition.
  • Relational Antonyms: Also known as converse antonyms, these pairs describe a reciprocal relationship where one implies the existence of the other. Examples include:
    • parent / child
    • teacher / student
    • buy / sell
    • employer / employee
    Relational antonyms are context-dependent and do not represent direct opposites in a binary sense, making them unsuitable for standard negation purposes.
General Principles of Standard Negation.
  • The negated sentence must preserve all elements (subject, tense, objects, adjuncts, etc.) of the original, except for the truth value of the main predicate.
  • When naturalness and logical negation conflict, logical correctness takes priority, but minimal rephrasing is allowed for fluency.
  • If the negated clause creates a contradiction with other parts of the sentence, the contradictory clause must be removed. For example, the standard negation of the sentence "While the spatial size of the entire universe is unknown, it is possible to measure the size of the observable universe, which is approximately 93 billion light-years in diameter."
    will be "While the spatial size of the entire universe is unknown, it isn’t possible to measure the size of the observable universe."
    The relative clause must be removed because its content directly contradicts the negated main clause.
Common Negation Errors and Corrections.
  • Original sentence: His characteristic style fuses samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Incorrect negation: His distinctive style doesn’t fuse samba, funk, rock or bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Correct negation: His distinctive style doesn’t fuse samba, funk, rock and bossa nova with lyrics that blend humor and satire with often esoteric subject matter.
    • Explanation: The verb "fuse" implies a combination of all listed elements. "and" must be preserved.
  • Original sentence: The mascot of Avon Center School is the "Koalaty Kid," while the mascot at Prairieview is an eagle and the mascot at Woodview is an owl.
    • Incorrect negation: Avon Center School’s mascot is not the "Koalaty Kid," Prairieview’s mascot is not an eagle, or Woodview’s mascot is not an owl.
    • Correct negation: Avon Center School’s mascot is not the "Koalaty Kid," while the mascot at Prairieview is an eagle and the mascot at Woodview is an owl.
    • Explanation: Two clauses connected by while are not coordinated propositions (as with and or or), but instead express contrastive information. Therefore, applying logical negation across both clauses is incorrect. Negation should apply only to the main clause (here, the first statement), while the contrasting clause remains affirmative.
Appendix K Code for Data Construction
K.1 Sentence-Negation Pair Dataset
To construct the sentence-negation pair dataset, we begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI, 2025) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective.
We select different model versions depending on the complexity of each task. For sentence merging, which demands nuanced contextual understanding and complex syntax, we use GPT-4. For grammar correction, where edits are more straightforward, GPT-3.5 is sufficient.
1def grammar_fix(claim):
2 messages = [{"role": "system", "content": "Fix grammatical errors."},
[…]
4 response = client.chat.completions.create(model="gpt-3.5-turbo", messages=messages)
5 fixed_text = response.choices[0].message.content.strip()
6 return fixed_text
Listing 1: Fixing Grammar with OpenAI API.
1def merge_sentences_with_gpt(claim):
2 messages = [{"role": "system", "content": "Merge sentences into a single one."},
[…]
4 response = client.chat.completions.create(model="gpt-4-turbo-preview", messages=messages)
5 merged_text = response.choices[0].message.content.strip()
6 return merged_text
Listing 2: Merging Sentences with OpenAI API.
K.2 Multiple Choice Dataset
To construct the multiple-choice dataset, we first segment the "summary" column of the Wikipedia Summary dataset, which often contains multiple sentences in a single entry, into individual sentences. To focus on the challenges of negation in complex sentences, we filter out sentences that are too short. This process is done with Python code.
Since conditional sentences (e.g., "If $P$, QQ") are rarely present in the Wikipedia summary dataset, we adopt a two-step approach: (1) prompting the model to generate conditional variants from given sentences (using OpenAI API, GPT-4o-mini), and (2) manually filtering or lightly editing the results to obtain valid conditionals.
Subsequently, we automatically generate contradictions and paraphrases for each sentence via the OpenAI API (GPT-4o) as well, followed by human review. The following scripts illustrate the procedures.
1import pandas as pd
2import re
3from datasets import load_dataset
4import random
5
6df = pd.DataFrame(load_dataset("jordiclive/wikipedia-summary-dataset")[’train’].shuffle(seed=42).select(range(10000))
7df = df.drop(columns=[’full_text’])
8
9def split_into_sentences(text):
10 sentences = re.split(r’(?<=[.!?]) +’, text)
11 return sentences
12
13df[’sentence’] = df[’summary’].apply(split_into_sentences)
14df = df.explode(’sentence’)
15df = $df[$’sentence’].apply(lambda x: len(x.split()) >= 30)]
16df = df.reset_index(drop=True)
17df.to_csv("file/wikipedia_summary_sentences.csv", index=False)
Listing 3: Sentence extraction and preprocessing from Wikipedia summaries.
1def generate_conditionals(sentence):
2 prompt = f"""
3 Based on the sentence below, write a conditional sentence that uses the main topic of the sentence.
4 The conditional sentence should express a hypothetical situation or cause-effect relationship related to the topic. It can be slightly complex in structure.
5 For example:
6 - If it rains tomorrow, I will stay home.
7 Sentence:
8 ’{sentence}’
9 """
10
11 completion = client.chat.completions.create(
12 model="gpt-4o-mini",
13 messages=[
14 {"role": "system", "content": "You are a helpful assistant that specializes in generating conditional sentences."},
15 {"role": "user", "content": prompt}
16 ]
17 )
18
19 return completion.choices[0].message.content
Listing 4: Conditionals sentence generation.
1def generate_contradiction(sentence):
2 prompt = f"""
3 You will be given a sentence. Generate a contradictory sentence that directly conflicts with the original sentence without using standard negation.
4
5 Definitions:
6 - Standard negation: Directly negating the main verb or using words like ’not’, ’no’, ’never’, or negative contractions such as \"isn’t\", \"doesn’t\", or \"can’t\".
7 - Contradiction: A sentence that logically conflicts with the original statement. The contradiction must be such that both sentences cannot logically be true at the same time under any circumstances.
8
9 Important:
10 - Do not change the main verb from the original sentence.
11 - Do not use ’never’ or other negative words to form the contradiction.
12 - Ensure the contradicted sentence logically excludes the possibility of the original sentence being true simultaneously.
13
14 Examples:
15 Original sentence: \"The tallest student won the award.\"
16 Contradicted sentence: \"The shortest student won the award.\"
17
18 Original sentence: \"The room was completely dark.\"
19 Contradicted sentence: \"The room was brightly lit.\"
20
21 Original sentence: \"The event took place in the morning.\"
22 Contradicted sentence: \"The event took place in the evening.\"
23
24 Original sentence: \"All people are dying.\"
25 Contradicted sentence: \"Some people are dying.\"
26
27 Now, generate a contradictory sentence without standard negation, without changing the main verb, and ensuring the two sentences are logically incompatible, for the following:
28
29 Original sentence: \"{sentence}\"
30
31 Contradicted sentence:
32 """
33
34 completion = client.chat.completions.create(
35 model="gpt-4o",
36 messages=[
37 {"role": "system", "content": "You are a helpful assistant tasked with generating logical contradictions. Do not use negation to make contradiction."},
38 {"role": "user", "content": prompt}
39 ]
40 )
41
42 return completion.choices[0].message.content
Listing 5: Contradiction generation.
1def generate_paraphrase(sentence):
2 prompt = f"""
3 Paraphrase the following sentence using synonyms or slight structural variations without changing its meaning.
4 Do not add or remove any main verbs. Keep the original intent of the sentence intact.
5
6 Original sentence: "{sentence}"
7
8 Paraphrased sentence:
9 """
10
11 completion = client.chat.completions.create(
12 model="gpt-4o",
13 messages=[
14 {"role": "system", "content": "You are a helpful assistant skilled at generating paraphrases while keeping the meaning of sentences unchanged."},
15 {
16 "role": "user",
17 "content": prompt
18 }
19 ]
20 )
21
22 return completion.choices[0].message.content
Listing 6: Paraphrase generation.
choice2_type Definition
Demonstration Set
Test Set
relative_part
negation inside relative clauses
(e.g., "who did not attend…").
12 312
pp_part
negation in participle clauses
(e.g., "not walking through the park…").
12 308
adverb_part
negation in adverbial clauses
(e.g., "because it was not raining").
12 310
compound_part negation applied to one clause within a compound sentence. 12 294
non-applicable
used when the sentence structure does not support
a valid local negation variant under our definition.
2 37
Total 50 1,261
Table 11: Choice 2 Types and Distributions.
Appendix L Thunder-NUBench Dataset Structure
Thunder-NUBench consists of two subsets: a sentence-negation pair dataset for supervised fine-tuning and a multiple-choice dataset for evaluation. Both datasets are built on English text and reviewed by authors following strict guidelines.
L.1 Sentence-Negation Pair Dataset
This subset contains pairs of affirmative and corresponding standard negation sentences. It includes the following fields:
  • index: the index of the data.
  • premise: the original sentence.
  • hypothesis: its logically negated form.
L.2 Multiple-Choice Dataset
This evaluation set presents each original sentence with four candidate transformations.
  • wikipedia_index: the original index of the Wikipedia Summary dataset.
  • index: the index of the data.
  • sentence: the original sentence.
  • choice1: standard negation (correct answer).
  • choice2: local negation (subclausal negation).
  • choice2_type: specifies the type of local negation.
  • choice2_element: a short description of the phrase or clause that was negated (e.g., "being built", "which crashed").
  • choice3: contradiction (non-negated, semantically incompatible).
  • choice4: paraphrase (semantically equivalent).
The details of choice2_type and distribution on demonstration and test sets are described in Table 11. It follows the definition in Table 3.
In addition to the Wikipedia Summary dataset, we supplement the evaluation set with conditionals (e.g., If $P$, $Q$) by manually searching Wikipedia articles where such constructions are more likely to occur (e.g., Newton’s laws of motion). Among the 100 conditional sentences included across demonstration and test sets, 20 are collected through manual search (marked with indices beginning with "S" in wikipedia_index). Meanwhile, the remaining 80 are sampled from the Wikipedia Summary dataset and converted into conditional form using the script in 4 (Appendix K).
Appendix M Prompt Templates Used in Evaluation
We design and employ two instruction formats for in-context learning. These instructions differ in the level of guidance they provide, ranging from a task definition to a detailed procedural instruction. Specifically, we use the following two instruction styles. These instruction formats correspond to the task instruction component shown in Figure 1.
  1. 1.
    Definition instruction: A concise definition of standard negation, which specifies the target operation,
    "Standard negation is sentential negation that reverses the truth value of the sentence by negating the main predicate(s) of the main clause(s). Keep the rest of the sentence content unchanged."
  2. 2.
    Detailed instruction: A step-by-step instruction that presents standard negation as an explicit procedure. This prompt describes how to identify the main clause and its main predicate, preserve other sentence elements, apply syntactic negation or complementary antonyms where appropriate, and negate complex sentences by reversing their logical structure. The full prompt format is shown in Listing 7.
Standard negation reverses the truth value of the main predicate in the main clause while keeping all other elements of the main clause unchanged.
Do not negate subordinate clauses or modify other parts of the sentence.
To do this:
1) Identify the main clause and its main verb (main predicate). Ignore subordinate clauses.
2) Preserve all other main-clause content.
3) Insert a negative particle such as "not" into the main verb, or replace it with a complementary antonym only if it forms an absolute binary (e.g., alive/dead, true/false, possible/impossible).
4) If the sentence contains multiple propositions connected by logical operators (e.g., and, or, conditional constructions), negate it in a way that reverses the entire proposition (e.g., A and B -> not A or not B; If A then B -> A and not B).
Listing 7: Detailed instruction format.
Below, we show the prompt formats used for completion-based and option-selection evaluations, including prompt–response structures for each evaluation setting.
Prompt] "{instruction format}
Generate the standard negation of the given sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
Negation:"
Response] "Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells."
Listing 8: Completion-based Format.
Prompt] "Given the following instruction and candidate answers, choose the single best answer.
{instruction format}
Generate the standard negation of the given sentence.
Sentence: Chromosome 2 is the second-largest human chromosome, spanning more than 242 million base pairs and representing almost eight percent of the total DNA in human cells.
A. Chromosome 2 isnt the second-largest human chromosome, which measures more than 242 million base pairs and represents almost eight percent of the entire DNA in human cells.
B. Chromosome 2 is the second-largest human chromosome, which doesnt span more than 242 million base pairs or represent nearly eight percent of the whole DNA in human cells.
C. Chromosome 2 is the smallest human chromosome, spanning fewer than 50 million base pairs and representing less than two percent of the total DNA in human cells.
D. Chromosome 2 is the second-biggest human chromosome, with over 242 million base pairs, making up nearly 8% of all DNA in human cells.
Your response should be one of A, B, C, D.
Only output the letter.
Answer:"
Response] "A"
Listing 9: Option-selection Format.
Appendix N Evaluation Setup
We use 8 NVIDIA GeForce RTX 3090 GPUs with 24GB of memory and CUDA 12.4 for all fine-tuning and evaluation.
N.1 Models
We evaluate a diverse set of pretrained and instruction-tuned models across multiple model families and scales, ranging from about a billion to 14 billion parameters. Our experiments include models from Gemma (Team et al., 2024), Qwen (Yang et al., 2025), LLaMA (Grattafiori et al., 2024), and Mistral families (Jiang et al., 2023), as well as API-based models from OpenAI (Achiam et al., 2023; Hurst et al., 2024) and Anthropic (Anthropic, 2024). API models are evaluated only in the option-selection setting, as they do not provide access to token-level log probabilities required for completion-based evaluation.
  • Gemma models:
    • gemma-2-2b, gemma-2-2b-it
    • gemma-2-9b, gemma-2-9b-it
  • Qwen models:
    • Qwen3-0.6B-Base, Qwen3-0.6B
    • Qwen3-1.7B-Base, Qwen3-1.7B
    • Qwen3-4B-Base, Qwen3-4B
    • Qwen3-8B-Base, Qwen3-8B
    • Qwen3-14B-Base, Qwen3-14B
    • Qwen3-32B
  • LLaMA models:
    • Llama-3.1-8B, Llama-3.1-8B-Instruct
    • Llama-3.2-1B, Llama-3.2-1B-Instruct
    • Llama-3.2-3B, Llama-3.2-3B-Instruct
  • Mistral models:
    • Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3
    • Mistral-Nemo-Base-2407 (12B size), Mistral-Nemo-Instruct-2407 (12B size)
    • Mistral-Small-24B-Base-2501, Mistral-Small-24B-Instruct-2501
  • API models:
    • GPT: GPT-4o mini, GPT-4o, GPT-4.1 mini, GPT-4.1
    • Claude: Haiku 4.5, Sonnet 4.5
N.2 Zero-shot and few-shot evaluation
For each model, we evaluate performance in both zero-shot and few-shot settings using the Language Model Evaluation Harness (Gao et al., 2024). In the few-shot scenario, we use examples from the demonstration set as in-context demonstrations. Results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and are reported for one, five, and ten examples from the demonstration set (1-shot, 5-shot, and 10-shot). We present the performance results on the test set for each model and prompt configuration.
N.3 Supervised fine-tuning
We conduct Supervised Fine-Tuning (SFT) using the LLaMA-Factory framework (Zheng et al., 2024) on the Sentence-Negation Pair dataset from Thunder-NUBench. The dataset is formatted in the Alpaca instruction style (Taori et al., 2023). To achieve parameter-efficient training, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 8, targeting all linear layers. The fine-tuning process is carried out for three epochs, using a batch size of 1, a gradient accumulation step of 8, cosine learning rate scheduling, and bfloat16 precision. We apply supervised fine-tuning only to models with fewer than 10B parameters, in order to focus on analyzing the effects of SFT on relatively smaller models, where instruction tuning is expected to have a more pronounced impact. After the SFT, we evaluate the model’s zero-shot performance to directly assess its ability to generalize from instruction tuning without being influenced by in-context examples. It is important to note that API models are not fine-tuned, as they are not compatible with the LLaMA-Factory framework.
Appendix O Total Model Performance on Thunder-NUBench
This section provides the full set of results on Thunder-NUBench for all evaluated models and training settings. Following the prompt templates in Appendix M, we report results separately for the definition instruction and detailed instruction styles.
For each instruction style, models are evaluated under three conditions: (1) zero-shot, (2) few-shot with 1, 5, and 10 in-context demonstrations (averaged across random seeds), and (3) supervised fine-tuning (SFT). For SFT, models are evaluated in the zero-shot setting to isolate the effect of task-specific training without the influence of in-context examples.
Tables 14 and 15 report the results of open-weight models under the definition and detailed instruction styles, respectively. Tables 16 and 17 present the corresponding results for API models under the same instruction settings.
Appendix P General Benchmark Performance after SFT
To assess whether supervised fine-tuning (SFT) on Thunder-NUBench affects performance on broader natural language understanding tasks, we evaluate models on six widely used benchmarks: ARC-Challenge, ARC-Easy, GSM8K, HellaSwag, OpenBookQA, and WinoGrande (Clark et al., 2018; Cobbe et al., 2021; Zellers et al., 2019; Mihaylov et al., 2018; Sakaguchi et al., 2021). These datasets cover a diverse range of domains, including commonsense reasoning, scientific knowledge, mathematics, and reading comprehension.
Table 18 and Table 19 summarize the results. We find that performance on general benchmarks remains broadly stable after SFT, indicating that fine-tuning on Thunder-NUBench does not substantially harm general capabilities.
Appendix Q Total Error Analysis of Model Predictions on the Thunder-NUBench
This section extends the error analysis presented in Section 5.3, providing the complete results. In particular, we examine (i) incorrect choice distributions and (ii) confusion rates for local negation categories, comparing models of different sizes and training paradigms (pretrained vs. instruction-tuned) under both zero-shot and few-shot conditions.
definition detailed
completion option completion option
Gemma2 Table 20 Table 21 Table 22 Table 23
Qwen3 Table 24 Table 25 Table 26 Table 27
Llama3 Table 28 Table 29 Table 30 Table 31
Mistral Table 32 Table 33 Table 34 Table 35
API Table 36 Table 37
Table 12: Complete prediction analysis by model family.
We report few-shot results using a fixed random seed (1234). Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make confusion analysis less interpretable. Each table follows the same instruction format, reporting error rates, incorrect choice distributions, and confusion rates across local negation subcategories under zero-shot, few-shot, and SFT conditions.
For API models, we also track cases labeled as “Answer Format Wrong.” This category captures instances where the model’s output does not follow the required answer format (selecting strictly one of A, B, C, or D). Because such responses cannot be mapped to a specific incorrect option, they are not included in the incorrect choice distribution or confusion rate. Instead, we report their raw counts alongside the other error statistics. This also serves as an indicator of the model’s ability to follow output-format instructions.
Appendix R Human Evaluation
The human evaluation was approved by the Institutional Review Board (IRB No. 2512/004-015). All participants provided informed consent through an official IRB-approved consent form prior to participation and were fully informed of the study’s purpose, procedures, potential risks and benefits, and their right to withdraw at any time without penalty. Participants were also notified that only aggregate statistics of their evaluation scores would be reported.
All responses were collected anonymously, and no personally identifiable information was included in the dataset. Participants were compensated KRW 50,000 for a two-hour session, which exceeds the minimum hourly wage (KRW 10,320 per hour) and constitutes adequate compensation. To minimize fatigue and ensure reliable responses, participants were given sufficient rest during the evaluation, and the total duration was limited to approximately two hours per session. Participants were required to be native or highly proficient English users (CEFR [^7] Common European Framework of Reference for languages. C-level or above), as verified during recruitment. A total of 11 adult participants aged in their 20s to 30s took part in the study, including 6 highly proficient English users at the CEFR C level and 5 native or dominant English speakers.
A total of 50 items were randomly sampled from the multiple-choice evaluation split of Thunder-NUBench, with stratified sampling to preserve the distribution of distractor types. Specifically, the sample included 12 items each for relative clause, participle clause, compound sentence, and adverbial clause (local negation type), and 2 items without local negation. All items were presented using the detailed instruction format. Participants were asked to select the option corresponding to standard negation for each item. No example items were provided during the evaluation.
Mean Median Min Max SD
0.78 0.82 0.58 0.96 0.139
Table 13: Human evaluation results.
Table 13 summarizes the results of the human evaluation conducted with 11 participants. Human accuracy ranges from 0.58 to 0.96, with a mean of 0.78. This range indicates that distinguishing standard negation from closely related alternatives can be challenging even for human readers, consistent with prior observations that negation processing requires careful semantic and syntactic reasoning. The average human accuracy is higher than the zero-shot performance of most models reported in this paper, suggesting that humans generally handle sentence-level negation more reliably without task-specific training.
zero-shot 1-shot 5-shot 10-shot after SFT
evaluation setting completion option
completion
(±SD)
option
(±SD)
completion
(±SD)
option
(±SD)
completion
(±SD)
option
(±SD)
completion option
gemma-2-2b 0.436 0.236
0.480
(±0.006)
0.328
(±0.004)
0.565
(±0.010)
0.373
(±0.009)
0.599
(±0.004)
0.398
(±0.011)
0.781 0.260
gemma-2-2b-it 0.516 0.504
0.514
(±0.008)
0.501
(±0.008)
0.569
(±0.007)
0.510
(±0.012)
0.594
(±0.003)
0.540
(±0.007)
0.741 0.543
gemma-2-9b 0.470 0.468
0.499
(±0.005)
0.530
(±0.009)
0.565
(±0.007)
0.588
(±0.006)
0.597
(±0.005)
0.626
(±0.005)
0.771 0.751
gemma-2-9b-it 0.508 0.554
0.562
(±0.007)
0.631
(±0.005)
0.653
(±0.007)
0.705
(±0.007)
0.689
(±0.005)
0.724
(±0.011)
0.767 0.814
Qwen3-0.6B-Base 0.453 0.349
0.505
(±0.008)
0.406
(±0.007)
0.581
(±0.005)
0.428
(±0.004)
0.616
(±0.005)
0.465
(±0.010)
0.708 0.479
Qwen3-0.6B 0.416 0.438
0.420
(±0.006)
0.466
(±0.004)
0.487
(±0.003)
0.502
(±0.006)
0.530
(±0.006)
0.528
(±0.007)
0.645 0.555
Qwen3-1.7B-Base 0.481 0.526
0.525
(±0.007)
0.559
(±0.007)
0.577
(±0.004)
0.641
(±0.004)
0.604
(±0.001)
0.676
(±0.007)
0.699 0.493
Qwen3-1.7B 0.398 0.532
0.406
(±0.006)
0.547
(±0.009)
0.476
(±0.007)
0.574
(±0.008)
0.514
(±0.005)
0.578
(±0.007)
0.668 0.608
Qwen3-4B-Base 0.454 0.602
0.489
(±0.002)
0.611
(±0.006)
0.535
(±0.004)
0.653
(±0.003)
0.560
(±0.007)
0.659
(±0.009)
0.733 0.709
Qwen3-4B 0.410 0.562
0.461
(±0.008)
0.619
(±0.005)
0.548
(±0.010)
0.676
(±0.005)
0.597
(±0.007)
0.708
(±0.006)
0.715 0.722
Qwen3-8B-Base 0.474 0.690
0.505
(±0.004)
0.692
(±0.003)
0.566
(±0.009)
0.741
(±0.008)
0.603
(±0.009)
0.758
(±0.007)
0.703 0.768
Qwen3-8B 0.442 0.589
0.487
(±0.011)
0.664
(±0.009)
0.556
(±0.008)
0.712
(±0.006)
0.597
(±0.006)
0.740
(±0.005)
0.722 0.784
Qwen3-14B-Base 0.489 0.708
0.547
(±0.006)
0.766
(±0.013)
0.613
(±0.005)
0.806
(±0.005)
0.653
(±0.007)
0.821
(±0.004)
Qwen3-14B 0.440 0.661
0.492
(±0.008)
0.677
(±0.009)
0.569
(±0.004)
0.732
(±0.007)
0.622
(±0.006)
0.751
(±0.003)
Qwen3-32B 0.470 0.730
0.534
(±0.011)
0.768
(±0.010)
0.650
(±0.009)
0.792
(±0.009)
0.690
(±0.006)
0.817
(±0.003)
Llama-3.1-8B 0.464 0.474
0.495
(±0.006)
0.511
(±0.005)
0.593
(±0.010)
0.580
(±0.011)
0.643
(±0.008)
0.629
(±0.010)
0.774 0.559
Llama-3.1-8B-
Instruct
0.464 0.538
0.540
(±0.005)
0.595
(±0.012)
0.628
(±0.006)
0.652
(±0.009)
0.668
(±0.006)
0.673
(±0.005)
0.771 0.644
Llama-3.2-1B 0.409 0.259
0.431
(±0.007)
0.263
(±0.011)
0.492
(±0.006)
0.250
(±0.010)
0.526
(±0.004)
0.262
(±0.009)
0.750 0.259
Llama-3.2-1B-
Instruct
0.431 0.301
0.468
(±0.006)
0.376
(±0.013)
0.526
(±0.007)
0.351
(±0.009)
0.548
(±0.009)
0.336
(±0.012)
0.761 0.286
Llama-3.2-3B 0.435 0.362
0.469
(±0.005)
0.443
(±0.004)
0.563
(±0.006)
0.472
(±0.003)
0.588
(±0.006)
0.483
(±0.004)
0.757 0.385
Llama-3.2-3B-
Instruct
0.452 0.512
0.456
(±0.007)
0.539
(±0.007)
0.548
(±0.005)
0.499
(±0.002)
0.590
(±0.004)
0.493
(±0.005)
0.795 0.651
Mistral-7B-v0.3 0.467 0.485
0.464
(±0.006)
0.494
(±0.015)
0.557
(±0.014)
0.631
(±0.007)
0.599
(±0.005)
0.653
(±0.007)
0.784 0.480
Mistral-7B-Instruct-
v0.3
0.611 0.657
0.643
(±0.006)
0.643
(±0.003)
0.694
(±0.005)
0.638
(±0.009)
0.715
(±0.008)
0.650
(±0.008)
0.793 0.692
Mistral-Nemo-
Base-2407 (12B)
0.460 0.412
0.490
(±0.010)
0.520
(±0.007)
0.600
(±0.003)
0.568
(±0.008)
0.649
(±0.007)
0.599
(±0.009)
Mistral-Nemo-
Instruct-2407 (12B)
0.487 0.604
0.528
(±0.004)
0.630
(±0.007)
0.633
(±0.009)
0.660
(±0.006)
0.677
(±0.007)
0.655
(±0.006)
Mistral-Small-24B-
Base-2501
0.474 0.568
0.500
(±0.003)
0.618
(±0.005)
0.597
(±0.006)
0.721
(±0.005)
0.647
(±0.007)
0.783
(±0.003)
Mistral-Small-24B-
Instruct-2501
0.526 0.686
0.569
(±0.006)
0.705
(±0.005)
0.637
(±0.009)
0.731
(±0.011)
0.688
(±0.006)
0.780
(±0.003)
total average 0.464 0.519 0.499 0.559 0.577 0.599 0.615 0.622 0.740 0.554
Table 14: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the definition instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
zero-shot 1-shot 5-shot 10-shot after SFT
evaluation setting completion option
completion
(±SD)
option
(±SD)
completion
(±SD)
option
(±SD)
completion
(±SD)
option
(±SD)
completion option
gemma-2-2b 0.438 0.276
0.494
(±0.005)
0.343
(±0.005)
0.570
(±0.004)
0.360
(±0.018)
0.604
(±0.004)
0.378
(±0.010)
0.768 0.259
gemma-2-2b-it 0.505 0.470
0.488
(±0.013)
0.474
(±0.009)
0.545
(±0.007)
0.487
(±0.011)
0.580
(±0.002)
0.500
(±0.009)
0.733 0.663
gemma-2-9b 0.469 0.466
0.513
(±0.005)
0.530
(±0.009)
0.570
(±0.006)
0.568
(±0.006)
0.604
(±0.005)
0.606
(±0.007)
0.772 0.797
gemma-2-9b-it 0.542 0.569
0.597
(±0.005)
0.635
(±0.005)
0.653
(±0.006)
0.695
(±0.010)
0.690
(±0.003)
0.715
(±0.009)
0.764 0.805
Qwen3-0.6B-Base 0.424 0.342
0.487
(±0.008)
0.397
(±0.004)
0.564
(±0.006)
0.416
(±0.008)
0.607
(±0.006)
0.461
(±0.010)
0.723 0.325
Qwen3-0.6B 0.413 0.459
0.411
(±0.004)
0.414
(±0.009)
0.487
(±0.003)
0.468
(±0.003)
0.535
(±0.007)
0.495
(±0.008)
0.679 0.526
Qwen3-1.7B-Base 0.462 0.483
0.511
(±0.007)
0.542
(±0.004)
0.574
(±0.005)
0.620
(±0.002)
0.597
(±0.005)
0.653
(±0.003)
0.724 0.609
Qwen3-1.7B 0.420 0.447
0.442
(±0.008)
0.502
(±0.013)
0.490
(±0.006)
0.547
(±0.007)
0.522
(±0.006)
0.556
(±0.010)
0.642 0.573
Qwen3-4B-Base 0.485 0.585
0.503
(±0.005)
0.581
(±0.006)
0.546
(±0.004)
0.636
(±0.006)
0.566
(±0.006)
0.649
(±0.009)
0.729 0.711
Qwen3-4B 0.439 0.577
0.489
(±0.010)
0.635
(±0.003)
0.559
(±0.008)
0.652
(±0.004)
0.600
(±0.005)
0.685
(±0.003)
0.688 0.711
Qwen3-8B-Base 0.455 0.668
0.513
(±0.004)
0.700
(±0.007)
0.557
(±0.005)
0.737
(±0.004)
0.591
(±0.009)
0.755
(±0.009)
0.714 0.823
Qwen3-8B 0.473 0.619
0.498
(±0.014)
0.695
(±0.006)
0.548
(±0.007)
0.693
(±0.007)
0.586
(±0.007)
0.723
(±0.007)
0.740 0.760
Qwen3-14B-Base 0.508 0.736
0.553
(±0.008)
0.771
(±0.010)
0.628
(±0.004)
0.797
(±0.007)
0.660
(±0.007)
0.816
(±0.006)
Qwen3-14B 0.488 0.720
0.522
(±0.007)
0.718
(±0.011)
0.601
(±0.008)
0.762
(±0.009)
0.636
(±0.005)
0.766
(±0.005)
Qwen3-32B 0.514 0.761
0.570
(±0.011)
0.814
(±0.004)
0.657
(±0.003)
0.808
(±0.001)
0.695
(±0.004)
0.821
(±0.005)
Llama-3.1-8B 0.451 0.431
0.486
(±0.006)
0.484
(±0.003)
0.564
(±0.008)
0.560
(±0.010)
0.620
(±0.004)
0.611
(±0.010)
0.802 0.672
Llama-3.1-8B-
Instruct
0.532 0.514
0.589
(±0.004)
0.588
(±0.010)
0.643
(±0.007)
0.626
(±0.013)
0.679
(±0.007)
0.648
(±0.004)
0.789 0.730
Llama-3.2-1B 0.403 0.259
0.439
(±0.009)
0.263
(±0.007)
0.497
(±0.007)
0.249
(±0.009)
0.527
(±0.005)
0.264
(±0.008)
0.747 0.259
Llama-3.2-1B-
Instruct
0.416 0.300
0.468
(±0.003)
0.347
(±0.009)
0.523
(±0.008)
0.339
(±0.010)
0.542
(±0.009)
0.330
(±0.009)
0.703 0.263
Llama-3.2-3B 0.447 0.374
0.478
(±0.003)
0.419
(±0.006)
0.568
(±0.006)
0.470
(±0.006)
0.590
(±0.004)
0.479
(±0.008)
0.761 0.364
Llama-3.2-3B-
Instruct
0.488 0.488
0.464
(±0.010)
0.518
(±0.014)
0.540
(±0.007)
0.492
(±0.005)
0.583
(±0.011)
0.481
(±0.006)
0.791 0.582
Mistral-7B-v0.3 0.479 0.502
0.469
(±0.005)
0.472
(±0.016)
0.555
(±0.012)
0.627
(±0.009)
0.597
(±0.006)
0.652
(±0.006)
0.792 0.477
Mistral-7B-Instruct-
v0.3
0.662 0.634
0.661
(±0.009)
0.643
(±0.009)
0.699
(±0.004)
0.635
(±0.008)
0.712
(±0.009)
0.643
(±0.008)
0.795 0.711
Mistral-Nemo-
Base-2407 (12B)
0.466 0.422
0.498
(±0.008)
0.522
(±0.015)
0.595
(±0.005)
0.557
(±0.006)
0.642
(±0.005)
0.593
(±0.007)
Mistral-Nemo-
Instruct-2407 (12B)
0.511 0.627
0.546
(±0.006)
0.643
(±0.005)
0.636
(±0.008)
0.658
(±0.005)
0.676
(±0.005)
0.654
(±0.006)
Mistral-Small-24B-
Base-2501
0.516 0.575
0.529
(±0.006)
0.633
(±0.007)
0.610
(±0.008)
0.731
(±0.004)
0.656
(±0.005)
0.786
(±0.005)
Mistral-Small-24B-
Instruct-2501
0.581 0.729
0.638
(±0.012)
0.758
(±0.007)
0.668
(±0.006)
0.757
(±0.007)
0.702
(±0.005)
0.804
(±0.005)
total average 0.481 0.520 0.513 0.557 0.580 0.591 0.615 0.612 0.740 0.564
Table 15: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the detailed instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
models zero-shot 1-shot (±SD) 5-shot (±SD) 10-shot (±SD)
GPT-4o mini 0.713 0.737 (±0.011) 0.773 (±0.005) 0.795 (±0.011)
GPT-4o 0.782 0.785 (±0.005) 0.800 (±0.012) 0.824 (±0.006)
GPT-4.1 mini 0.757 0.800 (±0.006) 0.840 (±0.007) 0.851 (±0.007)
GPT-4.1 0.881 0.863 (±0.003) 0.874 (±0.003) 0.891 (±0.001)
Haiku 4.5 0.772 0.775 (±0.005) 0.816 (±0.008) 0.837 (±0.006)
Sonnet 4.5 0.876 0.880 (±0.006) 0.875 (±0.002) 0.893 (±0.009)
total average 0.797 0.807 0.830 0.849
Table 16: Zero-shot and few-shot results of API models on Thunder-NUBench with the definition instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
models zero-shot 1-shot(±SD) 5-shot (±SD) 10-shot (±SD)
GPT-4o mini 0.756 0.768 (±0.007) 0.774 (±0.005) 0.790 (±0.011)
GPT-4o 0.863 0.847 (±0.006) 0.825 (±0.012) 0.826 (±0.006)
GPT-4.1 mini 0.902 0.873 (±0.004) 0.870 (±0.007) 0.870 (±0.007)
GPT-4.1 0.936 0.915 (±0.004) 0.901 (±0.003) 0.899 (±0.001)
Haiku 4.5 0.865 0.870 (±0.004) 0.870 (±0.008) 0.879 (±0.006)
Sonnet 4.5 0.915 0.914 (±0.004) 0.882 (±0.002) 0.825 (±0.009)
total average 0.873 0.865 0.854 0.848
Table 17: Zero-shot and few-shot results of API models on Thunder-NUBench with the detailed instruction. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column.
tasks ARC-Challenge ARC-Easy GSM8K HellaSwag OpenBookQA WinoGrande average
metric acc acc exact_match acc_norm acc_norm acc
gemma-2-2b 0.468 (+0.001) 0.797 (- 0.002) 0.259 (+0.002) 0.740 (- 0.001) 0.422 (- 0.002) 0.680 (+0.002) 0.561 (+0.000)
gemma-2-2b-it 0.498 (+0.003) 0.809 (+0.002) 0.438 (- 0.014) 0.716 (- 0.001) 0.434 (- 0.002) 0.680 (+0.007) 0.596 (- 0.001)
gemma-2-9b 0.616 (+0.005) 0.874 (+0.002) 0.672 (+0.035) 0.800 (+0.012) 0.472 (- 0.006) 0.741 (+0.009) 0.696 (+0.010)
gemma-2-9b-it 0.629 (+0.008) 0.858 (+0.010) 0.820 (+0.013) 0.800 (+0.000) 0.502 (- 0.008) 0.762 (- 0.010) 0.729 (+0.002)
Qwen3-0.6B-Base 0.326 (- 0.006) 0.668 (- 0.007) 0.510 (- 0.005) 0.517 (+0.002) 0.340 (+0.000) 0.594 (- 0.009) 0.493 (- 0.004)
Qwen3-0.6B 0.317 (- 0.010) 0.607 (- 0.011) 0.409 (+0.019) 0.473 (- 0.009) 0.318 (- 0.006) 0.560 (- 0.006) 0.447 (- 0.004)
Qwen3-1.7B-Base 0.416 (- 0.003) 0.733 (+0.013) 0.723 (- 0.034) 0.665 (+0.000) 0.392 (- 0.002) 0.639 (+0.006) 0.595 (- 0.003)
Qwen3-1.7B 0.403 (+0.009) 0.726 (+0.017) 0.692 (- 0.012) 0.604 (+0.004) 0.366 (- 0.004) 0.608 (+0.010) 0.567 (+0.004)
Qwen3-4B-Base 0.477 (+0.016) 0.789 (+0.006) 0.842 (- 0.008) 0.737 (+0.004) 0.408 (- 0.010) 0.706 (- 0.002) 0.660 (+0.001)
Qwen3-4B 0.507 (+0.017) 0.804 (+0.012) 0.840 (+0.026) 0.684 (+0.003) 0.402 (- 0.004) 0.661 (+0.004) 0.650 (+0.010)
Qwen3-8B-Base 0.532 (+0.002) 0.817 (+0.005) 0.850 (- 0.017) 0.785 (+0.003) 0.410 (+0.012) 0.723 (+0.002) 0.686 (+0.001)
Qwen3-8B 0.555 (+0.008) 0.835 (+0.005) 0.880 (+0.002) 0.750 (+0.006) 0.416 (+0.010) 0.677 (+0.017) 0.686 (+0.008)
Llama-3.1-8B 0.513 (+0.015) 0.814 (+0.014) 0.502 (+0.027) 0.790 (+0.007) 0.450 (+0.016) 0.739 (+0.004) 0.635 (+0.014)
Llama-3.1-8B-
Instruct
0.515 (+0.015) 0.819 (+0.006) 0.784 (- 0.012) 0.792 (- 0.004) 0.432 (+0.008) 0.735 (+0.004) 0.680 (+0.003)
Llama-3.2-1B 0.317 (- 0.012) 0.654 (- 0.010) 0.068 (- 0.011) 0.636 (- 0.004) 0.372 (+0.000) 0.603 (+0.010) 0.442 (- 0.005)
Llama-3.2-1B-
Instruct
0.358 (- 0.007) 0.683 (+0.002) 0.334 (- 0.002) 0.607 (- 0.006) 0.342 (+0.006) 0.597 (+0.004) 0.487 (- 0.001)
Llama-3.2-3B 0.422 (- 0.007) 0.743 (- 0.003) 0.262 (+0.025) 0.736 (+0.007) 0.428 (- 0.004) 0.698 (- 0.005) 0.548 (+0.002)
Llama-3.2-3B-
Instruct
0.436 (+0.003) 0.741 (+0.007) 0.649 (- 0.008) 0.705 (- 0.003) 0.362 (+0.002) 0.678 (- 0.001) 0.595 (+0.000)
Mistral-7B-v0.3 0.488 (+0.015) 0.796 (+0.011) 0.371 (- 0.053) 0.804 (+0.009) 0.436 (- 0.006) 0.736 (- 0.002) 0.605 (- 0.004)
Mistral-7B-Instruct-
v0.3
0.572 (- 0.008) 0.842 (- 0.007) 0.498 (- 0.043) 0.829 (- 0.004) 0.474 (- 0.016) 0.744 (- 0.014) 0.660 (- 0.015)
total average 0.601 (+0.001)
Table 18: General benchmark performance after supervised fine-tuning (SFT) on Thunder-NUBench with the definition instruction. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer. For each model and task, the main value reports the performance of the pre-SFT model, while the value in parentheses indicates the change after applying SFT.
tasks ARC-Challenge ARC-Easy GSM8K HellaSwag OpenBookQA WinoGrande average
metric acc acc exact_match acc_norm acc_norm acc
gemma-2-2b 0.468 (+0.000) 0.797 (+0.000) 0.259 (+0.000) 0.740 (+0.000) 0.422 (+0.000) 0.680 (+0.000) 0.561 (+0.000)
gemma-2-2b-it 0.498 (- 0.001) 0.809 (- 0.002) 0.438 (+0.000) 0.716 (- 0.001) 0.434 (+0.002) 0.680 (+0.002) 0.596 (+0.000)
gemma-2-9b 0.616 (+0.004) 0.874 (+0.001) 0.672 (+0.035) 0.800 (+0.013) 0.472 (- 0.008) 0.741 (+0.006) 0.696 (+0.009)
gemma-2-9b-it 0.629 (+0.006) 0.858 (+0.011) 0.820 (+0.005) 0.800 (+0.001) 0.502 (+0.002) 0.762 (- 0.004) 0.729 (+0.004)
Qwen3-0.6B-Base 0.326 (+0.000) 0.668 (+0.000) 0.510 (+0.000) 0.517 (+0.000) 0.340 (+0.000) 0.594 (+0.000) 0.493 (+0.000)
Qwen3-0.6B 0.317 (- 0.009) 0.607 (+0.016) 0.409 (+0.014) 0.473 (- 0.012) 0.318 (- 0.002) 0.560 (- 0.001) 0.447 (+0.001)
Qwen3-1.7B-Base 0.416 (+0.002) 0.733 (+0.003) 0.723 (- 0.035) 0.665 (+0.000) 0.392 (- 0.004) 0.639 (+0.002) 0.595 (- 0.005)
Qwen3-1.7B 0.403 (+0.003) 0.726 (+0.021) 0.692 (- 0.024) 0.604 (+0.003) 0.366 (+0.000) 0.608 (+0.009) 0.567 (+0.002)
Qwen3-4B-Base 0.477 (+0.026) 0.789 (+0.009) 0.842 (- 0.036) 0.737 (+0.005) 0.408 (- 0.004) 0.706 (+0.004) 0.660 (+0.001)
Qwen3-4B 0.507 (+0.008) 0.804 (+0.009) 0.840 (+0.008) 0.684 (+0.008) 0.402 (+0.006) 0.661 (+0.006) 0.650 (+0.008)
Qwen3-8B-Base 0.532 (+0.024) 0.817 (+0.011) 0.850 (- 0.040) 0.785 (+0.004) 0.410 (+0.014) 0.723 (- 0.002) 0.686 (+0.002)
Qwen3-8B 0.555 (+0.002) 0.835 (+0.000) 0.880 (- 0.002) 0.750 (+0.005) 0.416 (+0.006) 0.677 (+0.016) 0.686 (+0.005)
Llama-3.1-8B 0.513 (+0.003) 0.814 (+0.012) 0.502 (- 0.001) 0.790 (+0.010) 0.450 (+0.018) 0.739 (+0.001) 0.635 (+0.007)
Llama-3.1-8B-
Instruct
0.515 (+0.009) 0.819 (+0.003) 0.784 (- 0.013) 0.792 (- 0.006) 0.432 (+0.006) 0.735 (+0.009) 0.680 (+0.001)
Llama-3.2-1B 0.317 (- 0.009) 0.654 (- 0.005) 0.068 (- 0.024) 0.636 (- 0.008) 0.372 (- 0.004) 0.603 (+0.017) 0.442 (- 0.006)
Llama-3.2-1B-
Instruct
0.358 (- 0.004) 0.683 (+0.002) 0.334 (- 0.008) 0.607 (- 0.007) 0.342 (+0.008) 0.597 (+0.000) 0.487 (- 0.002)
Llama-3.2-3B 0.422 (- 0.005) 0.743 (- 0.003) 0.262 (+0.014) 0.736 (+0.004) 0.428 (- 0.006) 0.698 (+0.004) 0.548 (+0.001)
Llama-3.2-3B-
Instruct
0.436 (+0.001) 0.741 (+0.006) 0.649 (+0.001) 0.705 (- 0.001) 0.362 (+0.000) 0.678 (+0.000) 0.595 (+0.001)
Mistral-7B-v0.3 0.488 (+0.013) 0.796 (+0.011) 0.371 (- 0.094) 0.804 (+0.004) 0.436 (+0.008) 0.736 (+0.005) 0.605 (- 0.009)
Mistral-7B-Instruct-
v0.3
0.572 (- 0.009) 0.842 (- 0.004) 0.498 (- 0.057) 0.829 (- 0.003) 0.474 (- 0.006) 0.744 (- 0.014) 0.660 (- 0.016)
total average 0.601 (+0.000)
Table 19: General benchmark performance after supervised fine-tuning (SFT) on Thunder-NUBench with the detailed instruction. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer. For each model and task, the main value reports the performance of the pre-SFT model, while the value in parentheses indicates the change after applying SFT.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2-2b baseline zero-shot 0.564 74.96 19.41 5.63 25.64 32.47 63.27 53.87
1-shot 0.523 76.52 19.55 3.94 28.85 30.84 58.84 47.42
5-shot 0.420 81.47 17.01 1.51 21.47 25.00 51.02 44.19
10-shot 0.399 81.71 16.30 1.99 21.47 21.43 46.94 45.16
after SFT zero-shot 0.219 82.25 17.03 0.72 11.86 12.99 26.19 23.55
gemma-2-2b-it baseline zero-shot 0.485 75.29 22.09 2.62 22.12 23.38 48.64 56.77
1-shot 0.480 77.36 21.32 1.32 29.17 24.68 50.00 49.68
5-shot 0.427 78.07 20.07 1.86 23.08 23.05 46.94 44.84
10-shot 0.408 77.24 21.60 1.17 21.47 20.78 41.50 46.45
after SFT zero-shot 0.259 82.87 15.60 1.53 11.86 15.58 28.57 32.90
gemma-2-9b baseline zero-shot 0.530 75.30 19.61 5.09 24.36 31.49 61.56 48.06
1-shot 0.500 78.73 17.62 3.65 26.28 29.87 58.84 48.06
5-shot 0.431 83.79 14.00 2.21 22.44 27.27 50.00 49.68
10-shot 0.402 84.22 14.60 1.18 21.15 27.27 44.22 47.42
after SFT zero-shot 0.229 85.47 12.11 2.42 10.26 13.96 22.45 34.19
gemma-2-9b-it baseline zero-shot 0.493 76.01 22.06 1.93 24.68 27.92 52.72 49.68
1-shot 0.439 77.26 21.66 1.08 24.36 24.68 46.94 44.52
5-shot 0.352 78.38 20.72 0.90 17.31 17.86 31.97 46.77
10-shot 0.302 77.95 21.00 1.05 14.10 16.23 26.19 40.65
after SFT zero-shot 0.233 87.41 10.88 1.70 9.29 14.94 26.19 33.87
Table 20: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2-2b baseline zero-shot 0.765 33.20 26.66 40.15 22.12 29.55 29.25 23.87
1-shot 0.668 40.38 13.54 46.08 21.79 26.62 34.01 29.03
5-shot 0.623 53.69 12.47 33.84 27.88 34.42 40.48 35.48
10-shot 0.599 58.81 10.86 30.33 31.09 35.39 43.20 35.81
after SFT zero-shot 0.740 32.48 34.62 32.90 22.76 27.92 25.51 22.90
gemma-2-2b-it baseline zero-shot 0.496 74.88 8.80 16.32 37.18 26.30 66.67 24.19
1-shot 0.496 70.93 11.34 17.73 30.45 30.84 55.78 29.03
5-shot 0.473 67.67 12.56 19.77 31.09 33.12 43.88 24.52
10-shot 0.467 69.10 12.39 18.51 34.62 31.82 42.18 24.84
after SFT zero-shot 0.457 68.75 6.77 24.48 29.17 30.84 50.68 19.68
gemma-2-9b baseline zero-shot 0.532 62.89 7.60 29.51 36.22 29.87 57.82 15.16
1-shot 0.472 62.35 4.71 32.94 34.94 24.35 48.30 14.52
5-shot 0.402 71.79 6.11 22.09 33.01 23.70 45.92 17.10
10-shot 0.378 71.43 4.62 23.95 33.33 20.78 40.82 16.77
after SFT zero-shot 0.249 64.01 16.24 19.75 21.15 11.36 23.81 9.68
gemma-2-9b-it baseline zero-shot 0.447 73.00 22.02 4.97 39.42 22.40 51.70 21.61
1-shot 0.370 80.69 13.95 5.36 35.26 22.73 42.86 22.58
5-shot 0.286 83.66 11.63 4.71 28.85 17.53 31.63 20.97
10-shot 0.272 82.51 12.54 4.96 28.85 17.53 28.23 18.06
after SFT zero-shot 0.186 73.62 21.70 4.68 11.86 10.71 25.51 9.03
Table 21: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2-2b baseline zero-shot 0.562 74.89 19.04 6.06 26.60 34.74 60.88 52.26
1-shot 0.507 76.37 19.41 4.23 26.60 30.84 56.46 46.45
5-shot 0.425 82.09 16.04 1.87 20.83 25.97 52.72 45.16
10-shot 0.392 81.98 15.99 2.02 20.19 22.40 46.26 44.19
after SFT zero-shot 0.232 82.59 16.72 0.68 11.22 13.64 27.21 27.42
gemma-2-2b-it baseline zero-shot 0.495 78.37 18.75 2.88 24.68 27.60 54.76 53.55
1-shot 0.493 79.74 18.81 1.45 29.17 28.25 52.38 52.90
5-shot 0.450 80.81 17.61 1.58 25.64 24.35 51.70 49.03
10-shot 0.422 80.26 18.42 1.32 22.76 22.40 46.60 48.39
after SFT zero-shot 0.267 81.90 16.62 1.48 12.50 14.61 32.65 30.97
gemma-2-9b baseline zero-shot 0.531 74.63 20.60 4.78 25.00 33.77 62.59 43.23
1-shot 0.487 78.50 17.59 3.91 25.96 30.52 59.18 42.90
5-shot 0.425 83.21 14.37 2.43 21.47 25.65 51.36 48.06
10-shot 0.393 83.23 15.35 1.41 20.83 25.97 43.54 44.84
after SFT zero-shot 0.228 86.06 11.85 2.09 8.97 14.61 23.81 33.55
gemma-2-9b-it baseline zero-shot 0.458 76.95 21.14 1.91 18.59 27.60 56.80 43.23
1-shot 0.401 77.62 21.58 0.79 20.51 20.78 45.24 42.26
5-shot 0.345 78.85 20.23 0.92 16.35 17.53 33.67 44.84
10-shot 0.308 78.87 19.85 1.29 14.10 17.21 26.53 42.26
after SFT zero-shot 0.236 85.91 11.41 2.68 10.26 13.64 26.87 33.23
Table 22: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
gemma-2-2b baseline zero-shot 0.724 35.93 29.46 34.61 23.72 31.49 29.25 22.90
1-shot 0.660 42.91 15.02 42.07 22.76 28.90 35.37 30.00
5-shot 0.629 53.47 12.48 34.05 28.21 34.74 40.82 35.16
10-shot 0.612 57.64 11.53 30.83 32.05 34.42 41.84 37.42
after SFT zero-shot 0.741 32.55 34.58 32.87 22.76 27.92 25.51 23.23
gemma-2-2b-it baseline zero-shot 0.531 82.06 7.47 10.46 42.95 38.96 67.69 30.97
1-shot 0.539 70.84 8.84 20.32 32.37 31.82 63.27 30.97
5-shot 0.508 64.06 12.81 23.12 30.13 32.47 45.92 26.13
10-shot 0.504 61.64 12.11 26.26 31.41 29.22 43.20 24.84
after SFT zero-shot 0.337 72.00 10.12 17.88 21.15 25.65 35.71 18.06
gemma-2-9b baseline zero-shot 0.534 73.70 8.47 17.83 42.63 32.47 66.33 21.94
1-shot 0.476 64.83 6.50 28.67 36.22 25.32 50.00 16.45
5-shot 0.422 70.11 5.45 24.44 33.65 22.73 48.30 18.06
10-shot 0.403 69.88 3.94 26.18 34.62 20.45 44.22 17.42
after SFT zero-shot 0.203 75.00 11.72 13.28 22.12 12.01 16.33 12.26
gemma-2-9b-it baseline zero-shot 0.431 83.27 11.58 5.15 35.90 23.70 63.61 26.13
1-shot 0.361 83.30 10.77 5.93 35.58 21.75 43.54 23.55
5-shot 0.305 84.11 10.42 5.47 30.45 21.43 32.31 21.61
10-shot 0.284 82.12 12.85 5.03 30.13 18.51 28.57 19.03
after SFT zero-shot 0.195 69.92 17.07 13.01 16.67 13.96 12.59 12.90
Table 23: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen3-0.6B-Base baseline zero-shot 0.547 72.32 23.48 4.20 29.49 34.74 50.68 48.71
1-shot 0.487 74.27 22.15 3.58 31.09 28.90 46.60 42.90
5-shot 0.412 77.07 19.46 3.47 23.40 21.75 43.54 42.58
10-shot 0.390 79.27 18.90 1.83 22.76 22.08 40.82 42.26
after SFT zero-shot 0.292 83.97 15.22 0.82 18.59 17.53 30.27 34.84
Qwen3-0.6B baseline zero-shot 0.584 74.46 20.52 5.03 26.92 32.14 64.63 56.45
1-shot 0.571 76.81 19.17 4.03 29.81 34.42 62.24 55.16
5-shot 0.515 79.82 16.95 3.24 25.96 28.57 57.48 58.06
10-shot 0.470 80.74 15.71 3.55 25.00 25.97 48.64 57.10
after SFT zero-shot 0.355 84.82 13.84 1.34 20.51 20.78 38.10 45.16
Qwen3-1.7B-Base baseline zero-shot 0.519 73.70 22.02 4.28 28.21 34.74 56.80 38.71
1-shot 0.467 74.87 22.92 2.21 27.88 30.19 50.34 36.45
5-shot 0.416 76.57 21.14 2.29 22.12 26.95 42.86 40.00
10-shot 0.396 79.56 18.84 1.60 23.72 24.35 41.84 40.32
after SFT zero-shot 0.301 89.47 9.74 0.79 19.23 24.35 36.05 31.94
Qwen3-1.7B baseline zero-shot 0.602 67.46 24.64 7.91 28.53 37.99 65.31 36.77
1-shot 0.601 71.50 24.01 4.49 31.41 34.74 64.29 47.74
5-shot 0.517 76.61 20.64 2.75 26.92 28.25 58.84 50.65
10-shot 0.483 78.33 19.70 1.97 27.88 24.35 54.08 50.32
after SFT zero-shot 0.332 83.77 14.56 1.67 15.71 24.68 37.41 37.42
Qwen3-4B-Base baseline zero-shot 0.546 76.34 20.46 3.19 31.73 32.47 62.59 46.13
1-shot 0.515 74.73 22.03 3.24 30.13 30.52 54.42 44.19
5-shot 0.467 74.19 22.41 3.40 26.28 27.27 48.64 41.29
10-shot 0.443 76.21 21.65 2.15 25.64 25.65 47.96 40.65
after SFT zero-shot 0.267 89.61 10.09 0.30 16.03 20.13 30.95 31.94
Qwen3-4B baseline zero-shot 0.590 73.66 20.83 5.51 31.73 40.58 68.37 39.68
1-shot 0.534 76.97 18.13 4.90 35.26 33.77 62.24 39.03
5-shot 0.435 80.66 17.34 2.01 25.00 25.97 52.72 41.61
10-shot 0.408 81.91 16.34 1.75 23.72 26.95 48.30 39.35
after SFT zero-shot 0.285 85.79 12.81 1.39 16.03 18.83 24.83 40.97
Qwen3-8B-Base baseline zero-shot 0.526 73.76 22.47 3.77 27.56 33.12 61.56 38.71
1-shot 0.502 72.67 23.85 3.48 27.24 28.57 56.12 39.35
5-shot 0.420 73.91 22.68 3.40 20.83 23.05 44.22 40.32
10-shot 0.401 77.62 20.40 1.98 22.44 25.32 40.14 40.65
after SFT zero-shot 0.297 85.07 13.60 1.33 16.67 19.81 35.03 33.23
Qwen3-8B baseline zero-shot 0.558 71.59 22.87 5.54 27.24 34.74 65.31 38.71
1-shot 0.501 77.22 18.20 4.59 32.37 32.14 61.22 34.84
5-shot 0.436 80.73 16.73 2.55 26.92 30.84 48.98 39.03
10-shot 0.397 81.20 17.00 1.80 25.32 26.95 44.56 36.45
after SFT zero-shot 0.278 83.48 15.10 1.42 14.74 16.56 31.97 32.90
Qwen3-14B-Base baseline zero-shot 0.512 70.08 23.57 6.36 24.68 30.84 53.74 39.35
1-shot 0.450 72.01 23.06 4.93 24.68 26.62 45.58 37.42
5-shot 0.380 75.57 21.50 2.92 20.83 22.73 37.76 37.42
10-shot 0.347 76.89 21.05 2.06 20.51 20.13 31.97 37.42
Qwen3-14B baseline zero-shot 0.560 72.52 21.10 6.37 27.88 32.47 65.31 42.90
1-shot 0.500 78.92 16.64 4.44 32.37 31.17 59.18 40.97
5-shot 0.429 80.41 17.19 2.40 25.64 26.95 47.28 42.90
10-shot 0.381 82.29 15.62 2.08 23.72 25.00 36.05 44.52
Qwen3-32B baseline zero-shot 0.531 76.68 17.64 5.68 27.24 34.42 67.01 40.32
1-shot 0.462 77.49 17.70 4.81 26.28 28.25 55.78 38.06
5-shot 0.347 81.46 15.79 2.75 18.91 22.73 39.80 35.48
10-shot 0.316 83.17 14.32 2.51 19.87 20.78 32.31 35.48
Table 24: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen3-0.6B-Base baseline zero-shot 0.651 50.55 7.06 42.39 25.32 37.99 39.80 32.90
1-shot 0.603 55.26 6.84 37.89 25.64 37.01 48.98 26.45
5-shot 0.577 55.57 3.30 41.13 27.56 33.12 47.28 24.84
10-shot 0.525 60.12 3.78 36.10 29.81 30.52 45.92 24.52
after SFT zero-shot 0.521 67.43 11.42 21.16 32.37 39.94 42.18 30.65
Qwen3-0.6B baseline zero-shot 0.562 69.96 7.90 22.14 37.82 38.64 54.08 32.26
1-shot 0.532 59.31 14.75 25.93 26.28 29.87 47.28 27.42
5-shot 0.493 71.66 5.80 22.54 32.37 30.52 54.76 28.71
10-shot 0.469 69.88 6.60 23.52 32.37 29.87 49.32 24.19
after SFT zero-shot 0.445 67.02 10.34 22.64 24.68 27.27 47.96 23.87
Qwen3-1.7B-Base baseline zero-shot 0.474 65.22 9.87 24.92 31.41 23.70 52.04 21.29
1-shot 0.444 72.50 13.57 13.93 33.97 30.84 38.78 29.35
5-shot 0.361 75.82 12.53 11.65 29.81 25.32 36.39 21.61
10-shot 0.313 76.14 11.42 12.44 25.96 20.78 34.69 17.10
after SFT zero-shot 0.507 58.84 12.83 28.33 33.97 26.30 43.54 19.68
Qwen3-1.7B baseline zero-shot 0.468 70.85 15.76 13.39 26.92 31.82 46.26 32.26
1-shot 0.451 67.31 14.24 18.45 24.04 28.57 47.62 25.81
5-shot 0.424 72.47 11.80 15.73 26.60 23.70 51.70 25.48
10-shot 0.421 72.13 9.98 17.89 25.96 23.70 53.40 23.23
after SFT zero-shot 0.392 68.83 22.27 8.91 30.13 27.92 30.27 22.90
Qwen3-4B-Base baseline zero-shot 0.398 78.69 15.74 5.58 38.14 22.40 45.58 23.55
1-shot 0.386 77.82 19.30 2.87 40.71 29.22 36.73 17.42
5-shot 0.350 75.74 19.27 4.99 40.38 26.95 24.15 17.42
10-shot 0.336 77.36 17.45 5.19 39.10 26.95 24.49 16.45
after SFT zero-shot 0.291 89.37 8.45 2.18 32.05 18.18 43.20 14.52
Qwen3-4B baseline zero-shot 0.439 82.28 15.55 2.17 38.46 30.19 56.46 24.52
1-shot 0.375 80.97 17.34 1.69 33.97 25.32 46.26 20.32
5-shot 0.321 84.20 12.35 3.46 32.69 23.70 33.67 21.61
10-shot 0.301 83.64 13.19 3.17 30.77 24.68 26.87 21.29
after SFT zero-shot 0.278 87.14 9.43 3.43 28.53 22.73 37.76 11.29
Qwen3-8B-Base baseline zero-shot 0.310 79.03 13.81 7.16 29.17 20.78 33.33 18.06
1-shot 0.313 76.96 13.42 9.62 31.73 26.95 23.81 16.77
5-shot 0.254 76.25 15.00 8.75 28.21 21.10 16.67 13.55
10-shot 0.250 78.41 13.02 8.57 27.56 20.45 16.67 15.81
after SFT zero-shot 0.232 72.7 20.82 6.48 19.55 17.21 18.71 14.19
Qwen3-8B baseline zero-shot 0.411 84.17 12.55 3.28 41.67 31.82 49.32 20.32
1-shot 0.326 82.00 13.14 4.87 35.90 26.30 33.33 14.84
5-shot 0.279 83.81 10.80 5.40 34.29 25.32 21.09 15.48
10-shot 0.256 82.35 11.15 6.50 30.77 23.38 16.67 15.81
after SFT zero-shot 0.217 84.62 11.72 3.66 25.00 22.40 15.65 12.26
Qwen3-14B-Base baseline zero-shot 0.292 81.52 15.22 3.26 23.72 16.56 39.12 19.35
1-shot 0.236 75.17 18.79 6.04 21.15 16.56 26.53 9.35
5-shot 0.190 73.22 23.01 3.77 18.59 14.61 13.27 10.65
10-shot 0.177 75.34 21.97 2.69 17.95 14.61 9.52 12.58
Qwen3-14B baseline zero-shot 0.340 79.67 17.52 2.80 32.37 22.08 45.92 11.94
1-shot 0.316 82.96 12.28 4.76 30.77 21.75 43.20 13.23
5-shot 0.263 85.20 12.99 1.81 31.09 25.32 26.19 9.68
10-shot 0.250 84.76 12.70 2.54 33.01 21.75 20.75 11.61
Qwen3-32B baseline zero-shot 0.270 81.23 16.72 2.05 25.00 24.35 26.53 14.84
1-shot 0.232 80.82 18.15 1.03 24.68 23.05 19.73 9.68
5-shot 0.201 81.10 18.50 0.39 24.36 19.48 14.97 8.39
10-shot 0.181 77.63 21.93 0.44 19.55 16.88 12.24 9.03
Table 25: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen3-0.6B-Base baseline zero-shot 0.577 71.53 22.97 5.50 29.81 35.39 54.42 50.97
1-shot 0.502 73.46 21.80 4.74 29.17 30.84 50.34 42.26
5-shot 0.426 77.28 19.18 3.54 23.40 23.70 44.22 44.84
10-shot 0.401 79.01 18.61 2.38 22.12 23.70 42.18 42.90
after SFT zero-shot 0.277 83.67 15.47 0.86 17.31 17.53 28.23 32.58
Qwen3-0.6B baseline zero-shot 0.587 77.70 17.57 4.73 27.24 41.56 62.93 57.10
1-shot 0.590 76.48 18.82 4.70 30.77 38.64 64.29 53.23
5-shot 0.517 78.53 17.18 4.29 25.00 31.82 57.82 53.55
10-shot 0.471 81.31 15.15 3.54 25.64 27.92 52.38 52.58
after SFT zero-shot 0.321 84.44 14.07 1.48 17.63 20.13 35.03 39.35
Qwen3-1.7B-Base baseline zero-shot 0.538 71.24 22.71 6.05 29.81 33.12 58.50 37.42
1-shot 0.479 74.17 22.85 2.98 28.21 30.52 53.06 35.48
5-shot 0.417 77.38 20.53 2.09 23.40 26.30 41.84 41.94
10-shot 0.401 80.40 18.22 1.39 23.72 24.68 41.84 42.90
after SFT zero-shot 0.276 85.63 12.93 1.44 17.95 24.03 31.97 23.87
Qwen3-1.7B baseline zero-shot 0.581 67.76 25.27 6.97 27.88 37.66 63.27 34.52
1-shot 0.570 71.35 24.34 4.31 29.49 33.77 61.22 44.19
5-shot 0.504 76.54 21.10 2.36 26.60 26.95 58.16 48.06
10-shot 0.485 77.45 20.42 2.12 27.56 25.32 54.42 48.39
after SFT zero-shot 0.358 82.71 15.74 1.55 18.59 26.30 39.12 38.39
Qwen3-4B-Base baseline zero-shot 0.515 75.50 21.57 2.93 27.56 30.84 55.10 47.42
1-shot 0.494 75.12 22.15 2.73 29.49 29.55 52.72 41.94
5-shot 0.449 74.20 22.79 3.00 25.00 27.27 44.22 41.29
10-shot 0.440 76.22 21.80 1.98 25.64 25.97 45.58 41.61
after SFT zero-shot 0.271 86.26 12.87 0.88 16.67 19.48 31.63 29.03
Qwen3-4B baseline zero-shot 0.562 73.73 20.34 5.93 28.21 38.96 65.31 39.35
1-shot 0.495 74.68 20.19 5.13 28.85 29.87 57.82 36.77
5-shot 0.431 78.12 19.12 2.76 23.08 25.65 51.70 39.35
10-shot 0.398 79.08 18.53 2.39 21.79 25.65 46.60 36.45
after SFT zero-shot 0.313 83.76 14.47 1.78 16.67 21.10 28.91 41.29
Qwen3-8B-Base baseline zero-shot 0.545 74.09 21.69 4.22 29.17 33.44 61.22 43.55
1-shot 0.489 73.42 23.18 3.40 27.56 28.25 55.44 37.74
5-shot 0.436 73.09 23.27 3.64 22.44 24.35 44.90 40.32
10-shot 0.412 76.35 21.15 2.50 22.76 25.65 40.48 41.29
after SFT zero-shot 0.286 86.43 12.47 1.11 16.35 20.45 34.01 31.61
Qwen3-8B baseline zero-shot 0.527 71.84 23.04 5.12 26.60 35.06 61.22 34.19
1-shot 0.484 75.41 20.49 4.10 31.73 30.19 57.14 32.26
5-shot 0.442 80.43 17.59 1.97 27.56 30.52 49.32 39.68
10-shot 0.408 81.36 16.70 1.94 25.64 28.57 45.24 38.06
after SFT zero-shot 0.260 83.84 14.33 1.83 13.78 17.21 27.55 31.61
Qwen3-14B-Base baseline zero-shot 0.493 70.69 23.67 5.64 25.00 31.17 46.60 41.29
1-shot 0.443 74.37 22.58 3.05 25.64 27.27 46.26 37.10
5-shot 0.369 76.99 20.65 2.37 20.19 22.08 35.37 39.68
10-shot 0.335 77.07 21.75 1.18 20.51 19.81 28.91 37.42
Qwen3-14B baseline zero-shot 0.512 69.35 23.37 7.28 25.64 33.12 56.46 32.26
1-shot 0.470 75.89 18.89 5.23 31.09 30.84 55.78 30.32
5-shot 0.397 79.60 18.60 1.80 23.72 25.65 44.90 36.45
10-shot 0.361 79.56 18.24 2.20 23.72 24.03 33.67 37.10
Qwen3-32B baseline zero-shot 0.486 70.31 20.72 8.97 24.68 35.06 60.88 21.61
1-shot 0.418 74.95 18.60 6.45 23.72 27.60 53.06 25.81
5-shot 0.341 81.40 15.58 3.02 20.51 23.38 40.14 30.97
10-shot 0.309 82.01 15.17 2.83 19.23 21.75 32.99 30.65
Table 26: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Qwen3-0.6B-Base baseline zero-shot 0.658 48.80 4.46 46.75 27.88 37.66 38.78 28.39
1-shot 0.606 53.14 7.46 39.40 25.00 36.04 47.62 24.84
5-shot 0.588 52.77 3.24 43.99 26.60 33.12 43.20 25.48
10-shot 0.537 61.00 3.55 35.45 30.45 33.77 46.60 24.84
after SFT zero-shot 0.675 37.25 30.79 31.96 23.08 27.92 31.63 21.29
Qwen3-0.6B baseline zero-shot 0.541 74.93 7.92 17.16 43.91 40.91 57.48 25.48
1-shot 0.588 53.37 15.36 31.27 25.96 29.87 45.58 28.71
5-shot 0.535 65.04 6.81 28.15 31.09 31.49 52.72 29.03
10-shot 0.502 65.72 6.64 27.65 31.73 29.87 52.72 22.58
after SFT zero-shot 0.474 64.55 20.57 14.88 27.56 29.22 46.60 23.55
Qwen3-1.7B-Base baseline zero-shot 0.517 57.21 13.50 29.29 26.92 28.25 49.32 18.39
1-shot 0.457 69.97 13.89 16.15 31.73 32.79 40.14 27.42
5-shot 0.382 71.78 12.86 15.35 29.49 25.97 38.10 20.00
10-shot 0.343 72.29 11.55 16.17 25.32 23.70 36.73 17.10
after SFT zero-shot 0.391 68.15 13.18 18.66 30.45 26.30 41.16 12.58
Qwen3-1.7B baseline zero-shot 0.553 71.74 18.94 9.33 37.50 37.99 53.74 34.84
1-shot 0.508 68.12 14.06 17.81 30.45 31.82 55.44 25.81
5-shot 0.450 72.31 10.41 17.28 27.56 25.97 56.80 24.84
10-shot 0.455 71.25 10.10 18.64 30.13 25.65 55.78 23.23
after SFT zero-shot 0.427 66.05 26.90 7.05 29.81 29.87 34.35 22.58
Qwen3-4B-Base baseline zero-shot 0.416 79.01 14.89 6.11 36.86 24.68 48.64 25.81
1-shot 0.413 76.20 18.23 5.57 41.35 27.92 42.86 18.06
5-shot 0.366 75.49 17.57 6.94 39.42 25.97 30.27 18.06
10-shot 0.355 76.06 17.00 6.94 40.71 26.30 27.21 16.77
after SFT zero-shot 0.290 84.38 11.78 3.84 28.85 15.91 39.80 16.77
Qwen3-4B baseline zero-shot 0.424 81.27 17.23 1.50 37.18 29.22 51.36 24.84
1-shot 0.364 79.52 18.52 1.96 35.26 24.03 41.16 19.35
5-shot 0.347 82.88 14.61 2.51 32.37 26.30 38.10 22.26
10-shot 0.317 82.00 14.50 3.50 33.97 23.70 29.93 19.68
after SFT zero-shot 0.289 81.32 17.58 1.10 28.21 21.43 35.37 12.26
Qwen3-8B-Base baseline zero-shot 0.332 78.04 12.41 9.55 29.81 24.68 36.05 16.77
1-shot 0.305 77.86 13.80 8.33 30.45 26.62 25.51 15.16
5-shot 0.258 75.08 16.31 8.62 27.24 20.13 18.37 13.87
10-shot 0.251 78.23 12.93 8.83 27.24 19.81 18.03 15.81
after SFT zero-shot 0.177 74.89 16.14 8.97 17.63 14.61 9.52 12.58
Qwen3-8B baseline zero-shot 0.381 86.07 12.27 1.66 43.59 33.12 37.41 21.29
1-shot 0.305 83.38 14.03 2.60 36.54 24.03 26.87 17.42
5-shot 0.297 81.87 13.60 4.53 34.62 26.30 21.77 17.42
10-shot 0.276 81.32 12.93 5.75 31.73 23.70 18.71 18.06
after SFT zero-shot 0.240 85.48 10.56 3.96 29.17 21.43 19.05 14.84
Qwen3-14B-Base baseline zero-shot 0.264 79.58 15.92 4.50 20.83 16.56 27.89 21.61
1-shot 0.242 73.11 21.97 4.92 21.79 16.88 20.75 13.55
5-shot 0.210 75.09 21.51 3.40 22.12 15.58 13.95 13.23
10-shot 0.182 77.29 20.52 2.18 18.59 15.58 8.50 14.84
Qwen3-14B baseline zero-shot 0.280 79.04 19.55 1.42 28.85 23.05 25.17 14.19
1-shot 0.291 80.65 15.53 3.81 31.73 21.10 29.93 14.19
5-shot 0.241 86.51 11.18 2.30 30.13 20.78 20.75 14.19
10-shot 0.241 84.21 13.82 1.97 31.73 21.75 16.33 13.55
Qwen3-32B baseline zero-shot 0.240 78.81 19.54 1.66 20.83 24.35 18.03 14.52
1-shot 0.189 80.67 18.49 0.84 22.44 20.45 10.88 8.71
5-shot 0.194 80.82 18.78 0.41 21.47 20.45 12.93 9.68
10-shot 0.175 80.54 19.00 0.45 19.23 18.18 9.18 11.29
Table 27: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.1-8B baseline zero-shot 0.536 76.92 17.60 5.47 25.32 30.19 60.20 55.16
1-shot 0.501 80.06 16.14 3.80 24.68 29.22 57.14 55.16
5-shot 0.391 83.77 14.20 2.03 21.79 23.05 43.54 47.10
10-shot 0.356 83.07 15.14 1.78 19.23 21.43 38.78 42.90
after SFT zero-shot 0.226 88.07 10.18 1.75 10.58 15.91 24.15 31.61
Llama-3.1-8B- Instruct baseline zero-shot 0.536 73.82 23.96 2.22 25.00 31.82 58.50 48.71
1-shot 0.457 79.51 19.10 1.39 24.36 28.57 48.98 48.39
5-shot 0.367 85.31 14.25 0.43 18.91 22.40 40.14 48.06
10-shot 0.332 85.17 14.35 0.48 20.51 20.78 32.99 42.26
after SFT zero-shot 0.229 86.51 12.46 1.04 10.90 14.29 23.13 33.55
Llama-3.2-1B baseline zero-shot 0.591 71.01 24.03 4.97 21.79 29.55 63.95 58.71
1-shot 0.570 72.32 23.37 4.31 25.32 32.47 63.95 49.35
5-shot 0.514 75.77 20.52 3.70 20.83 27.27 63.95 49.68
10-shot 0.468 78.14 18.64 3.22 19.87 25.97 59.86 46.13
after SFT zero-shot 0.250 76.83 21.59 1.59 13.46 14.61 24.15 27.10
Llama-3.2-1B- Instruct baseline zero-shot 0.569 68.48 27.89 3.63 25.00 30.84 58.16 47.42
1-shot 0.526 69.83 27.75 2.41 25.32 26.95 54.76 45.16
5-shot 0.472 70.92 25.38 3.70 21.15 24.68 51.36 41.61
10-shot 0.447 72.70 23.94 3.37 22.12 22.40 47.96 42.26
after SFT zero-shot 0.239 81.06 17.94 1.00 13.14 15.26 21.43 30.00
Llama-3.2-3B baseline zero-shot 0.565 74.58 19.94 5.48 24.36 31.49 62.93 55.81
1-shot 0.538 78.17 18.88 2.95 31.09 31.82 60.88 50.32
5-shot 0.429 79.85 17.56 2.59 22.44 27.27 47.96 44.19
10-shot 0.413 81.96 15.16 2.88 23.72 26.95 45.92 43.55
after SFT zero-shot 0.244 84.36 14.33 1.30 10.58 16.88 24.83 32.58
Llama-3.2-3B- Instruct baseline zero-shot 0.548 76.70 21.71 1.59 27.88 33.12 54.42 58.39
1-shot 0.536 76.33 22.34 1.33 31.41 31.17 52.72 53.87
5-shot 0.448 78.76 20.18 1.06 24.68 26.95 40.82 53.23
10-shot 0.414 79.12 19.73 1.15 24.04 24.35 37.41 49.35
after SFT zero-shot 0.205 84.50 13.18 2.33 8.33 10.06 25.17 28.06
Table 28: Error rates, incorrect choice distributions, and local negation confusion rates for the Llama3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.1-8B baseline zero-shot 0.526 62.29 8.14 29.56 28.21 24.68 54.42 28.71
1-shot 0.493 70.42 6.11 23.47 36.22 30.52 47.28 29.68
5-shot 0.422 84.59 4.32 11.09 41.35 26.95 49.32 30.00
10-shot 0.372 88.27 4.48 7.25 36.86 25.32 40.48 32.90
after SFT zero-shot 0.441 62.77 17.63 19.60 27.24 20.45 41.16 25.81
Llama-3.1-8B- Instruct baseline zero-shot 0.462 70.67 18.52 10.81 27.56 24.68 53.40 30.00
1-shot 0.404 70.53 18.07 11.39 26.92 23.70 41.84 25.48
5-shot 0.332 81.62 11.22 7.16 25.64 18.83 40.48 27.42
10-shot 0.325 85.85 9.27 4.88 28.85 19.81 37.41 29.35
after SFT zero-shot 0.356 71.94 15.37 12.69 25.00 19.81 44.56 17.10
Llama-3.2-1B baseline zero-shot 0.741 32.44 34.69 32.87 22.76 27.92 25.51 22.90
1-shot 0.719 30.35 33.44 36.20 20.83 20.78 25.85 22.58
5-shot 0.751 33.16 34.85 32.00 24.36 24.68 27.89 25.81
10-shot 0.738 33.08 33.83 33.08 25.00 25.65 25.85 24.19
after SFT zero-shot 0.741 32.44 34.69 32.87 22.76 27.92 25.51 22.90
Llama-3.2-1B- Instruct baseline zero-shot 0.699 39.68 27.89 32.43 26.28 34.09 29.59 24.52
1-shot 0.608 53.59 19.17 27.25 25.32 39.29 41.50 28.71
5-shot 0.645 47.36 22.88 29.77 27.56 38.64 31.97 27.74
10-shot 0.678 43.98 25.26 30.76 28.21 33.77 33.33 27.74
after SFT zero-shot 0.714 35.78 29.22 35.00 23.72 30.52 26.53 24.52
Llama-3.2-3B baseline zero-shot 0.638 50.37 26.37 23.26 23.08 35.39 40.82 33.55
1-shot 0.555 68.43 14.00 17.57 33.33 34.74 46.94 41.94
5-shot 0.527 79.70 10.83 9.47 38.14 37.66 51.02 46.77
10-shot 0.516 83.23 10.15 6.62 40.38 36.36 49.66 50.65
after SFT zero-shot 0.615 46.97 28.77 24.26 22.44 30.52 35.71 30.65
Llama-3.2-3B- Instruct baseline zero-shot 0.488 80.00 7.97 12.03 27.24 35.39 64.29 35.16
1-shot 0.458 77.68 9.86 12.46 24.36 28.25 60.88 34.52
5-shot 0.502 74.25 8.85 16.90 29.49 27.27 67.35 30.97
10-shot 0.498 74.04 8.44 17.52 31.09 26.62 64.29 31.29
after SFT zero-shot 0.349 70.45 6.59 22.95 15.71 16.88 45.58 24.19
Table 29: Error rates, incorrect choice distributions, and local negation confusion rates for the Llama3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.1-8B baseline zero-shot 0.549 73.99 19.08 6.94 26.60 28.57 61.56 51.61
1-shot 0.516 75.42 20.28 4.30 25.96 25.32 58.50 51.61
5-shot 0.424 79.07 17.76 3.18 22.12 23.05 47.62 46.13
10-shot 0.381 81.08 16.63 2.29 20.51 22.08 41.50 43.87
after SFT zero-shot 0.198 88.00 11.60 0.40 9.62 13.31 21.09 28.06
Llama-3.1-8B- Instruct baseline zero-shot 0.468 71.86 26.44 1.69 23.08 29.87 47.62 38.71
1-shot 0.408 78.99 19.65 1.36 25.00 26.95 43.20 38.06
5-shot 0.347 84.70 14.84 0.46 19.55 23.05 37.76 41.29
10-shot 0.316 84.46 15.04 0.50 21.15 19.48 31.29 38.39
after SFT zero-shot 0.211 87.97 10.90 1.13 8.97 13.31 21.09 33.23
Llama-3.2-1B baseline zero-shot 0.597 69.85 24.70 5.44 24.04 32.47 67.01 49.68
1-shot 0.565 72.75 22.89 4.35 26.28 30.52 64.63 49.03
5-shot 0.510 76.21 20.22 3.58 20.83 26.95 64.97 48.71
10-shot 0.470 78.08 18.55 3.37 20.51 25.65 60.20 46.13
after SFT zero-shot 0.253 80.25 18.50 1.25 14.42 15.91 24.83 28.71
Llama-3.2-1B- Instruct baseline zero-shot 0.584 68.89 27.72 3.40 27.56 34.42 57.48 47.10
1-shot 0.528 69.67 27.18 3.15 24.36 27.60 53.06 47.42
5-shot 0.478 71.31 24.71 3.98 20.83 24.03 53.40 43.23
10-shot 0.453 72.33 23.99 3.68 21.79 21.43 50.00 42.58
after SFT zero-shot 0.297 79.41 18.98 1.60 14.74 18.83 33.33 30.65
Llama-3.2-3B baseline zero-shot 0.553 73.17 19.66 7.17 23.72 35.06 60.20 48.71
1-shot 0.522 77.05 19.30 3.65 27.88 30.52 59.86 48.39
5-shot 0.427 80.67 16.54 2.79 22.12 28.25 45.92 46.13
10-shot 0.413 82.73 14.40 2.88 25.00 27.60 44.56 44.19
after SFT zero-shot 0.240 81.13 16.89 1.99 9.29 14.94 24.49 31.61
Llama-3.2-3B- Instruct baseline zero-shot 0.512 76.63 21.98 1.39 25.64 34.09 47.62 54.84
1-shot 0.527 77.26 21.54 1.20 30.45 30.84 49.66 57.10
5-shot 0.456 80.35 18.61 1.04 24.36 28.57 42.86 55.48
10-shot 0.424 81.68 17.20 1.12 24.68 27.27 38.78 52.26
after SFT zero-shot 0.209 82.89 15.21 1.90 9.62 12.34 26.87 22.90
Table 30: Error rates, incorrect choice distributions, and local negation confusion rates for the Llama3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Llama-3.1-8B baseline zero-shot 0.569 66.43 8.64 24.93 32.69 31.49 59.18 33.55
1-shot 0.519 73.85 5.20 20.95 40.71 35.71 48.30 33.55
5-shot 0.448 85.84 3.89 10.27 43.27 31.82 49.32 34.52
10-shot 0.385 88.25 4.12 7.63 38.14 27.92 40.48 33.55
after SFT zero-shot 0.328 69.81 18.84 11.35 22.76 20.78 35.71 15.81
Llama-3.1-8B- Instruct baseline zero-shot 0.486 71.45 14.85 13.70 33.01 25.97 56.12 29.03
1-shot 0.402 67.26 17.16 15.58 28.85 21.10 36.39 25.48
5-shot 0.353 77.53 11.24 11.24 27.56 19.81 39.12 26.77
10-shot 0.349 83.41 9.09 7.50 31.41 19.48 39.12 30.32
after SFT zero-shot 0.270 65.69 21.70 12.61 14.74 14.61 29.25 15.16
Llama-3.2-1B baseline zero-shot 0.741 32.44 34.69 32.87 22.76 27.92 25.51 22.90
1-shot 0.726 29.29 33.66 37.05 21.47 18.83 24.49 22.90
5-shot 0.744 33.05 34.75 32.20 24.04 26.30 25.85 25.16
10-shot 0.742 32.48 34.19 33.33 23.72 25.00 25.85 24.84
after SFT zero-shot 0.741 32.44 34.69 32.87 22.76 27.92 25.51 22.90
Llama-3.2-1B- Instruct baseline zero-shot 0.700 40.43 27.86 31.71 25.96 35.39 29.59 25.81
1-shot 0.638 50.31 22.36 27.33 24.68 38.31 40.14 29.68
5-shot 0.658 46.63 23.37 30.00 27.24 38.64 31.63 29.03
10-shot 0.681 42.03 25.73 32.25 27.88 33.12 31.97 25.16
after SFT zero-shot 0.738 32.69 33.23 34.09 22.44 27.92 25.85 23.23
Llama-3.2-3B baseline zero-shot 0.626 58.81 18.76 22.43 27.24 41.56 47.96 35.48
1-shot 0.581 67.21 13.80 18.99 33.65 39.61 44.90 42.90
5-shot 0.526 79.94 10.71 9.35 37.18 38.64 51.02 46.77
10-shot 0.512 83.41 9.30 7.29 38.78 36.69 50.34 50.32
after SFT zero-shot 0.636 47.51 26.68 25.81 22.76 36.36 35.37 30.32
Llama-3.2-3B- Instruct baseline zero-shot 0.512 79.26 9.91 10.84 31.09 36.69 65.65 35.16
1-shot 0.466 80.07 10.22 9.71 28.53 30.19 60.88 35.16
5-shot 0.509 76.64 8.88 14.49 30.77 29.55 67.35 34.52
10-shot 0.509 74.61 7.79 17.60 32.05 28.25 65.65 31.94
after SFT zero-shot 0.418 58.82 10.25 30.93 17.63 18.18 48.64 18.06
Table 31: Error rates, incorrect choice distributions, and local negation confusion rates for the Llama3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B-v0.3 baseline zero-shot 0.533 75.74 20.24 4.02 25.96 31.82 60.20 49.35
1-shot 0.533 78.27 18.15 3.57 29.81 33.77 59.18 50.00
5-shot 0.427 82.71 15.80 1.49 22.12 25.65 51.02 47.42
10-shot 0.394 84.31 13.28 2.41 21.47 24.68 47.96 43.55
after SFT zero-shot 0.217 86.81 12.09 1.10 8.97 12.66 26.53 29.68
Mistral-7B-Instruct- v0.3 baseline zero-shot 0.389 65.10 34.08 0.82 16.67 18.83 29.25 39.68
1-shot 0.350 71.66 27.44 0.91 16.35 18.18 33.33 35.81
5-shot 0.305 72.47 27.27 0.26 12.50 15.58 28.91 34.52
10-shot 0.282 73.24 25.92 0.85 12.82 13.96 26.19 32.26
after SFT zero-shot 0.207 83.91 14.18 1.92 8.01 11.04 24.49 28.39
Mistral-Nemo- Base-2407 (12B) baseline zero-shot 0.540 75.62 21.44 2.94 25.64 33.44 60.20 50.00
1-shot 0.509 76.48 20.72 2.80 27.24 30.19 56.80 47.10
5-shot 0.399 81.11 16.70 2.19 21.15 23.70 42.86 46.13
10-shot 0.346 82.11 16.74 1.15 18.91 21.75 32.65 43.87
Mistral-Nemo- Instruct-2407 (12B) baseline zero-shot 0.513 75.43 21.64 2.94 24.04 29.22 56.80 50.32
1-shot 0.468 76.95 21.36 1.69 23.08 26.30 52.38 47.42
5-shot 0.369 78.28 20.43 1.29 19.23 20.45 38.44 41.29
10-shot 0.324 79.41 19.36 1.23 17.31 17.53 30.61 40.65
Mistral-Small-24B- Base-2501 baseline zero-shot 0.526 72.70 21.57 5.73 24.68 29.87 60.20 43.87
1-shot 0.502 78.52 18.48 3.00 26.60 29.87 58.16 48.71
5-shot 0.394 82.70 14.69 2.62 20.51 24.35 45.24 44.84
10-shot 0.348 83.37 14.58 2.05 17.63 21.75 37.76 42.90
Mistral-Small-24B- Instruct-2501 baseline zero-shot 0.474 78.43 19.57 2.01 25.96 29.87 53.06 45.16
1-shot 0.426 79.70 18.81 1.49 23.40 24.03 48.30 44.84
5-shot 0.355 81.43 17.67 0.89 18.91 20.78 38.10 41.61
10-shot 0.309 83.55 15.68 0.77 16.03 18.51 31.63 40.32
Table 32: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B-v0.3 baseline zero-shot 0.516 51.23 17.08 31.69 24.68 20.78 37.41 26.45
1-shot 0.521 53.42 10.20 36.38 27.88 23.38 36.73 27.10
5-shot 0.370 73.61 7.73 18.67 25.96 21.43 31.63 33.23
10-shot 0.336 73.82 7.78 18.40 25.00 19.81 26.19 31.29
after SFT zero-shot 0.520 54.57 24.85 20.58 27.88 29.87 33.67 25.81
Mistral-7B-Instruct- v0.3 baseline zero-shot 0.343 77.37 14.32 8.31 26.28 20.78 34.35 28.39
1-shot 0.355 69.42 20.09 10.49 28.53 23.05 29.25 20.97
5-shot 0.347 71.00 17.81 11.19 25.64 22.08 31.29 22.90
10-shot 0.351 72.46 20.77 6.77 24.68 22.08 32.65 25.81
after SFT zero-shot 0.308 77.58 17.01 5.41 23.72 19.48 35.37 20.32
Mistral-Nemo- Base-2407 (12B) baseline zero-shot 0.588 64.10 15.25 20.65 35.90 33.44 49.66 36.77
1-shot 0.488 83.41 10.24 6.34 39.10 33.12 54.42 41.61
5-shot 0.438 89.49 6.34 4.17 39.10 32.47 48.98 41.29
10-shot 0.400 89.88 6.15 3.97 33.97 29.55 40.14 44.52
Mistral-Nemo- Instruct-2407 (12B) baseline zero-shot 0.396 89.38 7.82 2.81 33.33 25.32 52.38 35.48
1-shot 0.368 89.22 8.62 2.16 30.77 24.68 46.94 33.55
5-shot 0.345 89.89 8.51 1.61 32.05 25.65 39.80 30.65
10-shot 0.338 90.85 7.75 1.41 32.69 29.22 35.37 29.35
Mistral-Small-24B- Base-2501 baseline zero-shot 0.432 62.20 18.72 19.08 29.81 21.75 39.12 20.65
1-shot 0.381 73.75 19.38 6.88 32.69 25.65 32.65 24.84
5-shot 0.274 85.80 9.86 4.35 25.32 20.13 21.77 29.35
10-shot 0.219 90.58 6.88 2.54 23.40 17.86 9.86 30.00
Mistral-Small-24B- Instruct-2501 baseline zero-shot 0.314 78.03 15.91 6.06 27.24 19.81 31.63 22.58
1-shot 0.290 81.15 14.48 4.37 30.77 18.83 27.89 19.68
5-shot 0.257 86.11 11.42 2.47 30.13 23.38 17.01 20.32
10-shot 0.217 89.05 7.66 3.28 27.24 19.81 8.50 23.55
Table 33: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B-v0.3 baseline zero-shot 0.521 74.12 20.24 5.63 26.60 32.14 55.44 45.81
1-shot 0.530 78.29 17.81 3.89 29.81 32.47 61.22 48.39
5-shot 0.432 82.94 15.23 1.83 23.08 24.35 52.38 48.71
10-shot 0.398 83.07 14.54 2.39 21.47 24.35 47.28 43.87
after SFT zero-shot 0.208 87.02 11.45 1.53 9.62 12.34 25.51 27.42
Mistral-7B-Instruct- v0.3 baseline zero-shot 0.338 68.78 30.52 0.70 17.31 18.18 23.47 36.77
1-shot 0.337 69.88 28.94 1.18 15.06 19.16 31.63 31.61
5-shot 0.296 73.19 26.54 0.27 13.14 16.56 28.57 31.29
10-shot 0.282 73.03 26.12 0.84 13.14 13.64 25.85 32.58
after SFT zero-shot 0.205 83.40 15.06 1.54 8.01 9.42 23.81 29.68
Mistral-Nemo- Base-2407 (12B) baseline zero-shot 0.535 71.81 24.04 4.15 25.96 30.84 53.74 48.39
1-shot 0.504 75.28 20.79 3.94 27.88 30.19 53.40 45.48
5-shot 0.409 78.68 18.41 2.91 21.47 24.35 42.52 44.84
10-shot 0.355 81.43 17.00 1.57 19.87 22.40 33.33 43.55
Mistral-Nemo- Instruct-2407 (12B) baseline zero-shot 0.489 71.47 25.28 3.24 24.36 29.55 48.30 42.58
1-shot 0.450 75.35 22.36 2.29 21.79 26.95 45.58 46.13
5-shot 0.362 78.73 19.96 1.32 18.91 20.45 36.05 42.26
10-shot 0.326 79.56 19.46 0.97 16.99 19.16 30.27 40.65
Mistral-Small-24B- Base-2501 baseline zero-shot 0.484 70.49 23.44 6.07 25.00 31.82 50.00 34.52
1-shot 0.466 76.32 18.91 4.77 25.64 28.25 54.08 39.35
5-shot 0.385 81.24 15.26 3.51 19.23 23.70 41.84 44.52
10-shot 0.339 83.14 14.99 1.87 17.31 21.10 36.05 41.94
Mistral-Small-24B- Instruct-2501 baseline zero-shot 0.420 76.94 20.60 2.46 25.32 32.79 44.22 31.29
1-shot 0.360 77.97 20.04 1.98 21.79 24.03 38.78 31.61
5-shot 0.328 82.57 16.95 0.48 17.95 19.48 35.03 39.35
10-shot 0.297 82.62 16.31 1.07 14.74 17.21 30.27 39.03
Table 34: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
Mistral-7B-v0.3 baseline zero-shot 0.498 59.39 15.13 25.48 24.04 25.97 43.88 28.71
1-shot 0.543 48.03 9.64 42.34 26.28 18.51 36.39 26.77
5-shot 0.364 71.02 7.63 21.35 24.36 20.45 30.95 30.97
10-shot 0.339 72.20 7.71 20.09 22.76 20.13 25.85 32.26
after SFT zero-shot 0.523 51.75 23.37 24.89 26.92 28.25 30.95 25.48
Mistral-7B-Instruct- v0.3 baseline zero-shot 0.366 80.04 10.63 9.33 31.73 20.78 35.71 32.58
1-shot 0.356 71.71 16.48 11.80 26.60 24.35 30.61 23.87
5-shot 0.355 69.42 16.07 14.51 24.68 21.10 31.97 24.19
10-shot 0.355 71.21 16.52 12.28 25.32 23.05 31.97 24.19
after SFT zero-shot 0.290 80.27 15.62 4.11 23.72 20.45 30.61 21.29
Mistral-Nemo- Base-2407 (12B) baseline zero-shot 0.578 66.67 14.54 18.79 37.18 40.91 47.62 33.55
1-shot 0.497 82.30 11.16 6.54 38.78 35.06 55.10 40.32
5-shot 0.443 88.35 5.91 5.73 38.46 30.52 49.32 43.23
10-shot 0.400 89.09 6.75 4.17 33.97 28.25 40.48 44.19
Mistral-Nemo- Instruct-2407 (12B) baseline zero-shot 0.374 85.56 10.83 3.61 29.49 27.92 42.86 31.94
1-shot 0.364 86.93 11.33 1.74 30.13 27.92 39.12 33.55
5-shot 0.344 91.01 7.83 1.15 31.41 26.95 39.80 31.29
10-shot 0.340 90.91 8.16 0.93 34.29 27.92 36.05 29.35
Mistral-Small-24B- Base-2501 baseline zero-shot 0.425 61.57 17.91 20.52 31.41 25.00 29.25 22.26
1-shot 0.357 70.00 21.78 8.22 29.81 20.13 29.59 23.55
5-shot 0.265 83.23 10.78 5.99 24.04 20.13 16.33 30.00
10-shot 0.215 89.67 6.64 3.69 22.12 17.53 8.84 30.32
Mistral-Small-24B- Instruct-2501 baseline zero-shot 0.271 80.41 13.16 6.43 32.37 22.73 11.22 22.90
1-shot 0.243 83.01 13.4 3.59 27.56 21.43 14.97 18.71
5-shot 0.240 87.09 9.93 2.98 27.88 21.10 12.93 23.55
10-shot 0.195 87.40 9.35 3.25 24.36 16.88 4.76 23.55
Table 35: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
GPT-4o mini baseline zero-shot 0.287 0 68.23 31.77 0 19.87 11.36 31.97 18.06
1-shot 0.251 0 68.99 30.38 0.63 20.51 10.71 26.87 13.55
5-shot 0.231 0 54.64 45.02 0.34 18.27 5.52 17.69 10.65
10-shot 0.205 0 52.51 47.10 0.39 16.67 7.14 11.90 8.71
GPT-4o baseline zero-shot 0.218 0 80.73 19.27 0 18.91 8.44 33.67 12.26
1-shot 0.209 0 88.26 11.36 0.38 23.08 8.77 27.55 17.10
5-shot 0.198 0 87.55 12.05 0.40 22.12 11.36 17.35 20.32
10-shot 0.175 0 85.45 14.09 0.45 20.19 10.39 12.24 18.39
GPT-4.1 mini baseline zero-shot 0.244 0 90.88 9.12 0 14.42 5.19 51.36 21.61
1-shot 0.204 0 88.33 11.67 0 18.59 7.14 30.27 18.71
5-shot 0.159 0 84.00 15.50 0.50 15.71 5.52 15.31 18.39
10-shot 0.151 0 77.89 21.58 0.53 13.78 5.84 8.84 19.68
GPT-4.1 baseline zero-shot 0.119 0 78.67 20.00 1.33 10.90 6.49 15.65 5.81
1-shot 0.136 0 81.87 18.13 0 17.31 9.09 10.88 8.39
5-shot 0.128 0 81.99 18.01 0 18.27 9.42 5.10 10.00
10-shot 0.109 0 79.71 19.57 0.72 15.06 8.44 2.04 10.00
Haiku 4.5 baseline zero-shot 0.228 0 88.54 11.46 0 22.12 15.26 38.78 8.06
1-shot 0.226 0 89.12 9.82 1.05 24.36 15.26 32.65 11.29
5-shot 0.190 0 90.79 9.21 0 29.49 17.53 13.61 10.00
10-shot 0.172 0 91.71 7.83 0.46 25.64 17.86 7.48 13.55
Sonnet 4.5 baseline zero-shot 0.125 0 85.35 12.74 1.91 16.35 14.61 8.50 4.19
1-shot 0.122 0 92.86 6.49 0.65 19.23 16.56 6.12 4.52
5-shot 0.127 0 81.25 12.50 6.25 16.67 12.99 4.76 7.74
10-shot 0.118 15 82.09 10.45 7.46 13.46 9.42 5.44 7.42
Table 36: Error rates, incorrect choice distributions, and local negation confusion rates for the API models under zero-shot and few-shot conditions, using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
Local
Negation
Contra-
diction
Para-
phrase
Relative
Clause
Participle
Clause
Compound
Sentence
Adverbial
Clause
GPT-4o mini baseline zero-shot 0.244 0 82.14 17.21 0.65 14.42 12.66 38.10 18.39
1-shot 0.226 0 81.05 17.89 1.05 19.55 8.12 33.33 15.16
5-shot 0.224 0 63.60 35.69 0.71 19.55 4.55 19.73 15.16
10-shot 0.210 0 60.38 39.25 0.38 16.99 7.14 15.65 12.58
GPT-4o baseline zero-shot 0.137 0 74.57 25.43 0 12.82 7.47 12.24 9.68
1-shot 0.147 0 79.46 20.00 0.54 18.27 10.06 13.27 6.45
5-shot 0.172 0 84.79 13.82 1.38 19.87 13.31 12.93 13.87
10-shot 0.185 0 83.69 15.45 0.86 24.68 12.01 11.22 15.48
GPT-4.1 mini baseline zero-shot 0.098 0 90.32 9.68 0 3.53 3.57 16.67 13.23
1-shot 0.133 0 84.52 15.48 0 14.42 5.84 18.37 8.06
5-shot 0.124 1 83.23 15.48 1.29 12.50 4.87 11.56 13.23
10-shot 0.130 0 75.61 23.78 0.61 12.50 4.22 11.56 12.26
GPT-4.1 baseline zero-shot 0.064 0 75.31 24.69 0 6.73 3.90 6.12 3.23
1-shot 0.082 0 91.26 8.74 0 13.14 8.12 4.42 4.84
5-shot 0.098 0 81.30 18.70 0 13.14 9.09 3.74 6.45
10-shot 0.098 0 84.68 15.32 0 13.78 8.44 1.70 10.00
Haiku 4.5 baseline zero-shot 0.135 1 89.94 10.06 0 16.99 15.58 13.27 3.87
1-shot 0.128 0 87.58 9.32 3.11 16.67 12.01 11.56 5.81
5-shot 0.126 0 93.71 5.66 0.63 22.12 14.29 6.80 5.16
10-shot 0.118 0 94.63 5.37 0 19.55 13.64 5.10 7.42
Sonnet 4.5 baseline zero-shot 0.085 1 91.51 7.55 0.94 10.58 12.34 6.12 2.58
1-shot 0.083 0 91.43 8.57 0 12.50 13.96 2.04 2.58
5-shot 0.121 26 82.68 14.17 3.15 14.74 12.01 3.74 3.55
10-shot 0.188 150 75.86 13.79 10.34 8.65 7.14 1.70 3.87
Table 37: Error rates, incorrect choice distributions, and local negation confusion rates for the API models under zero-shot and few-shot conditions, using detailed instruction.